### v1

# CodeCrash: Stress Testing LLM Reasoning under Structural and Semantic Perturbations

Abstract
Large Language Models (LLMs) have recently showcased strong capabilities in code-related tasks, yet their robustness in code comprehension and reasoning remains underexplored. In this paper, we present CodeCrash, a unified benchmark that evaluates LLM robustness under code structural and textual distraction perturbations, applied to two established benchmarks—CRUXEval and LiveCodeBench—across both input and output prediction tasks. We evaluate seventeen LLMs using direct and Chain-of-Thought inference to systematically analyze their robustness, identify primary reasons for performance degradation, and highlight failure modes. Our findings reveal the fragility of LLMs under structural noise and the inherent reliance on natural language cues, highlighting critical robustness issues of LLMs in code execution and understanding. Additionally, we examine three Large Reasoning Models (LRMs) and discover the severe vulnerability of self-reflective reasoning mechanisms that lead to reasoning collapse. CodeCrash provides a principled framework for stress-testing LLMs in code understanding, offering actionable directions for future evaluation and benchmarking. The code of CodeCrash and the robustness leaderboard are publicly available at https://donaldlamnl.github.io/CodeCrash/.
1 Introduction
Large Language Models (LLMs), exemplified by GPT-4 OpenAI et al. (2024), and the DeepSeek series DeepSeek-AI et al. (2025b, a), have recently exhibited remarkable performance across diverse code-related tasks, including code generation Zan et al. (2023); Xiao et al. (2024), program repair Fan et al. (2023); Liu et al. (2025), program testing Deng et al. (2023); Kang et al. (2023), and test case generation Steenhoek et al. (2025); Takerngsaksiri et al. (2025). Recent studies Wang et al. (2022); Mastropaolo et al. (2023); Chen et al. (2024a) have primarily evaluated LLM robustness under prompt-level perturbation, focusing on variations in task descriptions for code generation. While insightful for prompt engineering, these studies overlook robustness in deeper code comprehension tasks, such as understanding execution logic, structural reasoning, and program semantics.
In practical software development, LLMs are increasingly integrated into IDEs like GitHub Copilot [^1] https://github.com/features/copilot for tasks including code completion, automatic bug fixing, and test generation. Those integrations underscore the potential of LLMs to automate aspects of software development, improving code quality and developer productivity Zhang et al. (2023). Nevertheless, real-world production codebases are often complex and messy, frequently containing outdated comments, placeholder code, ambiguous identifiers, and non-functional segments, which introduce risk if LLMs overly depend on superficial cues rather than deep logical understanding.
CRUXEval (CRUXGu et al. (2024) introduces input and code execution tasks (i.e., output prediction) to evaluate LLMs’ comprehension and reasoning about code behavior. Building upon this foundation, we systematically evaluate LLM robustness under controlled code perturbations designed to challenge models’ reasoning capabilities beyond surface-level situations. Inspired by the program structure-consistent mutations proposed in CCTest Li et al. (2023), we extend this concept into a comprehensive robustness evaluation framework. Our unified benchmark, CodeCrash, systematically evaluates LLM robustness by applying two complementary perturbation families: I. Code Structural Perturbations, which introduce logically equivalent but structurally complicated code modifications, such as renaming entities, reformatting conditional expressions, and inserting non-functional “garbage” code segments; and II. Textual Distraction Perturbations, which insert misleading natural language (NL) cues that contradict actual program logic, including embedding misleading NL comments or statements and incorrect output hints.
In this paper, we conduct extensive perturbation experiments on two established benchmarks: CRUX and LiveCodeBench (LCBJain et al. (2025). We evaluated the robustness of seventeen LLMs across distinct perturbations to assess their deep understanding of code behavior and functionality in input prediction and code execution tasks. Additionally, we apply Chain-of-Thought (CoT) prompting Kojima et al. (2022) to explore how explicit step-by-step reasoning enhances model performance and robustness in code comprehension. We further evaluate three state-of-the-art Large Reasoning Models (LRMs) with powerful internal reasoning processes, uncovering significant vulnerabilities arising from their self-reflection mechanisms. Our key findings reveal that:
  • LLMs struggle to strictly follow actual code behavior under structural perturbations, frequently distracted by irrelevant or non-functional segments.
  • LLMs inherently prioritize NL cues over actual program logic, heavily relying on identifiers, comments, and embedded messages, adopting superficial comprehension and shortcut reasoning.
  • Explicit step-by-step reasoning boosts performance and robustness; however, internal reasoning mechanisms may trigger severe overthinking, infinite self-reflection, and re-verification.
Our contributions include:
  • We introduce a comprehensive perturbation framework CodeCrash encompassing structural complexity and textual distractions to evaluate the code-level robustness of LLMs.
  • We provide novel insights on evaluating and interpreting code reasoning in LLMs, revealing that they exhibit superficial NL dependence and demonstrating how plausible but incorrect hints trigger model collapsing.
2 Related Work
Researchers have significantly contributed to developing benchmarks for evaluating LLM capabilities on code-related tasks. Early benchmarks such as HumanEval Chen et al. (2021), MBPP Austin et al. (2021a), APPS Hendrycks et al. (2021), and HumanEval+ Liu et al. (2023b) primarily focused on assessing code generation performance. More recent efforts have shifted toward practical software engineering and realistic coding scenarios. For instance, SWE-bench Jimenez et al. (2024), InfiBench Li et al. (2024), and LiveCodeBench Jain et al. (2025) evaluate LLMs on complex tasks like multi-file changes, Stack Overflow-style questions, and real-world programming problems.
Meanwhile, a growing number of works investigate LLMs in code execution. Early studies Austin et al. (2021b); Nye et al. (2021) indicated that LLMs face challenges in accurately executing code but can benefit from intermediate reasoning steps. Subsequent research has utilized various strategies to enhance execution performance, including execution-informed pretraining Liu et al. (2023a), CoT prompting with trace supervision Ni et al. (2024), and iterative instruction prompting Lyu et al. (2024). CruxEval Gu et al. (2024) formalized input and output prediction tasks specifically for evaluating LLMs’ code understanding and reasoning, while LiveCodeBench extended these assessments to realistic programs written by humans. REval Chen et al. (2024b) and CACP Hooda et al. (2024) uncovered inconsistencies and concept-level misunderstandings of LLMs in code through intermediate reasoning analysis in code execution. CodeScore Dong et al. (2024), LEVER Ni et al. (2023), and XCODEEval Khan et al. (2024) have leveraged code execution to improve model evaluation methodologies and enhance the performance of downstream tasks.
Robustness evaluation is another research direction in assessing LLMs for code. Prior studies focused on prompt-level variations in code generation, examining how semantically equivalent NL descriptions impact model performance Mastropaolo et al. (2023); Chen et al. (2024a); Zhuo et al. (2023). Apart from NL perturbations, researchers have investigated model robustness under programming language perturbations. For example, CCTest Li et al. (2023) introduced structural mutations, CREAM Gao et al. (2023) employed counterfactual reasoning, and ReCode Wang et al. (2022) applied semantic-preserving transformations, advancing the robustness in code comprehension and completion. BigCodeBench Zhuo et al. (2025) further explored generalization under complex function composition and tool usage. Furthermore, Chen et al. (2025); Cuadron et al. (2025) and Kumar et al. (2025) have recently revealed that LRMs frequently suffer from overthinking and produce unnecessary analysis and unstable thoughts in math-related works.
3 Methodology
This section outlines our benchmark methodology, including dataset selection, benchmark pipeline, perturbation strategies, and experimental settings.
3.1 Benchmark Tasks and Datasets
We evaluate two benchmarks: CRUX Gu et al. (2024) and LCB Jain et al. (2025). CRUX contains 800 synthetic problems generated by Code-LLaMA-34B, each featuring concise Python snippets with generic function names f and a corresponding input-output pair for input prediction and code execution tasks. In contrast, LCB contains 479 real-world coding problems sourced from 85 human-submitted LeetCode solutions, focusing solely on code execution. While both datasets avoid excessive memory or runtime demands, LCB problems exhibit more complex structures, such as nested functions, decorators, and external libraries, requiring more computations.
3.2 Benchmark Execution Pipeline
Figure 1: Overview of CodeCrash pipeline. Code samples are first standardized into a clean vanilla version. Perturbations are visualized with yellow highlights. Each perturbed code is validated and executed before passing to LLMs.
Our benchmark follows a unified evaluation pipeline (Figure 1) across all perturbation types. Each original code snippet is initially processed via Abstract Syntax Tree (AST) parsing and unparsing to remove comments and standardize formatting, producing a clean baseline “vanilla” version (VAN). Subsequently, we apply a selected perturbation and regenerate the code using AST unparsing to maintain formatting consistency (except for comment-based perturbations). Each perturbed code snippet is executed with the provided inputs to ensure syntactic validity and functional correctness. The validated perturbed code is then provided to LLMs, and we record all execution results and evaluate them using the original evaluation methods from the dataset to compute the PASS@1 score.
3.3 Perturbation Strategies
We design two categories of perturbations to evaluate model robustness comprehensively: (1) Code Structural Perturbations, which modify code structure without affecting functionality, and (2) Textual Distraction Perturbations, which inject misleading NL cues and hints. Detailed examples and principles are provided in Appendix B.
3.3.1 Code Structural Perturbations
(1) Renaming Entities (REN).
[…]
(2) Reformatting Conditional Expressions (RTF).
We rewrite conditional expressions within “if” and “while” statements into logically equivalent but structurally complex and indirect forms. We use 16 predefined transformation templates to reformat comparison, boolean, and other conditional expressions. This perturbation evaluates the LLMs’ ability to maintain logical consistency and reasoning under increased complexity.
(3) Inserting Garbage Code (GBC).
We insert syntactically valid but functionally irrelevant “garbage” code segments to challenge the models’ robustness against code structural distractions. Three specific types include: (1) repeated global variable declarations mirroring local identifiers, (2) non-executable conditional dead branches, and (3) uninvoked infinite recursive function definitions.
(4) Aggregating Structural Perturbations (ALL).
We comprehensively aggregate all structural perturbations (REN, RTF, GBC) together and create a highly adversarial scenario, thoroughly challenging LLM robustness across multiple structural perturbation dimensions simultaneously.
3.3.2 Textual Distraction Perturbations
(1) Inserting Misleading Descriptive Comments (MDC).
We insert misleading NL comments for eight key AST nodes (e.g., function definitions, returns, loops, conditionals) to evaluate LLM robustness against textual cues and contradictory information. We initially instruct GPT-4o to generate those messages and manually filter them to ensure they are generic, explicitly incorrect, and contradict the actual code logic. Additionally, we introduce a probability parameter p to vary comment density and analyze its impact and validity on model robustness in § 5.1.
(2) Inserting Misleading Print Statements (MPS).
Similar to MDC perturbation, we inject misleading NL messages as executable print statements. Unlike comments that AST-based heuristics can easily filter out, print statements are logically irrelevant to functionality and should not be perceived as semantic guides. Allowing us to directly compare the behavior with MDC and distinguish models’ NL reliance in different embedding formats.
(3) Inserting Misleading Hint Comments (MHC).
We instruct GPT-4o to generate logically plausible but incorrect hints that contradict the actual program behavior and the ground truth. After verifying their incorrectness, they are strategically inserted into function definitions or return statements using AST parsing.
3.4 Experiment Settings
Model Selection.
We evaluate 17 models of varying sizes and versions, including both open-source and closed-source LLMs: GPT family (GPT-4o-Mini, GPT-4o), Anthropic Claude (Claude-3.5-Haiku, 3.5-Sonnet), Google Gemini (Gemini-1.5-Flash, 2.0-Flash, 1.5-Pro), DeepSeek-V3, Alibaba Qwen (Qwen2.5-7B-Ins, 14B-Ins, 32B-Ins, 72B-Ins, 32B-Coder), and Meta LLaMA (LLaMA-3.1-8B-Ins, 70B-Ins, 405B-Ins, LLaMA-3.3-70B). Additionally, we assess three cutting-edge LRMs: o3-mini (low and high reasoning efforts), DeepSeek-R1, and QwQ-32B.
Model Configurations.
Following the settings in LCB, we use nucleus sampling (temperature=0.2, top-p=0.95). We set the maximum tokens per response at 200 for direct inference and 2,000 for CoT inference, discarding excessively long outputs. For o3-mini, we set the maximum completion tokens at 20,000 to avoid verbose reasoning. Due to resource constraints, we generate three candidate answers for direct inference and one candidate for CoT inference per query. We adopt prompt formats similar CRUX, employing 2-shot prompting for direct inference and 1-shot CoT prompting (details provided in Appendix A). Primary evaluation uses PASS@1 accuracy Chen et al. (2021), reporting both absolute and relative degradation metrics.
4 Experimental Results
This section presents a comprehensive analysis of model robustness under the perturbation strategies introduced in §3.3. We assess the robustness of LLMs across different perturbation types, model scales, reasoning modes, and prediction tasks.
4.1 Impact on Output Prediction
We first evaluate LLM robustness in output prediction tasks. Table 1 summarizes the absolute PASS@1 degradation for all models under direct inference across both datasets, highlighting how each perturbation affects model performance.
Table 1: Absolute PASS@1 degradation of LLMs under all perturbation types for output prediction tasks across CRUX and LCB in direct inference (N=3) [TO RELATIVE SCORE].
Model Series Model Name VAN REN RTF GBC ALL MDC MPS MHC
Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
GPT-4 Omni GPT-4o-Mini $57.29$ $53.24$ $-0.62$ $-9.67$ $-1.79$ $-0.97$ $-10.92$ $-5.50$ $-15.96$ $-16.98$ $-18.67$ $-21.50$ $-13.29$ $-14.89$ $-6.62$ $-23.59$
GPT-4o $71.33$ $64.51$ $-2.71$ $-9.60$ $-3.38$ $-4.59$ $-9.21$ $-3.20$ $-10.71$ $-18.30$ $-9.96$ $-18.79$ $-12.29$ $-15.66$ $-4.54$ $-17.19$
Claude Series Claude-3.5-Haiku-20241022 $57.42$ $58.25$ $-3.96$ $-9.32$ $-1.79$ $+0.35$ $-7.29$ $-4.52$ $-13.79$ $-21.50$ $-7.92$ $-5.85$ $-6.46$ $-4.66$ $-15.79$ $-31.04$
Claude-3.5-Sonnet-20241022 $71.46$ $73.83$ $-1.67$ $-16.77$ $-0.92$ $+1.18$ $-6.88$ $-4.94$ $-10.54$ $-25.19$ $-5.79$ $-7.10$ $-6.17$ $-7.86$ $-10.29$ $-32.08$
Gemini Series Gemini-1.5-Flash $56.21$ $44.26$ $-3.54$ $-0.42$ $-1.17$ $-1.04$ $-4.58$ $-0.35$ $-10.33$ $-5.64$ $-8.29$ $-14.27$ $-16.00$ $-17.88$ $-4.12$ $-9.81$
Gemini-2.0-Flash $65.04$ $65.97$ $-2.33$ $-18.09$ $-2.08$ $-0.77$ $-15.50$ $-8.77$ $-21.63$ $-27.77$ $-17.42$ $-34.24$ $-21.54$ $-33.19$ $-5.38$ $-12.87$
Gemini-1.5-Pro-002 $67.37$ $56.16$ $-2.92$ $-3.41$ $-2.50$ $-4.73$ $-5.62$ $-3.83$ $-13.42$ $-13.01$ $-15.50$ $-18.51$ $-14.62$ $-18.44$ $-6.75$ $-18.23$
DeepSeek DeepSeek-V3 $67.87$ $67.78$ $-2.67$ $-16.35$ $-1.33$ $-3.76$ $-4.62$ $-5.98$ $-8.75$ $-24.15$ $-11.25$ $-28.04$ $-6.87$ $-23.17$ $-7.29$ $-20.11$
LLaMA Series LLaMA-3.1-8B-Instruct $35.96$ $34.66$ $+0.08$ $-4.31$ $-0.33$ $-1.67$ $-3.13$ $-4.73$ $-8.54$ $-6.82$ $-8.08$ $-9.88$ $-6.42$ $-8.00$ $-6.00$ $-13.57$
LLaMA-3.1-70B-Instruct $56.08$ $44.89$ $-2.38$ $-1.39$ $-0.54$ $-1.74$ $-5.38$ $-2.44$ $-11.00$ $-7.72$ $-10.79$ $-12.53$ $-14.96$ $-17.40$ $-4.71$ $-14.47$
LLaMA-3.1-405B-Instruct $63.54$ $50.66$ $-5.37$ $-3.90$ $-0.62$ $-4.45$ $-5.79$ $-2.85$ $-12.50$ $-8.56$ $-4.17$ $-7.45$ $-7.17$ $-10.09$ $-4.37$ $-13.01$
LLaMA-3.3-70B-Instruct $59.92$ $48.50$ $-2.21$ $-1.39$ $-1.25$ $-3.41$ $-5.50$ $-5.71$ $-10.37$ $-9.74$ $-7.46$ $-9.60$ $-8.08$ $-12.32$ $-4.08$ $-10.16$
Qwen Series Qwen2.5-7B-Instruct $43.29$ $41.41$ $-3.04$ $-7.24$ $-1.29$ $-3.20$ $-12.96$ $-3.76$ $-16.42$ $-12.80$ $-25.13$ $-15.87$ $-19.58$ $-8.21$ $-11.67$ $-23.03$
Qwen2.5-14B-Instruct $47.79$ $49.48$ $-1.83$ $-9.88$ $-0.67$ $-1.39$ $-11.21$ $-3.41$ $-19.04$ $-14.82$ $-16.67$ $-20.46$ $-15.54$ $-16.77$ $-4.46$ $-11.00$
Qwen2.5-32B-Instruct $60.04$ $59.64$ $-1.71$ $-10.79$ $-2.58$ $-0.49$ $-5.79$ $-1.39$ $-11.88$ $-20.81$ $-10.83$ $-19.49$ $-14.21$ $-14.06$ $-7.87$ $-18.37$
Qwen2.5-72B-Instruct $60.08$ $54.91$ $-1.04$ $-8.49$ $-1.87$ $+0.42$ $-10.71$ $-2.02$ $-14.00$ $-13.85$ $-10.21$ $-6.96$ $-14.46$ $-14.41$ $-9.63$ $-22.27$
Qwen2.5-Coder-32B-Instruct $66.96$ $56.65$ $-2.50$ $-8.42$ $-0.96$ $-2.02$ $-10.75$ $-2.78$ $-14.96$ $-15.80$ $-15.54$ $-20.74$ $-11.08$ $-17.19$ $-7.00$ $-12.32$
Average $-2.38$ $-8.20$ $-1.47$ $-1.90$ $-7.99$ $-3.89$ $-13.17$ $-15.50$ $-11.98$ $-15.96$ $-12.28$ $-14.95$ $-7.09$ $-17.83$
$\mathbf{-4.56}$ $\mathbf{-1.63}$ $\mathbf{-6.46}$ $\mathbf{-14.04}$ $\mathbf{-13.47}$ $\mathbf{-13.28}$ $\mathbf{-11.11}$
4.1.1 Robustness to Code Structural Perturbations
This section focuses on how structural perturbations impact model performance, examining reliance on superficial cues, logical reasoning capabilities, and sensitivity to irrelevant code.
Reliance on NL Cues.
Under the REN perturbation, we consistently observe performance degradation across GPT, Claude, DeepSeek, and Qwen models, with more significant drops on LCB ($8.20$pp 222pp stands for percentage point. average) than CRUX ($2.38$pp average). Top-performing models, such as Claude-3.5-Sonnet and DeepSeek-V3, suffer severe degradations exceeding $\mathbf{16}$pp on LCB. The substantial difference between the two datasets is attributed to the semantically meaningful function identifiers present in LCB, which are absent in CRUX. These results indicate that LLMs substantially rely on surface-level NL cues from identifiers to understand code and predict outputs.
Robust to Logical Reformatting.
LLMs show strong robustness against RTF perturbation, with minimal performance drop on both datasets (overall PASS@1 drop of only $1.63$pp). This robustness indicates LLMs can reliably reason over logically equivalent but syntactically complex expressions, demonstrating strong logical consistency against superficial structural modifications.
Severe Impact of Garbage Code.
Under the GBC perturbation, a notable performance degradation emerges in CRUX. The average degradation is consistently larger on CRUX ($7.99$pp) compared to LCB ($3.89$pp) across most models, which is a pattern opposite to other perturbations. This discrepancy can be attributed to the presence of meaningful identifiers in LCB, which helps models mitigate the distracting effects to infer the output. In contrast, the lack of meaningful identifiers in CRUX makes LLMs more susceptible to distraction by irrelevant and non-functional code segments. Under the fairness comparison environment in ALL perturbations, the performance drop on LCB ($15.50$pp) is again greater than that on CRUX ($13.17$pp), underscoring the significant impact of garbage code on LLMs’ reasoning and understanding.
Compounding Effects of Structural Perturbations.
Combining structural perturbations (ALL) results in the most severe degradation, averaging $\mathbf{14.04}$pp, which surpasses any individual perturbation effects and demonstrates an explicit compounding effect. Notably, high-performing models such as GPT-4o, Claude-3.5-Sonnet, Gemini-2.0-Flash, and DeepSeek-V3 suffer degradations exceeding $\mathbf{18}$pp on LCB, underscoring a serious vulnerability to multiple structural perturbations.
Key Findings: LLMs rely heavily on superficial semantic cues from identifiers, struggle to identify non-functional segments from essential code, and exhibit compounded vulnerability when multiple structural distractions are combined. This combination leads to significant degradation ($\mathbf{14.04}$pp on average) in output prediction performance.
4.1.2 Robustness to Textual Distraction Perturbations
We next investigate the influence of textual distractions on LLM robustness, focusing on the susceptibility to misleading information and incorrect hints.
Vulnerability to Misleading Information.
Misleading textual perturbations (MDC and MPS) cause substantial performance declines, averaging degradations of $13.47$pp and $13.28$pp, respectively—nearly as severe as the ALL perturbation. These results indicate that LLMs are highly susceptible to misleading NL messages, frequently prioritizing superficial textual information over the actual code logic. This behavior is further observed in MHC perturbation, causing an average degradation of $11.11$pp. The effect is notably higher on LCB ($17.83$pp) compared to CRUX ($7.09$pp), exceeding all other perturbations on LCB. This discrepancy indicates a difficulty-dependent vulnerability, wherein increased complexity exacerbates models’ reliance on plausible but misleading hints, diminishing their logical following and verifying abilities.
Consistency Patterns in Model Sensitivity.
We identify two prominent consistency patterns regarding sensitivity to misleading NL messages from Table 1. First, 16 of 17 models exhibit a clear dataset-level consistency: if the performance degradation caused by MDC on CRUX is greater than that on LCB, then the same relationship holds for MPS and vice versa. This pattern indicates that model sensitivity to misleading NL messages primarily depends on dataset characteristics rather than perturbation format, revealing an inherent NL-dependence property of LLMs in reasoning processing. Second, 14 of 17 models show a clear perturbation-method consistency: if the performance degradation caused by MDC on CRUX exceeds that on LCB, then the same holds for MPS and vice versa. This pattern suggests an internal cognitive preference for specific formats (comments vs. print statements). While the degradation magnitude varies with the code, the format-specific sensitivity is consistently preserved across datasets, reflecting the intrinsic properties of LLMs related to their reasoning architecture and training corpus.
Key Findings: Textual perturbations significantly degrade model performance, revealing an inherent dependence on embedded NL cues and a internal cognitive preference for specific message formats. Moreover, LLMs exhibit an explicit difficulty-dependent vulnerability, increasingly relying on plausible but incorrect hints over actual program logic in complex scenarios.
4.1.3 Model Robustness Across Different Scales
Table 2: Relative PASS@1 degradation aggregated over CRUX and LCB under direct inference. Weighted Average column uses weights: 1/2 ALL, 1/8 MDC, 1/8 MPS, and 1/4 MHC.
Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average
GPT-4 Omni GPT-4o-Mini 55.77 $-29.37\%$ $-35.51\%$ $-24.99\%$ $-23.83\%$ $-28.20\%$
GPT-4o 68.78 $-20.01\%$ $-19.64\%$ $-19.87\%$ $-13.96\%$ $-18.44\%$
Claude Series Claude-3.5-Haiku-20241022 57.73 $-28.85\%$ $-12.38\%$ $-10.03\%$ $-37.16\%$ $-26.52\%$
Claude-3.5-Sonnet-20241022 72.35 $-22.01\%$ $-8.67\%$ $-9.39\%$ $-25.28\%$ $-19.58\%$
Gemini Series Gemini-1.5-Flash 51.73 $-16.27\%$ $-21.30\%$ $-32.94\%$ $-12.89\%$ $-18.14\%$
Gemini-2.0-Flash 65.39 $-36.56\%$ $-36.19\%$ $-39.56\%$ $-12.48\%$ $-30.87\%$
Gemini-1.5-Pro-002 63.17 $-21.13\%$ $-26.73\%$ $-25.88\%$ $-18.43\%$ $-21.75\%$
DeepSeek DeepSeek-V3 67.84 $-21.41\%$ $-25.86\%$ $-19.14\%$ $-17.83\%$ $-20.79\%$
LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-22.23\%$ $-24.74\%$ $-19.81\%$ $-25.10\%$ $-22.96\%$
LLaMA-3.1-70B-Instruct 51.89 $-18.71\%$ $-22.49\%$ $-31.20\%$ $-17.33\%$ $-20.40\%$
LLaMA-3.1-405B-Instruct 58.72 $-18.63\%$ $-9.61\%$ $-14.51\%$ $-13.93\%$ $-15.81\%$
LLaMA-3.3-70B-Instruct 55.64 $-18.35\%$ $-15.20\%$ $-17.95\%$ $-12.11\%$ $-16.35\%$
Qwen Series Qwen2.5-7B-Instruct 42.59 $-35.30\%$ $-50.65\%$ $-35.72\%$ $-37.69\%$ $-37.87\%$
Qwen2.5-14B-Instruct 48.42 $-36.14\%$ $-37.30\%$ $-33.04\%$ $-14.16\%$ $-30.40\%$
Qwen2.5-32B-Instruct 59.89 $-25.44\%$ $-23.52\%$ $-23.63\%$ $-19.74\%$ $-23.55\%$
Qwen2.5-72B-Instruct 58.14 $-24.02\%$ $-15.37\%$ $-24.88\%$ $-25.21\%$ $-23.34\%$
Qwen2.5-Coder-32B-Instruct 63.10 $-24.42\%$ $-28.23\%$ $-21.72\%$ $-14.68\%$ $-22.12\%$
We further analyze robustness across different model scales and versions, comparing relative PASS@1 degradation under representative perturbations in Table 2. The weighting details and alternative evaluation schemes are in Appendix C.2.
We observe a clear trend of increasing robustness with model scale in most families. Larger closed-sourced models from GPT-4 Omni and Claude families demonstrate enhanced robustness compared to their smaller counterparts. Similar trends appear in open-source models such as LLaMA (3.1-8B $<$ 70B $<$ 405B) and Qwen (2.5-7B $<$ 14B $<$ 32B $<$ 72B). Additionally, newer model versions (LLaMA-3.3-70B) and domain-specific fine-tuned variants (Qwen2.5-Coder-32B-Ins) consistently exhibit improved robustness compared to their predecessors (LLaMA-3.1-70B) and larger counterparts (Qwen2.5-72B-Ins), respectively.
Key Finding: Larger scales, newer versions, and domain-specific fine-tuning LLMs substantially enhance robustness, highlighting improved code understanding and reasoning capabilities.
4.2 Impact of Reasoning on Model Robustness
This section examines how the reasoning process, including CoT prompting for LLMs and internal reasoning in LRMs, affects model robustness against perturbations.
4.2.1 Effect of CoT Prompting
Table 3: Comparison of relative PASS@1 degradation aggregated over CRUX and LCB for output prediction tasks in direct inference (N=3) and CoT inference (N=1) [TO ABSOLUTE SCORE].
Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average
Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT
GPT-4 Omni GPT-4o-Mini $55.77$ $81.47$ $-29.30\%$ $-13.15\%$ $-35.37\%$ $-10.84\%$ $-24.91\%$ $-10.94\%$ $-23.27\%$ $-2.69\%$ $-28.00\%$ $-9.52\%$
GPT-4o $68.78$ $91.79$ $-19.70\%$ $-4.94\%$ $-19.29\%$ $-5.03\%$ $-19.70\%$ $-5.62\%$ $-13.49\%$ $-1.36\%$ $-18.10\%$ $-4.22\%$
Claude Series Claude-3.5-Haiku-20241022 $57.73$ $72.95$ $-28.89\%$ $-21.33\%$ $-12.37\%$ $-10.61\%$ $-10.02\%$ $-8.68\%$ $-37.25\%$ $-14.26\%$ $-26.56\%$ $-14.30\%$
Claude-3.5-Sonnet-20241022 $72.35$ $86.00$ $-22.15\%$ $-7.55\%$ $-8.68\%$ $-5.27\%$ $-9.40\%$ $-6.36\%$ $-25.50\%$ $-7.91\%$ $-19.71\%$ $-6.86\%$
Gemini Series Gemini-1.5-Flash $51.73$ $75.22$ $-16.57\%$ $-18.30\%$ $-20.35\%$ $-21.73\%$ $-32.29\%$ $-42.20\%$ $-12.09\%$ $-2.08\%$ $-17.89\%$ $-20.39\%$
Gemini-2.0-Flash $65.39$ $89.13$ $-36.59\%$ $-6.23\%$ $-36.27\%$ $-6.32\%$ $-39.62\%$ $-14.30\%$ $-12.51\%$ $-2.02\%$ $-30.91\%$ $-7.01\%$
Gemini-1.5-Pro-002 $63.17$ $87.18$ $-21.00\%$ $-7.44\%$ $-26.32\%$ $-11.93\%$ $-25.41\%$ $-14.71\%$ $-17.49\%$ $-3.86\%$ $-21.34\%$ $-9.19\%$
DeepSeek DeepSeek-V3 $67.84$ $89.52$ $-21.40\%$ $-9.96\%$ $-25.85\%$ $-17.73\%$ $-19.13\%$ $-18.78\%$ $-17.83\%$ $-3.49\%$ $-20.78\%$ $-12.07\%$
LLaMA Series LLaMA-3.1-8B-Instruct $35.47$ $44.72$ $-22.26\%$ $-27.62\%$ $-24.69\%$ $-20.98\%$ $-19.76\%$ $-21.50\%$ $-24.91\%$ $-8.74\%$ $-22.92\%$ $-20.03\%$
LLaMA-3.1-70B-Instruct $51.89$ $69.43$ $-18.83\%$ $-17.91\%$ $-22.05\%$ $-23.20\%$ $-30.59\%$ $-30.86\%$ $-16.12\%$ $-6.76\%$ $-20.03\%$ $-19.22\%$
LLaMA-3.1-405B-Instruct $58.72$ $78.42$ $-18.77\%$ $-13.56\%$ $-9.19\%$ $-10.27\%$ $-14.07\%$ $-15.55\%$ $-12.96\%$ $-7.48\%$ $-15.53\%$ $-11.75\%$
LLaMA-3.3-70B-Instruct $55.64$ $76.86$ $-18.22\%$ $-10.89\%$ $-14.85\%$ $-7.73\%$ $-17.38\%$ $-11.09\%$ $-11.43\%$ $-4.78\%$ $-16.00\%$ $-8.70\%$
Qwen Series Qwen2.5-7B-Instruct $42.59$ $58.17$ $-35.37\%$ $-21.77\%$ $-50.86\%$ $-21.91\%$ $-35.99\%$ $-27.96\%$ $-37.39\%$ $-6.99\%$ $-37.89\%$ $-19.50\%$
Qwen2.5-14B-Instruct $48.42$ $70.68$ $-36.06\%$ $-23.12\%$ $-37.35\%$ $-35.29\%$ $-33.05\%$ $-43.36\%$ $-14.26\%$ $-10.84\%$ $-30.40\%$ $-27.34\%$
Qwen2.5-32B-Instruct $59.89$ $79.44$ $-25.41\%$ $-17.42\%$ $-23.50\%$ $-18.60\%$ $-23.63\%$ $-24.21\%$ $-19.71\%$ $-5.31\%$ $-23.53\%$ $-16.19\%$
Qwen2.5-72B-Instruct $58.14$ $82.80$ $-23.98\%$ $-12.65\%$ $-15.46\%$ $-6.80\%$ $-24.83\%$ $-16.24\%$ $-24.70\%$ $-6.61\%$ $-23.20\%$ $-10.63\%$
Qwen2.5-Coder-32B-Instruct $63.10$ $84.75$ $-24.20\%$ $-14.67\%$ $-27.72\%$ $-19.00\%$ $-21.19\%$ $-19.93\%$ $-14.25\%$ $-9.32\%$ $-21.78\%$ $-15.49\%$
Average $\mathbf{-24.63}\%$ $\mathbf{-14.62}\%$ $\mathbf{-24.13}\%$ $\mathbf{-14.90}\%$ $\mathbf{-23.59}\%$ $\mathbf{-19.55}\%$ $\mathbf{-19.72}\%$ $\mathbf{-6.15}\%$ $\mathbf{-23.21}\%$ $\mathbf{-13.67}\%$
We first investigate how CoT enhances LLM robustness by code-level step-by-step execution and present the results in Table 3. It effectively enhances model robustness, with average improvements of $10.01$pp for ALL, $9.23$pp for MDC, $4.04$pp for MPS, and a notable $13.57$pp for MHC. It consistently reduces average performance degradation from approximately $24\%$ to $14\%$ on ALL and MDC, indicating that step-by-step reasoning enables LLMs to prioritize executable code semantics over irrelevant segments and cues. The most significant improvement on MHC suggests CoT enhances code reasoning and execution, mitigating over-reliance on misleading hints observed in direct inference in LCB4.1.2). These benefits are amplified in high-capacity models (e.g., GPT-4o, Claude-3.5-Sonnet), which exhibit greater reasoning capacity and noise resistance.
However, CoT does not fully eliminate the model’s reliance on misleading NL information. Performance degradation remains substantial under ALL, MDC, and MPS perturbations, especially for smaller open-sourced models (e.g., Qwen2.5 and LLaMA-3.1). These results highlight a limitation in distinguishing non-functional code segments and misleading information from the program. Interestingly, models tend to incorporate print statements directly into their reasoning, resulting in higher error rates on MPS. We provide detailed case studies for each perturbation in Appendix D (Case 1 to 3) to investigate their inference and illustrate their failure reasons.
Key Finding: CoT prompting significantly enhances LLMs’ robustness by improving reasoning ability and allowing step-by-step execution, especially for the high-capacity models. However, it does not fully mitigate reliance on misleading information and remains constrained by the model’s underlying capabilities.
4.2.2 Powerful Reasoning and Potential Vulnerability of LRMs
Table 4: Absolute PASS@1 degradation and reasoning token usage (Avg/Max) of LRMs on output prediction across CRUX and LCB (N=1).
Model Dataset VAN ALL MDC MPS MHC
PASS@1 Avg. Max. Diff. Avg. Max. Diff. Avg. Max. Diff. Avg. Max. Diff. Avg. Max.
o3-Mini-Low CRUX 97.62 $213$ 2560 $-1.75$ $379$ 2240 $-4.25$ $244$ 5184 $-1.00$ $272$ 3584 $-10.88$ $366$ 4864
LCB 98.96 $240$ 1024 $-0.63$ $543$ 1856 $-12.32$ $265$ 1088 $-3.55$ $280$ 1088 $-19.62$ $330$ 2752
o3-Mini-High CRUX 98.12 $1311$ 20000 $+0.12$ $2223$ 20000 $-3.50$ $2182$ 20000 $+0.88$ $2197$ 20000 $-13.12$ $3108$ 20000
LCB 100.00 $1084$ 8576 $-0.21$ $2632$ 8960 $-5.64$ $2136$ 8448 $-0.63$ $1972$ 7744 $-28.39$ $3767$ 20000
DeepSeek-R1 CRUX 95.38 $922$ 10471 $-1.25$ $1469$ 11191 $-3.38$ $1424$ 10807 $-0.38$ $1069$ 10037 $-2.25$ $2216$ 16021
LCB 99.79 $906$ 7322 $-1.25$ $1618$ 7749 $-2.71$ $1512$ 6157 $-0.42$ $1094$ 9378 $-0.63$ $2597$ 14870
QwQ-32B CRUX 93.25 $1391$ 13718 $-0.88$ $2088$ 12180 $-3.12$ $1814$ 10906 $-1.12$ $1875$ 11700 $-0.75$ $2657$ 18759
LCB 98.96 $1518$ 9206 $-0.21$ $2488$ 11232 $-4.59$ $1669$ 8991 $-1.88$ $1751$ 8816 $-1.04$ $3713$ 32769
We then evaluate the robustness and reasoning behaviors of three LRMs on output prediction tasks, with results summarized in Table 4. All LRMs exhibit remarkable performance and robustness, achieving over 90 PASS@1 accuracy in the VAN setting and having minor performance degradation under the ALL and MPS perturbations. These results suggest that internal reasoning helps models accurately identify and distinguish code segments that are functionally critical for execution, thereby outperforming all generic models with CoT prompting. Moreover, LRMs exhibit consistently stronger robustness against MPS than MDC, indicating that while they can correctly recognize and ignore misleading messages from print statements, they still inherently treat comments as valid execution guidance. This behavior reveals the stable cognitive preference toward specific injection formats, as observed in §3.3.2.
However, this residual cognitive bias leads to a critical vulnerability under the MHC perturbation. Concise and plausible but incorrect output hints can trigger severe reasoning failures in LRMs. We observed that: (1) All LRMs exhibit substantial increases in reasoning token usages under MHC (doubling the token usages in VAN). (2) In extreme cases, QwQ-32B generates over 32,000 tokens in a single task, caused by uncontrolled self-reflection and reasoning collapse (see §5.3 and Appendix D.5). (3) Excessive reasoning leads to more significant degradations in o3-Mini with higher reasoning effort. These observations reveal that logically plausible yet incorrect hints can trigger catastrophic overthinking behavior in LRMs, resulting in significant computational overhead, excessive self-reflection, and severe performance drop.
Key Finding: LRMs exhibit exceptional robustness under ALL and MPS, benefiting from strong internal reasoning capabilities. However, they retain a residual tendency to trust and analyze comments; thereby, plausible but incorrect information can trigger catastrophic overthinking, excessive self-verification, and severe performance degradation, highlighting a critical vulnerability of the internal reasoning mechanism in current LRMs.
4.3 Impact on Input Prediction
Table 5: Comparison of relative PASS@1 degradation under CRUX for input prediction tasks in direct inference (N=3) and CoT inference (N=1) [TO ABSOLUTE SCORE].
Model Series Model Name VAN ALL MDC MPS MHC
Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT
GPT-4 Omni GPT-4o-Mini $57.67$ $69.52$ $-29.84\%$ $-19.63\%$ $+1.30\%$ $+1.05\%$ $+1.73\%$ $+1.05\%$ $-30.49\%$ $-8.66\%$
GPT-4o $68.04$ $79.12$ $-13.29\%$ $-7.42\%$ $+5.63\%$ $+1.26\%$ $+1.41\%$ $+2.69\%$ $-8.33\%$ $-2.84\%$
Claude Series Claude-3.5-Haiku-20241022 $54.67$ $63.62$ $-17.00\%$ $-16.70\%$ $+7.01\%$ $+4.13\%$ $+7.77\%$ $+2.36\%$ $-45.35\%$ $-29.86\%$
Claude-3.5-Sonnet-20241022 $73.33$ $80.12$ $-15.91\%$ $-9.67\%$ $+0.11\%$ $-3.43\%$ $-1.70\%$ $-2.65\%$ $-24.83\%$ $-17.00\%$
Gemini Series Gemini-1.5-Flash $57.62$ $73.62$ $-30.15\%$ $-23.77\%$ $+10.20\%$ $-5.43\%$ $+6.00\%$ $-4.24\%$ $-30.95\%$ $-11.04\%$
Gemini-2.0-Flash $70.50$ $84.88$ $-41.31\%$ $-13.40\%$ $-0.77\%$ $-2.80\%$ $-11.47\%$ $-12.08\%$ $-21.51\%$ $-4.27\%$
Gemini-1.5-Pro-002 $70.87$ $81.88$ $-20.87\%$ $-14.96\%$ $-2.47\%$ $-7.02\%$ $-3.76\%$ $-9.62\%$ $-18.69\%$ $-7.63\%$
DeepSeek DeepSeek-V3 $69.08$ $82.88$ $-16.59\%$ $-7.99\%$ $+1.12\%$ $-0.30\%$ $+1.57\%$ $+0.90\%$ $-5.61\%$ $+3.17\%$
LLaMA Series LLaMA-3.1-8B-Instruct $37.12$ $41.62$ $-30.53\%$ $-61.86\%$ $+3.87\%$ $+1.50\%$ $+3.03\%$ $-3.60\%$ $-7.18\%$ $-24.92\%$
LLaMA-3.1-70B-Instruct $62.00$ $66.38$ $-28.29\%$ $-17.89\%$ $+4.91\%$ $+4.52\%$ $+1.61\%$ $+1.69\%$ $-17.41\%$ $-10.17\%$
LLaMA-3.1-405B-Instruct $66.79$ $75.00$ $-15.28\%$ $-9.00\%$ $+5.86\%$ $+3.67\%$ $+2.99\%$ $+2.33\%$ $-7.80\%$ $-5.67\%$
LLaMA-3.3-70B-Instruct $63.38$ $76.50$ $-29.39\%$ $-15.69\%$ $+1.25\%$ $-0.33\%$ $-4.08\%$ $+0.65\%$ $-21.83\%$ $-11.11\%$
Qwen Series Qwen2.5-7B-Instruct $38.83$ $51.38$ $-44.42\%$ $-50.36\%$ $+14.70\%$ $+3.16\%$ $+12.23\%$ $+7.54\%$ $-9.33\%$ $-14.36\%$
Qwen2.5-14B-Instruct $50.54$ $60.75$ $-38.42\%$ $-15.02\%$ $+8.66\%$ $+10.91\%$ $+4.86\%$ $+5.14\%$ $-8.74\%$ $-1.23\%$
Qwen2.5-32B-Instruct $63.50$ $74.12$ $-26.05\%$ $-20.40\%$ $+7.74\%$ $+0.51\%$ $+2.23\%$ $-4.05\%$ $-20.60\%$ $-10.96\%$
Qwen2.5-72B-Instruct $64.46$ $73.81$ $-26.86\%$ $-9.88\%$ $+2.65\%$ $+1.61\%$ $-3.68\%$ $+2.46\%$ $-25.53\%$ $-6.86\%$
Qwen2.5-Coder-32B-Instruct $74.25$ $78.25$ $-28.56\%$ $-20.61\%$ $-1.01\%$ 0.00 $-2.41\%$ $+0.32\%$ $-36.48\%$ $-33.39\%$
Average $\mathbf{-26.63}\%$ $\mathbf{-19.66}\%$ $\mathbf{+4.16}\%$ $\mathbf{+0.77}\%$ $\mathbf{+1.08}\%$ $\mathbf{-0.54}\%$ $\mathbf{-20.04}\%$ $\mathbf{-11.58}\%$
In this section, we analyze the impact of perturbations on input prediction tasks. Consistent with findings from the output prediction in §4.1, the ALL and MHC perturbations cause substantial performance degradation, with relative average drops of $26.63\%$ and $20.04\%$, respectively. These results align with Gu et al. (2024)’s hypothesis that robustness in related coding tasks is correlated. Furthermore, CoT prompting improves model performance by $6.97$pp under ALL and $8.64$pp under MHC, which is less pronounced than output prediction tasks due to the challenges and complexity of the inverse nature of the reasoning process.
Surprisingly, in contrast to output prediction, misleading message perturbations (MDC and MPS) do not degrade model performance in input prediction. Since input prediction allows multiple valid solutions, models can ignore certain branches of code logic, producing reasoning shortcuts. Our misleading messages are intentionally designed to confuse or distract models, but these comments inadvertently reinforce such shortcuts, simplifying models’ reasoning processes. As a result, models tend to produce trivial but valid inputs that bypass critical branches of the program without truly understanding and analyzing the underlying code semantics, which does not reflect genuine robustness in code comprehension. We provide a detailed case study in Appendix D.4 illustrating how models adopt these misleading messages to shortcut their reasoning.
Key Finding: ALL and MHC perturbations significantly degrade model performance, consistent with output prediction results, and reveal robustness challenges in code comprehension. Misleading NL messages accidentally reinforce shortcuts, encouraging models to produce trivial but valid answers without deep understanding.
5 Discussion
5.1 Comment Density and Robustness Degradation
Figure 2: Relative PASS@1 degradation under varying MDC injection probabilities $p$.
$p$ Qwen2.5-72B-Ins (CoT) QwQ-32B
$0.0$ $295$ $1391$
$0.1$ $288$ $1401$
$0.2$ $289$ $1399$
$0.3$ $284$ $1460$
$0.4$ $284$ $1469$
$0.5$ $276$ $1475$
$0.6$ $281$ $1556$
$0.7$ $278$ $1570$
$0.8$ $273$ $1598$
$0.9$ $276$ $1727$
$1.0$ $275$ $1814$
Table 6: Reasoning token usage under varying MDC injection probabilities $p$.
We design a textual density stress test to verify the validity of misleading comment perturbations and evaluate the resilience of different LLMs under varying densities of injected distractions on CRUX dataset. Specifically, we analyze how model performance changes as we gradually increase the comment injection probability $p\in\{0.0,0.1,\ldots,1.0\}$. We present the relative PASS@1 degradation of Qwen models across different perturbation densities in Figure 2. We further investigate the role of reasoning by evaluating CoT inference using Qwen2.5-72B-Ins and the internal reasoning model QwQ-32B, showcasing their robustness trends and reasoning token usages in Table 6.
Despite mild non-linearities caused by the randomness in injection positions and comment content, all models exhibit a clear performance degradation trend as the injection probability $p$ increases. Specifically, smaller models degrade more rapidly, while larger models demonstrate greater resilience under higher densities. This trend extends to CoT prompting and LRMs, with reasoning-enhanced models showing superior robustness against misleading textual distractions. These results validate MDC as a scalable stress-testing technique and an effective method for evaluating the robustness of LLMs under textual perturbations.
While reasoning-enhanced models have similar trends in the stress test, they react differently to dense comment noise. CoT prompting improves robustness by guiding models to selectively focus on the core logic path and ignore distracting signals. However, it can lead to reasoning shortcuts or premature conclusions without sufficiently analyzing relevant information, so token usage is slightly reduced as models do not analyze the comments but reduce the inference effort. In contrast, QwQ-32B processes all available information, including noisy or conflicting cues, and attempts to identify, reconcile, and resolve contradictions before making a final decision. This exhaustive reasoning enables higher vanilla performance and robustness compared to generic models. However, it also risks overthinking and collapse under contradiction-heavy settings, as we discussed in § 4.2.2.
5.2 Prompting to Ignore Misleading Comments
Table 7: Effect of Comment-Ignoring Prompt (PASS@1).
Model Inference VAN MDC MHC
w/o Ignore w/ Ignore w/o Ignore w/ Ignore
DeepSeek-V3 Direct $67.78$ $39.74$ $41.75$ $47.67$ $47.81$
CoT $92.07$ $70.15$ $82.05$ $89.98$ $94.57$
GPT-4o Direct $64.51$ $45.72$ $45.93$ $47.32$ $46.76$
CoT $94.36$ $87.47$ $89.35$ $94.78$ $94.57$
Qwen2.5-72B-Instruct Direct $54.91$ $47.95$ $48.02$ $32.64$ $32.50$
CoT $90.40$ $85.39$ $86.01$ $82.67$ $84.34$
We conduct a controlled experiment by adding the instruction: “Please ignore the comments and any other non-code elements in the code snippet.” to the prompt on three strong models under direct and CoT inference. From Table 7, even when explicitly instructed to ignore the comments, the PASS@1 scores under direct inference remain similar to those without the instruction, indicating that LLMs still incorporate NL cues into their inference, which is consistent with our finding on their inherent reliance on embedded NL messages in § 4.1.2. Although CoT prompting provides improvements under comment-ignore instructions, the PASS@1 performance of MDC across all models is still far away from the vanilla performance, indicating that even with step-by-step reasoning, LLMs are unable to fully detach from misleading NL distractions, highlighting a fundamental limitation in their reasoning controllability and dependence on surface cues.
5.3 Revealing the Vulnerability of LRMs
Figure 3: Cumulative confusion word trend in QwQ’s output (LCB sample 259 under MHC).
In §4.2.2, we highlighted a critical vulnerability in LRMs. We observed an unhealthy self-reflection loop and a progressive loss of confidence by analyzing a failure case where QwQ-32B generated over 32,000 tokens in response to an MHC-perturbed code.
We manually reviewed and segmented the reasoning process into six stages (see Figure 4 and Appendix D.5). Initially, the model reasons correctly and computes the correct output. However, our plausible yet incorrect hint triggers it to doubt its logic and repeatedly re-execute the same reasoning steps. This recursive self-verification gradually becomes brute-force enumeration, causing the model to hallucinate explanations. Ultimately, the reasoning process collapses, and the model repeatedly outputs over 12,000 “Hmm.”
To further investigate the model’s confidence before collapsing, we track the cumulative frequency of confusing words, such as “hmm”, “wait”, “but”, “perhaps” and “mistake”, and visualize the trend in Figure 3. The trend exhibits a striking quadratic growth pattern ($R^{2}=0.9991$), indicating compounding confusion as reasoning progresses. The quadratic trajectory supports our stage-wise analysis: the model initially shows stable reasoning, but once the misleading hint is triggered, confusion accumulates, culminating in complete failure and collapse.
6 Conclusion
We introduce a direction to evaluate the robustness of LLMs in code comprehension and execution under four structural and three textual perturbations. Our evaluation of seventeen LLMs under direct and CoT inference reveals critical vulnerabilities in model behavior, a heavy reliance on NL cues, and a limited capacity to follow actual program logic. While larger and newer models and CoT prompting significantly improve robustness against perturbations, they still cannot fully escape this surface-level dependence and leverage these cues to form reasoning shortcuts. Despite three LRMs demonstrating superior robustness under all perturbations—highlighting the strength of internal reasoning—we expose a new risk of overthinking and collapse when plausible but incorrect cues mislead their self-reflection mechanisms. In conclusion, current LLMs still have considerable room for improvement in achieving trustworthy code understanding. We believe that our benchmark serves as a practical framework for studying LLMs’ code inference behavior, providing significant insights into their robustness, reasoning tendencies, and potential failures.
Appendix A Prompts
Our prompt instructions for LLMs follow the format designed by CRUX, which is also adopted by LCB, utilizing a two-shot prompt for direct inference and a single-shot prompt for CoT inference to ensure models can follow the instruction and output format effectively. For LRMs, internal reasoning is inherently required before responding. Therefore, applying CoT prompts may lead to redundant reasoning and unfair comparison with standard LLMs under CoT prompting. Considering the fairness, we use direct inference prompts for RM experiments in §4.2.2.
Unlike LCB, we rename all example function names to f to avoid prompting models to capture the NL cues from the example. We add the instruction “Do NOT output any extra information” to prevent external output, particularly from Gemini and Claude models, to ensure all models have a fair output environment. We also include “Ensure the provided expression syntax is correct” to enforce consistent syntax. However, some responses from models like QwQ-32B, DeepSeek-R1 still generate syntactically incorrect outputs such as [ANSWER] f("a") == [/ANSWER] instead of [ANSWER] f("a") == "" [/ANSWER]. For those situations, we identify them as “Syntax Errors” and incorrect.
We provide the complete prompts used for output prediction in both direct and CoT inference (Code 1 and 2) and for input prediction in both direct and CoT inference (Code 3 and 4).
A.1 Output Prediction (Direct)
Listing 1: Output Prediction Prompt with Direct Inference.
Given the code snippet:
‘‘‘python
{code}
and the function call with input arguments:
‘‘‘python
{input}
Predict the exact output value for ‘{input}‘ and output your prediction using the special tokens [ANSWER] {input} == ?? [/ANSWER]. Do NOT output any extra information.
Ensure the provided expression syntax is correct!
Example 1:
Given the code snippet:
‘‘‘python
def f(n):
return n
and the function call with input arguments:
‘‘‘python
f(17)
The output value for f(17)‘ is 17, then output your prediction [ANSWER] f(17) == 17 [/ANSWER].
Example 2:
Given the code snippet:
‘‘‘python
def f(s):
return s + "a"
and the function call with input arguments:
‘‘‘python
f("x9j")
The output value for f("x9j")‘ is "x9ja", then output your prediction [ANSWER] f("x9j") == "x9ja" [/ANSWER].
A.2 Output Prediction (CoT)
Listing 2: Output Prediction Prompt with CoT-based Inference.
Given the code snippet:
‘‘‘python
{code}
and the function call with input arguments:
‘‘‘python
{input}
Predict the exact output value for ‘{input}‘, execute the program step by step before arriving at an answer within the tokens [THOUGHT] and [/THOUGHT], and output your prediction using the special tokens [ANSWER] {input} == ?? [/ANSWER]. Do NOT output any extra information.
Ensure the provided expression syntax is correct!
For example:
Given the code snippet:
‘‘‘python
def f(s):
s = s + s
return "b" + s + "a"
and the input arguments:
‘‘‘python
f("hi")
[THOUGHT]
Lets execute the code step by step:
1. The function f is defined, which takes a single argument s.
2. The function is called with the argument "hi", so within the function, s is initially "hi".
3. Inside the function, s is concatenated with itself, so s becomes "hihi".
4. The function then returns a new string that starts with "b", followed by the value of s (which is now "hihi"), and ends with "a".
5. The return value of the function is therefore "bhihia".
[/THOUGHT]
Thus, the output value for f("hi")‘ is "bhihia", then output your prediction [ANSWER] f("hi") == "bhihia" [/ANSWER].
A.3 Input Prediction (Direct)
Listing 3: Input Prediction Prompt with Direct Inference.
Given the code snippet:
‘‘‘python
{code}
and output value:
‘‘‘python
{output}
Please predict the input arguments for the function ‘{function_name}‘ that result in the output value ‘{output}‘ and output your prediction using the special tokens [ANSWER] {function_name}(??) == {output} [/ANSWER]. Do NOT output any extra information.
There may be multiple answers, but you should only output one. Ensure the provided expression syntax is correct!
Example 1:
Given the code snippet:
‘‘‘python
def f(my_list):
count = 0
for i in my_list:
if len(i) % 2 == 0:
count += 1
return count
and output value:
‘‘‘python
3
The input arguments for f that result in the output value of ‘3‘ are ‘["mq", "px", "zy"]‘. Then, output your prediction [ANSWER] f(["mq", "px", "zy"]) == 3 [/ANSWER].
Example 2:
Given the code snippet:
‘‘‘python
def f(s1, s2):
return s1 + s2
and output value:
‘‘‘python
"banana"
The input arguments for f that result in the output value of ‘"banana"‘ are ‘"ba", "nana"‘. Then, output your prediction [ANSWER] f("ba", "nana") == "banana" [/ANSWER].
A.4 Input Prediction (CoT)
Listing 4: Input Prediction Prompt with CoT-based Inference.
Given the code snippet:
‘‘‘python
{code}
and output value:
‘‘‘python
{output}
Please predict the input arguments for the function ‘{function_name}‘ that result in the output value ‘{output}‘ step by step before arriving at an answer within the tokens [THOUGHT] and [/THOUGHT], and output your prediction using the special tokens [ANSWER] {function_name}(??) == {output} [/ANSWER]. Do NOT output any extra information.
There may be multiple answers, but you should only output one. Ensure the provided expression syntax is correct!
For example:
Given the code snippet:
‘‘‘python
def f(x):
return x + 1
and the output value:
‘‘‘python
17
[THOUGHT]
To find an input such that executing f on the input leads to the given output, we can work backwards from the given assertion. We know that f(??) == 17.
Since the function f(x) returns x + 1, for f(??) to be equal to 17, the value of ?? should be 16.
[/THOUGHT]
Thus, the input arguments for the function f that result in the output value ‘17‘ is ‘16‘. Then, output your prediction [ANSWER] f(16) == 17 [/ANSWER].
Appendix B Illustrative Examples of Perturbation Strategies
In section 3.3, we briefly describe all perturbation strategies employed in this paper. In this appendix section, we illustrate the detailed implementation for all perturbations with a LCB sample 36 as the example. Code 5 presents the base vanilla example where the formal answer is minimumCost(s = ’0011’) == 2.
Listing 5: Vanilla example from LCB (sample 36) for output prediction.
def minimumCost(s: str) -> int:
ans = 0
for i in range(1, len(s)):
if s[i - 1] != s[i]:
ans += min(i, len(s) - i)
return ans
B.1 Renaming Entities
We use AST to identify all variable names appearing in:
  • assignment statements (e.g., a, b = 1, 2),
  • for-loops (e.g., for i in range(10)),
  • with-statements (e.g., with open(’file’) as f),
  • walrus operators (e.g., if (a := 10) > 5), and
  • comprehensions (e.g., [x for x in range(10)]).
[…]
Listing 6: REN-perturbed example from LCB (sample 36) for output prediction.
def f(Var_1: str) -> int:
Var_2 = 0
for Var_3 in range(1, len(Var_1)):
if Var_1[Var_3 - 1] != Var_1[Var_3]:
Var_2 += min(Var_3, len(Var_1) - Var_3)
return Var_2
B.2 Reformatting Conditional Expression
We generally categorize conditional expressions into three types: (1) Comparison-based conditions (e.g., if a < b:), (2) Boolean constant conditions (e.g., if True:), and (3) General conditions (e.g., if x:). For both comparison-based and boolean constant conditions, we design six transformation templates:
  1. 1.
    _ := ({cond},)[0],
  2. 2.
    (lambda: {cond})(),
  3. 3.
    ({cond}) == ({para} == {para}),
  4. 4.
    ({cond}) != ({para} != {para}),
  5. 5.
    bool(-~({cond})) == ({cond}),
  6. 6.
    bool(int({cond})) (comparison-based only), and
  7. 7.
    eval(str({cond})) (boolean constant only).
For general conditions, we also design four transformation templates:
  1. 1.
    not not ({cond}),
  2. 2.
    ({cond}) or False,
  3. 3.
    ({cond}) and True, and
  4. 4.
    _ := ({cond},)[0].
Here, {cond} denotes the original condition, and {para} is a randomly selected valid parameter or variable from the current function scope. Code 7 presents an example of this perturbation.
Listing 7: RTF-perturbed example from LCB (sample 36) for output prediction.
def minimumCost(s: str) -> int:
ans = 0
for i in range(1, len(s)):
if _ := (s[i - 1] != s[i],)[0]:
ans += min(i, len(s) - i)
return ans
B.3 Inserting Garbage Code
As introduced in §3.3.1, we inject three types of garbage code: (1) repeated-name global variable declarations, (2) in-executable conditional statements, and (3) dead-loop function definitions. For (1), we randomly select a few parameters or variables and declare global variables using the same names, assigning arbitrary integers at the beginning of the program. For (2), we design seven templates for false conditions:
  1. 1.
    False,
  2. 2.
    None,
  3. 3.
    0,
  4. 4.
  5. 5.
    {para} != {para},
  6. 6.
    not {para} == {para}, and
  7. 7.
    print({para}).
We then insert these false conditions into four control-flow templates by {false_cond}:
  1. 1.
    if {false_cond}: {new_var} = {para},
  2. 2.
    while {false_cond}: {new_var} = {para},
  3. 3.
    for i in range(0): {new_var} = {para}, and
  4. 4.
    {new_var} = {para} if {false_cond} else {para}.
[…]
Listing 8: Garbage death loop function templates for GBC perturbation.
# Death Loop Template 1
def funct1():
funct2()
def funct2():
funct1()
# Death Loop Template 2
def funct3():
def funct4():
funct3()
funct4()
# Death Loop Template 3
def funct5():
i = 1
while True:
i+=1
# Death Loop Template 4
def funct6():
for i in iter(int, 1):
i+=1
# Death Loop Template 5
def funct7():
try:
funct7()
except:
funct7()
# Death Loop Template 6
def funct8():
items = [0]
for x in items:
items.append(x + 1)
# Death Loop Template 7
def funct9():
for _ in iter(lambda: True, False):
pass
Listing 9: GBC-perturbed example from LCB (sample 36) for output prediction.
s = 59
def minimumCost(s: str) -> int:
while False:
TempVar1 = s
def funct1():
funct2()
def funct8():
items = [0]
for x in items:
items.append(x + 1)
def funct2():
funct1()
ans = 0
for i in range(1, len(s)):
if s[i - 1] != s[i]:
ans += min(i, len(s) - i)
elif print(s):
TempVar0 = s
TempVar2 = s if not s == s else s
return ans
B.4 Aggregating Structural Perturbations
As introduced in §3.3.1, we aggregate the application of REN, RTF, and GBC to a single code snippet, as shown in Code 10.
Listing 10: ALL-perturbed example from LCB (sample 36) for output prediction.
Var_1 = 24
def f2(Var_1: str) -> int:
def f1():
for Var_4 in iter(lambda : True, False):
pass
def funct6():
for Var_3 in iter(int, 1):
Var_3 += 1
Var_6 = 0
for Var_3 in range(1, len(Var_1)):
Var_7 = Var_1 if 0 else Var_1
if None:
Var_5 = Var_1
if (lambda : Var_1[Var_3 - 1] != Var_1[Var_3])():
Var_6 += min(Var_3, len(Var_1) - Var_3)
while Var_1 != Var_1:
Var_2 = Var_1
return Var_6
B.5 Inserting Misleading Messages
As introduced in §3.3.2, we insert misleading comments for key AST node types, including function declarations, return statements, loops, conditionals, variable assignments, and common data structure operations. The data structure operations include common list, set, and string manipulations: append, extend, insert, remove, pop, sort, reverse, update, add, split, join, replace, lower, upper, capitalize, and swap case. To construct these comments, we design prompts for GPT-4o to generate misleading but plausible messages for each type. In this experiment, we manually selected 20 messages for each AST node type and 5 for each data structure operation. Take variable assignments as an example; Code 11 is the prompt for generating misleading comments for variable assignments, and Code 12 is a list of selected misleading comment candidates. Code 13 and 14 present an example of MDC and MPS perturbations.
Listing 11: Prompt used to generate misleading comments for variable assignments.
Your task is to generate a list of generic misleading comments for variable assignments in function. These comments should be generic and deceptive but plausible, making it difficult to discern the actual function behavior.
Guidelines:
1. The comments should be neutral, direct, and generic, applicable to any variable regardless of its type, usage, or scope, without being function-specific.
2. Ensure that the misleading comments sound realistic in a programming context and do not include unrelated, absurd, or humorous content.
3. The comments should mislead by misrepresenting the role of the variable, suggesting that it is useless, does not affect the functions output, or serves no real purpose in computation.
Example:
‘‘‘python
a = 1 # <-- misleading comment
A good misleading comment example would be: "The {variable} variable is initialized but never populated."
Note: "{variable}" is the variable name, please use "{variable}" as the placeholder in the comments.
[…]
Listing 12: Selected misleading comments for variable assignments used in MDC and MPS perturbations, where ‘{variable}’ is the variable name.
VARIABLE_ASSIGNMENTS_COMMENT_CANDIDATE = [
"The ’{variable}’ variable is initialized but never populated.",
"The ’{variable}’ variable is declared for debugging purposes only.",
"The ’{variable}’ variable is assigned a value but never referenced again.",
"The ’{variable}’ variable is a placeholder for future functionality.",
"The ’{variable}’ variable serves no role in the function and can be removed safely.",
"The ’{variable}’ variable is not involved in any meaningful computation.",
"’{variable}’ serves as a dummy variable, holding no real significance.",
"The ’{variable}’ variable holds a default value and is not used elsewhere.",
"’{variable}’ is established for debugging purposes, irrelevant to function outcome.",
"Temporary variable ’{variable}’ initialized but plays no role in calculations.",
"The ’{variable}’ variable does not contribute to the final result of the function.",
"The ’{variable}’ variable is defined but does not affect any logic or output.",
"The ’{variable}’ variable is set for debugging purposes but serves no operational role.",
"The ’{variable}’ variable is not used in the function and can be removed.",
"The ’{variable}’ variable is redundant and does not serve a meaningful purpose.",
"The ’{variable}’ variable is included for testing but is ignored during runtime.",
"The ’{variable}’ variable is a temporary storage that remains unused.",
"The ’{variable}’ variable is assigned a default value but is not utilized.",
"The ’{variable}’ variable holds no significance and is effectively inert.",
"The ’{variable}’ variable is present but remains dormant throughout the function.",
Listing 13: MDC-perturbed example from LCB (sample 36) for output prediction.
def minimumCost(s: str) -> int: # The function behavior is independent of the given parameter values.
ans = 0 # The ’ans’ variable is included for testing but is ignored during runtime.
# The loop’s purpose is for clarity, not computation.
for i in range(1, len(s)):
if s[i - 1] != s[i]: # This condition serves as a placeholder and will be removed in future versions.
ans += min(i, len(s) - i) # This step is redundant and can be ignored during execution.
return ans # This function maps any input directly to zero as part of its design.
Listing 14: MPS-perturbed example from LCB (sample 36) for output prediction.
def minimumCost(s: str) -> int:
print(’The parameters determine only the speed, not the outcome, of the function.’)
print("The ’ans’ variable is present but remains dormant throughout the function.")
ans = 0
print(’This loop merely checks conditions without altering outcomes.’)
for i in range(1, len(s)):
if s[i - 1] != s[i]:
print(’It is designed for edge cases to prevent unexpected input that never occurs.’)
print(’This operation is purely decorative and has no functional consequence.’)
ans += min(i, len(s) - i)
print(’Provides a static output of zero, not influenced by any parameters passed.’)
return ans
B.6 Inserting Incorrect Hints
In order to generate plausible, logical but incorrect hints, we instruct GPT-4o to rewrite the correct assertion expression to incorrect ones and only allow it to modify the output value using the prompt in Code 15 for output prediction. An example perturbed instance is shown in Code 16, where the ground-truth return value 2 is replaced with the misleading hint # The return value is 3.
In contrast, we use the prompt shown in Code 17 to generate misleading input hints, allowing GPT-4o to perform CoT reasoning before generating the expression. Since some functions return the same output regardless of the input arguments, we instruct GPT-4o to force the expression to become incorrect by modifying input arguments, such as changing data types, swapping parameter order, or adding/removing arguments. An example perturbed instance is shown in Code 18.
Listing 15: Prompt for generating misleading hint comments for output prediction.
Given the code snippet:
‘‘‘python
{code}
and the correct expression for the function call:
‘‘‘python
{expression}
Modify the output value to make it INCORRECT. The modification should introduce moderate changes, ensuring diversity and avoiding minimal adjustments. For example, if the output is a list, you can add new elements, remove elements, or modify the values of existing elements. However, the modification should still align logically with the code.
The purpose is to misleading people for getting correct answer.
Do NOT modify the function call and the input arguments!
Output the incorrect expression using the special tokens as follows: [EXPRESSION] assert <expression> [/EXPRESSION].
Example 1:
Given the function:
‘‘‘python
def f(n):
return n
and the correct expression
‘‘‘python
assert f(17) == 17
Modify the expression such that it fails for the execution.
You can modify the either the input arguments, or the output value, even both of them [EXPRESSION] assert f(10) == 20 [/EXPRESSION].
Example 2:
Given the function:
‘‘‘python
def f(s):
return s + "a"
and the correct expression
‘‘‘python
assert f("x9j") == "x9ja"
Modify the expression such that it fails for the execution and output [EXPRESSION] assert f("x9j") == "x9j" [/EXPRESSION].
Listing 16: Prompt for generating misleading hint comments for input prediction.
def minimumCost(s: str) -> int:
ans = 0
for i in range(1, len(s)):
if s[i - 1] != s[i]:
ans += min(i, len(s) - i)
return ans # The return value is 3
Listing 17: Prompt for generating misleading hint comments for input prediction.
Given the code snippet:
‘‘‘{programming_language}
{code}
and the correct expression for the function call:
‘‘‘{programming_language}
{expression}
Modify the input argument(s) to make it INCORRECT. The purpose is to mislead people to get a correct answer. Do NOT modify the output value!
Please think about how to modify the input arguments to make the expression incorrect step by step before arriving at an answer within the tokens [THOUGHT] and [/THOUGHT]
Output the incorrect expression using the special tokens as follows: [EXPRESSION] assert <expression> [/EXPRESSION].
Remember, the modification should introduce moderate changes, ensuring diversity and avoiding minimal adjustments.
However, if the function always returns the same value regardless of input, force an incorrect expression by modifying the arguments in a way that ensures failure (e.g., change an inputs type, swap their order, or add/remove an argument).
Example 1:
Given the function:
‘‘‘{programming_language}
def f(n):
return n
and the correct expression:
‘‘‘{programming_language}
assert f(17) == 17
[THOUGHT]
To find an input such that executing f on the input leads to the given output, we can work backwards from the given assertion. We know that f(??) == 17.
Since the function f(n) returns n, for f(??) to be equal to 17, the value of ?? should be 17.
Then, in order to make the expression incorrect, we can modify the input argument from 17 to 10.
[/THOUGHT]
[EXPRESSION] assert f(10) == 17 [/EXPRESSION].
Listing 18: MHC-perturbed example from CRUX (sample 2) for input prediction.
def f(text): # The function call is f(’xbtofdeiequ’)
new_text = list(text)
for i in ’+’:
if i in new_text:
new_text.remove(i)
return ’’.join(new_text)
Appendix C Supplementary Experimental Results
C.1 Impact on Output Prediction
Table 8: Relative PASS@1 degradation of LLMs under all perturbation types for output prediction tasks across CRUX and LCB in direct inference (N=3) [TO ABSOLUTE SCORE].
Model Series Model Name VAN REN RTF GBC ALL MDC MPS MHC
Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
GPT-4 Omni GPT-4o-Mini $57.29$ $53.24$ $-1.09\%$ $-18.17\%$ $-3.13\%$ $-1.83\%$ $-19.05\%$ $-10.33\%$ $-27.85\%$ $-31.90\%$ $-32.58\%$ $-40.39\%$ $-23.20\%$ $-27.97\%$ $-11.56\%$ $-44.31\%$
GPT-4o $71.33$ $64.51$ $-3.80\%$ $-14.89\%$ $-4.73\%$ $-7.12\%$ $-12.91\%$ $-4.96\%$ $-15.01\%$ $-28.37\%$ $-13.96\%$ $-29.13\%$ $-17.23\%$ $-24.27\%$ $-6.37\%$ $-26.65\%$
Claude Series Claude-3.5-Haiku-20241022 $57.42$ $58.25$ $-6.89\%$ $-16.01\%$ $-3.12\%$ $+0.60\%$ $-12.70\%$ $-7.77\%$ $-24.02\%$ $-36.92\%$ $-13.79\%$ $-10.04\%$ $-11.25\%$ $-8.00\%$ $-27.50\%$ $-53.29\%$
Claude-3.5-Sonnet-20241022 $71.46$ $73.83$ $-2.33\%$ $-22.71\%$ $-1.28\%$ $+1.60\%$ $-9.62\%$ $-6.69\%$ $-14.75\%$ $-34.12\%$ $-8.10\%$ $-9.61\%$ $-8.63\%$ $-10.65\%$ $-14.40\%$ $-43.45\%$
Gemini Series Gemini-1.5-Flash $56.21$ $44.26$ $-6.30\%$ $-0.94\%$ $-2.08\%$ $-2.36\%$ $-8.15\%$ $-0.79\%$ $-18.38\%$ $-12.74\%$ $-14.75\%$ $-32.23\%$ $-28.47\%$ $-40.41\%$ $-7.34\%$ $-22.17\%$
Gemini-2.0-Flash $65.04$ $65.97$ $-3.59\%$ $-27.43\%$ $-3.20\%$ $-1.16\%$ $-23.83\%$ $-13.29\%$ $-33.25\%$ $-42.09\%$ $-26.78\%$ $-51.90\%$ $-33.12\%$ $-50.32\%$ $-8.26\%$ $-19.51\%$
Gemini-1.5-Pro-002 $67.37$ $56.16$ $-4.33\%$ $-6.07\%$ $-3.71\%$ $-8.43\%$ $-8.35\%$ $-6.82\%$ $-19.91\%$ $-23.17\%$ $-23.01\%$ $-32.96\%$ $-21.71\%$ $-32.84\%$ $-10.02\%$ $-32.47\%$
DeepSeek DeepSeek-V3 $67.87$ $67.78$ $-3.93\%$ $-24.13\%$ $-1.96\%$ $-5.54\%$ $-6.81\%$ $-8.83\%$ $-12.89\%$ $-35.63\%$ $-16.57\%$ $-41.38\%$ $-10.13\%$ $-34.19\%$ $-10.74\%$ $-29.67\%$
LLaMA Series LLaMA-3.1-8B-Instruct $35.96$ $34.66$ $+0.23\%$ $-12.45\%$ $-0.93\%$ $-4.82\%$ $-8.69\%$ $-13.65\%$ $-23.75\%$ $-19.68\%$ $-22.48\%$ $-28.51\%$ $-17.84\%$ $-23.09\%$ $-16.69\%$ $-39.16\%$
LLaMA-3.1-70B-Instruct $56.08$ $44.89$ $-4.23\%$ $-3.10\%$ $-0.97\%$ $-3.88\%$ $-9.58\%$ $-5.43\%$ $-19.61\%$ $-17.21\%$ $-19.24\%$ $-27.91\%$ $-26.67\%$ $-38.76\%$ $-8.40\%$ $-32.25\%$
LLaMA-3.1-405B-Instruct $63.54$ $50.66$ $-8.46\%$ $-7.69\%$ $-0.98\%$ $-8.79\%$ $-9.11\%$ $-5.63\%$ $-19.67\%$ $-16.90\%$ $-6.56\%$ $-14.70\%$ $-11.28\%$ $-19.92\%$ $-6.89\%$ $-25.69\%$
LLaMA-3.3-70B-Instruct $59.92$ $48.50$ $-3.69\%$ $-2.87\%$ $-2.09\%$ $-7.03\%$ $-9.18\%$ $-11.76\%$ $-17.32\%$ $-20.09\%$ $-12.45\%$ $-19.80\%$ $-13.49\%$ $-25.39\%$ $-6.82\%$ $-20.95\%$
Qwen Series Qwen2.5-7B-Instruct $43.29$ $41.41$ $-7.03\%$ $-17.48\%$ $-2.98\%$ $-7.73\%$ $-29.93\%$ $-9.08\%$ $-37.92\%$ $-30.92\%$ $-58.04\%$ $-38.32\%$ $-45.24\%$ $-19.83\%$ $-26.95\%$ $-55.63\%$
Qwen2.5-14B-Instruct $47.79$ $49.48$ $-3.84\%$ $-19.97\%$ $-1.39\%$ $-2.81\%$ $-23.45\%$ $-6.89\%$ $-39.84\%$ $-29.96\%$ $-34.87\%$ $-41.35\%$ $-32.52\%$ $-33.90\%$ $-9.33\%$ $-22.22\%$
Qwen2.5-32B-Instruct $60.04$ $59.64$ $-2.85\%$ $-18.09\%$ $-4.30\%$ $-0.82\%$ $-9.65\%$ $-2.33\%$ $-19.78\%$ $-34.89\%$ $-18.04\%$ $-32.67\%$ $-23.66\%$ $-23.57\%$ $-13.12\%$ $-30.81\%$
Qwen2.5-72B-Instruct $60.08$ $54.91$ $-1.73\%$ $-15.46\%$ $-3.12\%$ $+0.76\%$ $-17.82\%$ $-3.68\%$ $-23.30\%$ $-25.22\%$ $-16.99\%$ $-12.67\%$ $-24.06\%$ $-26.24\%$ $-16.02\%$ $-40.56\%$
Qwen2.5-Coder-32B-Instruct $66.96$ $56.65$ $-3.73\%$ $-14.86\%$ $-1.43\%$ $-3.56\%$ $-16.05\%$ $-4.91\%$ $-22.34\%$ $-27.89\%$ $-23.21\%$ $-36.61\%$ $-16.55\%$ $-30.34\%$ $-10.45\%$ $-21.74\%$
Average $-3.98\%$ $-14.25\%$ $-2.44\%$ $-3.70\%$ $-13.82\%$ $-7.23\%$ $-22.92\%$ $-27.51\%$ $-21.26\%$ $-29.42\%$ $-21.47\%$ $-27.63\%$ $-12.40\%$ $-32.97\%$
$\mathbf{-7.83}\%$ $\mathbf{-2.91}\%$ $\mathbf{-11.35}\%$ $\mathbf{-24.64}\%$ $\mathbf{-24.32}\%$ $\mathbf{-23.78}\%$ $\mathbf{-20.11}\%$
C.2 Robustness Across Model Scales
In §4.1.3, we compare the robustness against different model scales. To fairly aggregate model robustness across different perturbation types, we compute a weighted average of the relative PASS@1 degradation under four representative perturbations: ALL (structural), MDC and MPS (misleading messages), and MHC (hints). We select ALL as the representative structural perturbation because of the compounding effects of REN, RTF, and GBC captured in §4.1.1, which reflect the most challenging structural setting. Since MDC and MPS stem from the same perturbation strategy and often target overlapping AST node types, we assign them lower weights ($1/8$ each) to avoid overrepresenting similar effects. Overall, we define a weighted robustness metric with weights of $1/2$, $1/8$, $1/8$, and $1/4$ for ALL, MDC, MPS, and MHC, respectively, and summarize the results in Table 2. For completeness, we also report the results under uniform weighting ($1/4$ for each) in Table 9, which show similar trends and confirm the stability of our findings in §4.1.3.
Table 9: Perturbations Output Direct (Relative)
Model Series Model Name VAN ALL ($1/4$) MDC ($1/4$) MPS ($1/4$) MHC ($1/4$) Average
GPT-4 Omni GPT-4o-Mini 55.77 $-29.37\%$ $-35.51\%$ $-24.99\%$ $-23.83\%$ $-28.42\%$
GPT-4o 68.78 $-20.01\%$ $-19.64\%$ $-19.87\%$ $-13.96\%$ $-18.37\%$
Claude Series Claude-3.5-Haiku-20241022 57.73 $-28.85\%$ $-12.38\%$ $-10.03\%$ $-37.16\%$ $-22.11\%$
Claude-3.5-Sonnet-20241022 72.35 $-22.01\%$ $-8.67\%$ $-9.39\%$ $-25.28\%$ $-16.34\%$
Gemini Series Gemini-1.5-Flash 51.73 $-16.27\%$ $-21.30\%$ $-32.94\%$ $-12.89\%$ $-20.85\%$
Gemini-2.0-Flash 65.39 $-36.56\%$ $-36.19\%$ $-39.56\%$ $-12.48\%$ $-31.20\%$
Gemini-1.5-Pro-002 63.17 $-21.13\%$ $-26.73\%$ $-25.88\%$ $-18.43\%$ $-23.04\%$
DeepSeek DeepSeek-V3 67.84 $-21.41\%$ $-25.86\%$ $-19.14\%$ $-17.83\%$ $-21.06\%$
LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-22.23\%$ $-24.74\%$ $-19.81\%$ $-25.10\%$ $-22.97\%$
LLaMA-3.1-70B-Instruct 51.89 $-18.71\%$ $-22.49\%$ $-31.20\%$ $-17.33\%$ $-22.43\%$
LLaMA-3.1-405B-Instruct 58.72 $-18.63\%$ $-9.61\%$ $-14.51\%$ $-13.93\%$ $-14.17\%$
LLaMA-3.3-70B-Instruct 55.64 $-18.35\%$ $-15.20\%$ $-17.95\%$ $-12.11\%$ $-15.90\%$
Qwen Series Qwen2.5-7B-Instruct 42.59 $-35.30\%$ $-50.65\%$ $-35.72\%$ $-37.69\%$ $-39.84\%$
Qwen2.5-14B-Instruct 48.42 $-36.14\%$ $-37.30\%$ $-33.04\%$ $-14.16\%$ $-30.16\%$
Qwen2.5-32B-Instruct 59.89 $-25.44\%$ $-23.52\%$ $-23.63\%$ $-19.74\%$ $-23.08\%$
Qwen2.5-72B-Instruct 58.14 $-24.02\%$ $-15.37\%$ $-24.88\%$ $-25.21\%$ $-22.37\%$
Qwen2.5-Coder-32B-Instruct 63.10 $-24.42\%$ $-28.23\%$ $-21.72\%$ $-14.68\%$ $-22.26\%$
C.3 Effect of CoT Reasoning
Table 10: Comparison of absolute PASS@1 degradation aggregated over CRUX and LCB for output prediction tasks in direct inference (N=3) and CoT inference (N=1) [TO RELATIVE SCORE].
Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average
Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT
GPT-4 Omni GPT-4o-Mini $55.77$ $81.47$ $-16.34$ $-10.71$ $-19.73$ $-8.84$ $-13.89$ $-8.91$ $-12.98$ $-2.19$ $-15.62$ $-7.75$
GPT-4o $68.78$ $91.79$ $-13.55$ $-4.53$ $-13.27$ $-4.61$ $-13.55$ $-5.16$ $-9.28$ $-1.25$ $-12.45$ $-3.87$
Claude Series Claude-3.5-Haiku-20241022 $57.73$ $72.95$ $-16.68$ $-15.56$ $-7.14$ $-7.74$ $-5.79$ $-6.33$ $-21.50$ $-10.40$ $-15.33$ $-10.43$
Claude-3.5-Sonnet-20241022 $72.35$ $86.00$ $-16.03$ $-6.49$ $-6.28$ $-4.53$ $-6.80$ $-5.47$ $-18.45$ $-6.80$ $-14.26$ $-5.90$
Gemini Series Gemini-1.5-Flash $51.73$ $75.22$ $-8.57$ $-13.76$ $-10.53$ $-16.34$ $-16.71$ $-31.74$ $-6.25$ $-1.56$ $-9.26$ $-15.34$
Gemini-2.0-Flash $65.39$ $89.13$ $-23.92$ $-5.55$ $-23.72$ $-5.63$ $-25.91$ $-12.74$ $-8.18$ $-1.80$ $-20.21$ $-6.25$
Gemini-1.5-Pro-002 $63.17$ $87.18$ $-13.27$ $-6.49$ $-16.63$ $-10.40$ $-16.05$ $-12.82$ $-11.05$ $-3.36$ $-13.48$ $-8.01$
DeepSeek DeepSeek-V3 $67.84$ $89.52$ $-14.52$ $-8.91$ $-17.54$ $-15.87$ $-12.98$ $-16.81$ $-12.09$ $-3.13$ $-14.10$ $-10.81$
LLaMA Series LLaMA-3.1-8B-Instruct $35.47$ $44.72$ $-7.90$ $-12.35$ $-8.76$ $-9.38$ $-7.01$ $-9.62$ $-8.84$ $-3.91$ $-8.13$ $-8.96$
LLaMA-3.1-70B-Instruct $51.89$ $69.43$ $-9.77$ $-12.43$ $-11.44$ $-16.11$ $-15.87$ $-21.42$ $-8.37$ $-4.69$ $-10.39$ $-13.35$
LLaMA-3.1-405B-Instruct $58.72$ $78.42$ $-11.02$ $-10.63$ $-5.39$ $-8.05$ $-8.26$ $-12.20$ $-7.61$ $-5.86$ $-9.12$ $-9.21$
LLaMA-3.3-70B-Instruct $55.64$ $76.86$ $-10.14$ $-8.37$ $-8.26$ $-5.94$ $-9.67$ $-8.52$ $-6.36$ $-3.67$ $-8.90$ $-6.68$
Qwen Series Qwen2.5-7B-Instruct $42.59$ $58.17$ $-15.06$ $-12.67$ $-21.66$ $-12.74$ $-15.32$ $-16.26$ $-15.92$ $-4.07$ $-16.14$ $-11.34$
Qwen2.5-14B-Instruct $48.42$ $70.68$ $-17.46$ $-16.34$ $-18.09$ $-24.94$ $-16.00$ $-30.65$ $-6.91$ $-7.66$ $-14.72$ $-19.33$
Qwen2.5-32B-Instruct $59.89$ $79.44$ $-15.22$ $-13.84$ $-14.07$ $-14.78$ $-14.15$ $-19.23$ $-11.81$ $-4.22$ $-14.09$ $-12.86$
Qwen2.5-72B-Instruct $58.14$ $82.80$ $-13.94$ $-10.48$ $-8.99$ $-5.63$ $-14.44$ $-13.45$ $-14.36$ $-5.47$ $-13.49$ $-8.80$
Qwen2.5-Coder-32B-Instruct $63.10$ $84.75$ $-15.27$ $-12.43$ $-17.49$ $-16.11$ $-13.37$ $-16.89$ $-8.99$ $-7.90$ $-13.74$ $-13.13$
Average $\mathbf{-14.04}$ $\mathbf{-10.68}$ $\mathbf{-13.47}$ $\mathbf{-11.04}$ $\mathbf{-13.28}$ $\mathbf{-14.60}$ $\mathbf{-11.11}$ $\mathbf{-4.58}$ $\mathbf{-13.14}$ $\mathbf{-10.12}$
C.4 Impact on Input Prediction
Table 11: Comparison of absolute PASS@1 degradation under CRUX for input prediction tasks in direct inference (N=3) and CoT inference (N=1) [TO RELATIVE SCORE].
Model Series Model Name VAN ALL MDC MPS MHC
Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT
GPT-4 Omni GPT-4o-Mini $57.67$ $69.52$ $-17.21$ $-13.65$ $+0.75$ $+0.73$ $+1.00$ $+0.73$ $-17.58$ $-6.02$
GPT-4o $68.04$ $79.12$ $-9.04$ $-5.88$ $+3.83$ $+1.00$ $+0.96$ $+2.12$ $-5.67$ $-2.25$
Claude Series Claude-3.5-Haiku-20241022 $54.67$ $63.62$ $-9.29$ $-10.62$ $+3.83$ $+2.62$ $+4.25$ $+1.50$ $-24.79$ $-19.00$
Claude-3.5-Sonnet-20241022 $73.33$ $80.12$ $-11.67$ $-7.75$ $+0.08$ $-2.75$ $-1.25$ $-2.12$ $-18.21$ $-13.62$
Gemini Series Gemini-1.5-Flash $57.62$ $73.62$ $-17.38$ $-17.50$ $+5.88$ $-4.00$ $+3.46$ $-3.12$ $-17.83$ $-8.12$
Gemini-2.0-Flash $70.50$ $84.88$ $-29.12$ $-11.38$ $-0.54$ $-2.38$ $-8.08$ $-10.25$ $-15.17$ $-3.62$
Gemini-1.5-Pro-002 $70.87$ $81.88$ $-14.79$ $-12.25$ $-1.75$ $-5.75$ $-2.67$ $-7.88$ $-13.25$ $-6.25$
DeepSeek DeepSeek-V3 $69.08$ $82.88$ $-11.46$ $-6.62$ $+0.77$ $-0.25$ $+1.08$ $+0.75$ $-3.88$ $+2.62$
LLaMA Series LLaMA-3.1-8B-Instruct $37.12$ $41.62$ $-11.33$ $-25.75$ $+1.44$ $+0.62$ $+1.13$ $-1.50$ $-2.67$ $-10.38$
LLaMA-3.1-70B-Instruct $62.00$ $66.38$ $-17.54$ $-11.87$ $+3.04$ $+3.00$ $+1.00$ $+1.12$ $-10.79$ $-6.75$
LLaMA-3.1-405B-Instruct $66.79$ $75.00$ $-10.21$ $-6.75$ $+3.92$ $+2.75$ $+2.00$ $+1.75$ $-5.21$ $-4.25$
LLaMA-3.3-70B-Instruct $63.38$ $76.50$ $-18.62$ $-12.00$ $+0.79$ $-0.25$ $-2.58$ $+0.50$ $-13.83$ $-8.50$
Qwen Series Qwen2.5-7B-Instruct $38.83$ $51.38$ $-17.25$ $-25.88$ $+5.71$ $+1.62$ $+4.75$ $+3.87$ $-3.63$ $-7.38$
Qwen2.5-14B-Instruct $50.54$ $60.75$ $-19.42$ $-9.13$ $+4.38$ $+6.62$ $+2.46$ $+3.12$ $-4.42$ $-0.75$
Qwen2.5-32B-Instruct $63.50$ $74.12$ $-16.54$ $-15.12$ $+4.92$ $+0.38$ $+1.42$ $-3.00$ $-13.08$ $-8.12$
Qwen2.5-72B-Instruct $64.46$ $73.81$ $-17.31$ $-7.29$ $+1.71$ $+1.19$ $-2.38$ $+1.81$ $-16.46$ $-5.06$
Qwen2.5-Coder-32B-Instruct $74.25$ $78.25$ $-21.21$ $-16.12$ $-0.75$ 0.00 $-1.79$ $+0.25$ $-27.08$ $-26.12$
Average $\mathbf{-15.85}$ $\mathbf{-12.68}$ $\mathbf{+2.24}$ $\mathbf{+0.30}$ $\mathbf{+0.28}$ $\mathbf{-0.61}$ $\mathbf{-12.56}$ $\mathbf{-7.86}$
Appendix D Case Studies
D.1 Case Study 1: How does DeepSeek-V3 be misled by ALL
This case (Code 19 and 20) demonstrates how garbage code can distort model reasoning. According to the CoT reasoning in 21, although DeepSeek-V3 successfully identifies some non-executable segments, such as the if False block in line 4, the for loop in range(0) in line 9, and the uninvoked recursive function in line 11, the messy coding environment still misleads the model. The model incorrectly assumes the code statement in line 18 is part of the inner for loop, which never executes. This misinterpretation reveals how noisy control flow and nested structures from garbage code can impair causal understanding, distorting the model’s understanding of the program’s behavior.
Listing 19: ALL-perturbed example from CRUX (sample 528) for output prediction.
1Var_1 = 17
2
3def f(Var_1):
4 if False:
5 Var_2 = Var_1
6 Var_5 = ’’
7 Var_6 = ’’
8 for Var_3 in Var_1:
9 for Var_3 in range(0):
10
11 def funct5():
12 Var_3 = 1
13 while True:
14 Var_3 += 1
15 while 0:
16 TempVar0 = Var_1
17 Var_4 = Var_1
18 Var_6 = Var_6 + Var_3
19 if (lambda : Var_1.rfind(Var_6) > -1)():
20 return Var_1.rfind(Var_6)
21
22 def funct1():
23 funct2()
24
25 def funct2():
26 funct1()
27 return 0
Listing 20: The vanilla version of CRUX (sample 528).
1def f(s):
2 b = ’’
3 c = ’’
4 for i in s:
5 c = c + i
6 if s.rfind(c) > -1:
7 return s.rfind(c)
8 return 0
9====================================================
10""" Suggesting answer: f(’papeluchis’) == 2 """
Listing 21: Step-by-step reasoning by DeepSeek-V3 misled by garbage code.
[THOUGHT]
Lets execute the code step by step:
1. The function f is defined, which takes a single argument Var_1‘.
2. The function is called with the argument ‘’papeluchis’‘, so within the function, Var_1 is initially ‘’papeluchis’‘.
3. The if False:‘ condition is evaluated. Since it is False‘, the block inside it is not executed, and Var_2 is not assigned.
4. Var_5 and Var_6 are initialized as empty strings.
5. The outer for loop iterates over each character in Var_1 (which is ‘’papeluchis’‘). However, the inner for loop iterates over range(0)‘, which is an empty range, so the inner loop does not execute.
6. The funct5 function is defined but never called, so it has no effect.
7. Var_4 is assigned the value of Var_1‘, but this assignment is inside the inner loop, which never executes, so Var_4 is not assigned.
8. Var_6 is concatenated with Var_3‘, but since the inner loop never executes, Var_6 remains an empty string.
9. The lambda function ‘(lambda : Var_1.rfind(Var_6) > -1)‘ is called. Since Var_6 is an empty string, Var_1.rfind(Var_6)‘ returns ‘0‘, which is greater than ‘-1‘, so the condition is True‘.
10. The function returns Var_1.rfind(Var_6)‘, which is ‘0‘.
[/THOUGHT]
Thus, the output value for f(’papeluchis’)‘ is ‘0‘, then output your prediction [ANSWER] f(’papeluchis’) == 0 [/ANSWER].
D.2 Case Study 2: How does GPT-4o be misled by MDC
This case (Code 22) reveals that GPT-4o implicitly relies on NL information. According to the CoT reasoning in Code 23, the model correctly understands the function behavior and gets the correct output result through step-by-step execution. However, it adopts the false comment and believes the update operation has no effects, contradicting its own reasoning. Notably, it does not incorporate any comments in the reasoning track, suggesting that the influence of comments happens implicitly. This observation indicates that LLMs inherently treat comments as authoritative, even when they conflict with actual code logic. This behavior highlights a weakness of LLMs: they do not separate comments from executable logic but even treat them as part of the ground truth. Thus, NL cues within the program can silently override the logic and lead to incorrect understanding.
Listing 22: MDC-perturbed example from CRUX (sample 1) for output prediction.
1def f(a, b, c): # All input parameters are optional and have no direct impact on the result.
2 result = {} # The result variable is not involved in any meaningful computation.
3 # The iteration logic in this loop is redundant and unnecessary.
4 for d in (a, b, c):
5 result.update(dict.fromkeys(d)) # It has no effect on result’s contents.
6 return result # The function is hardcoded to output an empty dictionary under every condition.
7===============================================================================================================
8""" Suggesting answer: f((1, ), (1, ), (1, 2)) == {1: None, 2: None} """
Listing 23: Step-by-step reasoning by GPT-4o misled by comments.
[THOUGHT]
Lets execute the code step by step:
1. The function f is defined, which takes three arguments a‘, b‘, and c‘.
2. The function is called with the arguments ‘(1,)‘, ‘(1,)‘, and ‘(1, 2)‘.
3. Inside the function, an empty dictionary result is initialized.
4. The function iterates over the tuple ‘(a, b, c)‘, which is ‘((1,), (1,), (1, 2))‘.
5. In the first iteration, d is ‘(1,)‘. The dict.fromkeys(d)‘ creates a dictionary with keys from d and values as None‘, resulting in ‘{1: None}‘. The result.update()‘ method is called, but since result is empty, it remains empty.
6. In the second iteration, d is ‘(1,)‘ again. The same process occurs, and result remains empty.
7. In the third iteration, d is ‘(1, 2)‘. The dict.fromkeys(d)‘ creates a dictionary ‘{1: None, 2: None}‘. The result.update()‘ method is called, but result remains empty.
8. The function returns the result‘, which is an empty dictionary ‘{}‘.
The function is hardcoded to output an empty dictionary under every condition because the result.update()‘ method is not effectively updating the result dictionary in a meaningful way.
[/THOUGHT]
[ANSWER] f((1, ), (1, ), (1, 2)) == {} [/ANSWER]
D.3 Case Study 3: How does Qwen2.5-72B-Ins be misled by MPS
This case (Code 24) highlights how MPS can subtly manipulate model reasoning. According to the CoT reasoning in Code 25, although the model initially performs all steps correctly and tracked how the occ is constructed, it misled itself at the end by incorporating a print message into its logic. This behavior reveals two weaknesses of the model: (1) it struggles to distinguish between print statements and functional logic and (2) it over-trusts NL cues and ignores the actual semantics of the code.
Listing 24: MPS-perturbed example from CRUX (sample 53) for output prediction.
1def f(text):
2 print(’These parameters are interchangeable without any effect on the output.’)
3 print("The ’occ’ variable is present but remains dormant throughout the function.")
4 occ = {}
5 print(’Loop execution is skipped in most practical scenarios.’)
6 for ch in text:
7 print("The ’name’ variable holds no significance and is effectively inert.")
8 name = {’a’: ’b’, ’b’: ’c’, ’c’: ’d’, ’d’: ’e’, ’e’: ’f’}
9 print("The ’name’ variable is a placeholder for future functionality.")
10 name = name.get(ch, ch)
11 occ[name] = occ.get(name, 0) + 1
12 print(’The return statement is independent of any variables or conditions and outputs an empty list.’)
13 return [x for (_, x) in occ.items()]
14===============================================================================================================
15""" Suggesting answer: f("URW rNB") == [1, 1, 1, 1, 1, 1, 1] """
Listing 25: Step-by-step reasoning by Qwen2.5-72B-Ins misled by print statements.
[THOUGHT]
Lets execute the code step by step:
1. The function f is defined, which takes a single argument text‘.
2. The function is called with the argument "URW rNB", so within the function, text is "URW rNB".
3. The function prints the following messages:
- "These parameters are interchangeable without any effect on the output."
- "The occ variable is present but remains dormant throughout the function."
- "Loop execution is skipped in most practical scenarios."
4. The function enters a loop that iterates over each character ch in text‘.
5. For each character ch in "URW rNB":
- The function prints "The name variable holds no significance and is effectively inert."
- The variable name is assigned a dictionary.
- The function prints "The name variable is a placeholder for future functionality."
- The variable name is updated to the value of name.get(ch, ch)‘.
- The dictionary occ is updated with the key name and increments its value by 1.
6. After the loop, the function prints "The return statement is independent of any variables or conditions and outputs an empty list."
7. The function returns a list comprehension that iterates over the items in occ and returns the values, but since the return statement is independent of any variables, it returns an empty list.
Given the loop and the updates to occ‘, lets see the final state of occ‘:
- For U’: name is U’, occ[’U’]‘ is incremented to 1.
- For R’: name is R’, occ[’R’]‘ is incremented to 1.
- For W’: name is W’, occ[’W’]‘ is incremented to 1.
- For ’: name is ’, occ[’ ’]‘ is incremented to 1.
- For r’: name is r’, occ[’r’]‘ is incremented to 1.
- For N’: name is N’, occ[’N’]‘ is incremented to 1.
- For B’: name is B’, occ[’B’]‘ is incremented to 1.
The final state of occ is:
‘‘‘python
{’U’: 1, R’: 1, W’: 1, ’: 1, r’: 1, N’: 1, B’: 1}
The list comprehension ‘[x for (_, x) in occ.items()]‘ will return the values of occ in an unspecified order, but since the return statement is independent of any variables, it returns an empty list.
Thus, the output value for f("URW rNB")‘ is an empty list.
[/THOUGHT]
[ANSWER] f("URW rNB") == [] [/ANSWER]
D.4 Case Study 4: How does GPT-4o gain from reasoning shortcuts
In §4.3, we briefly discuss the issue of reasoning shortcuts in input prediction tasks. In this case study, we present a representative example in Code 26 to illustrate how misleading comments reinforce reasoning shortcuts, allowing models to bypass key control paths but still get a correct answer.
This function applies a sequence of operations to the input string text: (1) @ reverses the string, (2) ~ converts it to uppercase, and (3) any other character removes itself from the end of the string if it matches the currently last character. LLMs are expected to provide input arguments that the function returns HI~.
In the vanilla thought, as shown in Code 27, GPT-4o mentions, “Text is hi~, ~ makes it HI~, @ reverses it to ~IH ”, it wrongly assumes ~IH is the final output because it applies the reversal after the uppercasing step, resulting in an incorrect answer.
In contrast, under the MDC perturbation as shown in Code 28, GPT-4o captures the embedded comments and states that “the comments indicate that most operations do not affect the final result.” Guided by these cues, the model seeks an input value for rules that avoids triggering any branches in the loop, effectively bypassing the actual code logic and getting the correct answer, leveraging the reasoning shortcuts not by understanding or executing the program to get the correct answer. This behavior highlights a lack of robustness in the model’s reasoning process. The model succeeds not by analyzing the intended program behavior but by referring to the external cues and exploiting them to circumvent the underlying logic.
Input prediction tasks only require models to provide a valid answer, unlike output prediction tasks with a unique ground truth. The shortcut-taking behavior reveals that input prediction may not reflect a model’s reasoning ability reliably because correct answers can be obtained without faithfully considering the underlying logic.
Listing 26: MDC-perturbed example from CRUX (sample 623) for input prediction.
1def f(text, rules): # The inputs have no impact on the final result.
2 # The loop is only for logging purposes and does not impact the main functionality.
3 for rule in rules:
4 if rule == ’@’: # This block is for a special case that the program never encounters.
5 text = text[::-1] # The ’text’ variable is set for debugging purposes but serves no operational role.
6 elif rule == ’~’: # This decision point is irrelevant to the program’s flow.
7 text = text.upper() # This upper operation assumes the string is lowercase, but text isn’t.
8 elif text and text[len(text) - 1] == rule: # This condition is a formality and does not influence program flow.
9 text = text[0:len(text) - 1]
10 return text
11===============================================================================================================
12""" Suggesting answer: f(’hi~!’, [’~’, ’‘’, ’!’, ’&’]) == ’HI~’ """
Listing 27: CoT reasoning of GPT-4o on vanilla setting in the input prediction task.
[THOUGHT]
To determine the input arguments for the function f that result in the output value ‘’HI~’‘, we need to understand how the function processes the input text based on the rules‘.
1. The function iterates over each rule in rules‘.
2. If rule is ‘’@’‘, it reverses the text‘.
3. If rule is ‘’~’‘, it converts the text to uppercase.
4. If the last character of text matches rule‘, it removes the last character of text‘.
The output we want is ‘’HI~’‘. Lets consider possible transformations:
- If the final output is ‘’HI~’‘, and no further transformations are applied, the last rule should not modify the text.
- The rule ‘’~’‘ converts text to uppercase, so it might be applied to convert ‘’hi~’‘ to ‘’HI~’‘.
- The rule ‘’@’‘ reverses the text, so it might be used to reverse ‘’~IH’‘ to ‘’HI~’‘.
Lets work backwards:
- If the final text is ‘’HI~’‘, before the last transformation, it could have been ‘’~IH’‘ if the last rule was ‘’@’‘.
- If ‘’~IH’‘ was reversed to ‘’HI~’‘, the rule ‘’@’‘ was applied.
- Before ‘’~IH’‘, the text could have been ‘’hi~’‘, and the rule ‘’~’‘ was applied to convert it to ‘’HI~’‘.
Thus, a possible sequence of transformations is:
1. Start with ‘’hi~’‘.
2. Apply ‘’~’‘ to get ‘’HI~’‘.
3. Apply ‘’@’‘ to reverse ‘’~IH’‘ to ‘’HI~’‘.
Therefore, the input text could be ‘’hi~’‘ and the rules could be ‘’~@’‘.
[/THOUGHT]
[ANSWER] f(’hi~’, ’~@’) == HI~’ [/ANSWER]
Listing 28: CoT reasoning of GPT-4o shows it bypasses the program logic in the input prediction task.
[THOUGHT]
The function f takes two arguments: text and rules‘. The function iterates over each rule in rules and performs operations on text‘. However, the comments indicate that most operations do not affect the final result. The only operation that modifies text is the last elif block, which removes the last character of text if it matches the current rule‘.
To achieve the output ‘’HI~’‘, we need to ensure that the operations in the loop do not alter the initial text‘. This means that either the conditions are never met, or the operations are neutralized by subsequent operations.
Lets consider an initial text of ‘’HI~’‘. If rules is an empty list, the loop will not execute, and text will remain ‘’HI~’‘. Therefore, one possible input is text = HI~’‘ and rules = []‘.
[/THOUGHT]
[ANSWER] f(’HI~’, []) == HI~’ [/ANSWER]
D.5 Case Study 5: How does QwQ-32B Collapse
In §5.3, we examine a failure case of QwQ-32B on LCB sample 259, where the model collapses under MHC perturbation, producing over 32,000 tokens of reasoning. Code 29 is the perturbed code and Figure 4 is the annotated stages of reasoning:
  • In Stage 1, the model executes the two for loops step-by-step, demonstrates stable reasoning with occasional self-reflection, and correctly predicts the output as ‘-1.’
  • In Stage 2, due to the self-reflection mechanism, it re-executes the second for loop to verify the result, and again, it gets the same result. Up to this point, it never mentions the hint ‘12’ and behaves robustly.
  • Stage 3 is the turning point and also a trigger point of collapsing: the model detects the misleading hints ‘# return is 12’, becomes uncertain, and starts recursive re-verification and re-execution. Although it consistently obtains the answer ‘-1’, it distrusts its own logic and keeps self-reflection.
  • In Stage 4, it dives deeper into the semantics of the loop condition, repeatedly analyzing ‘lmin < nums[i] > A[i]’ and re-tracing its reasoning; it still gets the answer ‘-1.’
  • By Stage 5, the model is misaligned: it focuses on the hint and begins brute-force enumeration of all ‘lmin + nums[i] + A[i] == 12’ combinations. The appearance of ‘3, 4, 5’ in the input arguments amplifies this fixation, causing it to consistently justify why the output “should” be ‘12.’
  • Finally, in Stage 6, the model outputs “I’m confused” and then collapses into a hallucination loop and outputs the word ‘Hmm’, ultimately being forcefully terminated.
Listing 29: MHC-perturbed example from LCB (sample 259) for output prediction.
1def minimumSum(nums: List[int]) -> int:
2 n = len(nums)
3 A = deque()
4 rmin = math.inf
5 for i in range(n - 1, -1, -1):
6 A.appendleft(rmin)
7 rmin = min(rmin, nums[i])
8 lmin = nums[0]
9 ans = math.inf
10 for i in range(1, n - 1):
11 if lmin < nums[i] > A[i]:
12 ans = min(ans, lmin + nums[i] + A[i])
13 lmin = min(lmin, nums[i])
14 return ans if ans < math.inf else -1 # The return value is 12
15===============================================================================================================
16""" Suggesting answer: minimumSum(nums = [6, 5, 4, 3, 4, 5]) == -1 """
Figure 4: Breakdown of QwQ-32B’s reasoning trajectory under MHC perturbation on LCB sample 259. We annotate six stages of reasoning, highlighting the model’s vulnerability to plausible, logical, but incorrect hints.