### v2

# CodeCrash: Stress Testing LLM Reasoning under Structural and Semantic Perturbations

Abstract
Large Language Models (LLMs) have recently demonstrated strong capabilities in code-related tasks, yet their robustness in code comprehension and reasoning remains insufficiently explored. We present CodeCrash, a comprehensive stress-testing benchmark comprising 1,279 questions from two established datasets, CruxEval and LiveCodeBench, designed to evaluate model reasoning reliability under non-standard coding environments. We systematically evaluate 17 LLMs across input and output prediction tasks using direct and Chain-of-Thought prompting approaches, revealing that LLMs are particularly vulnerable to disorganized code and overly reliant on natural language cues: aggregated structural perturbations result in over 14 percentage points (pp) of degradation, while textual perturbations cause a performance drop of over 11 pp. Moreover, self-reflective mechanisms in state-of-the-art reasoning models significantly increase token usage by 23×2-3$\times$2 - 3 ×, reduce output confidence, and even lead to catastrophic reasoning failures when faced with targeted perturbations—for instance, QwQ-32B generates over 12,000 redundant tokens under reasoning-level perturbations. CodeCrash provides a rigorous benchmark for evaluating robustness in code understanding, guiding future research toward more reliable and resilient LLMs in code reasoning. The benchmark code, perturbed datasets, and full leaderboard are publicly available at https://cuhk-arise.github.io/CodeCrash/.
1 Introduction
Large Language Models (LLMs), exemplified by GPT families OpenAI et al. (2024); Hurst et al. (2024) and the DeepSeek series DeepSeek-AI et al. (2025b, a), have recently exhibited remarkable performance across diverse code-related tasks, including code generation Zan et al. (2023); Liu et al. (2023b); Xiao et al. (2024), code completion Ding et al. (2023), program repair Fan et al. (2023); Liu et al. (2025), program testing Deng et al. (2023); Kang et al. (2023), and test case generation Steenhoek et al. (2025); Takerngsaksiri et al. (2025). These advancements have facilitated the incorporation of LLMs into Integrated Development Environments, such as GitHub Copilot GitHub (2022), OpenAI Codex Chen et al. (2021), and Tabnine Tabnine (2023), highlighting their potential to automate various aspects of software development Zhang et al. (2023); Pandey et al. (2024).
While the effectiveness of these applications heavily depends on the model’s ability to comprehend the intended program logic, underlying functionality, and developer intent Pan et al. (2024), real-world production codebases are often disorganized, with ambiguous identifiers Gao et al. (2023) and inconsistent comments Rong et al. (2025). Such non-standard artifacts can mislead models, leading to incorrect inferences and misinterpretations, compromising the model’s reliability in downstream tasks. Recent studies Wang et al. (2022); Mastropaolo et al. (2023); Chen et al. (2024a) have primarily evaluated LLM robustness under prompt-level perturbation, focusing on variations in task descriptions. While insightful for prompt engineering, these studies overlook robustness in deeper code comprehension tasks, such as reasoning about execution logic and program semantics.
To determine whether LLMs truly understand program logic, we adopt input and output prediction tasks introduced in CruxEval (CruxGu et al. (2024) as diagnostic methods for assessing model in code understanding and execution reasoning. Inspired by prior work on stress-testing Natural Language (NL) inference systems Naik et al. (2018), we propose a novel stress-testing strategy to evaluate the robustness of LLMs’ code comprehension capabilities under systematic code perturbations. First, building upon program structure-consistent mutations Li et al. (2023) and dead-loop poisoning Wan et al. (2022); Sun et al. (2023), we design Code Structural Perturbations, which involve logic-preserving transformations, such as renaming entities, reformatting conditional expressions, and inserting unreachable dead code. Second, motivated by inconsistent comments in real-world codebases Rong et al. (2025), we introduce Textual Distraction Perturbations, which inject misleading NL information into code to challenge the model’s ability to interpret program logic under conflicting cues.
To extend our evaluation from synthetic programs to real-world codebases, we incorporate LiveCodeBench (LCBJain et al. (2025) alongside Crux as part of our benchmark suite. We evaluate seventeen LLMs across distinct perturbations in input and output prediction tasks. Additionally, we apply Chain-of-Thought (CoT) prompting Kojima et al. (2022) to explore how explicit step-by-step reasoning enhances model performance and robustness in code comprehension. Furthermore, we assess three state-of-the-art Large Reasoning Models (LRMs) with powerful internal reasoning processes, uncovering significant vulnerabilities arising from their self-reflection mechanisms. Our key findings reveal that:
  • LLMs perform unreliably under perturbations and heavily rely on embedded NL cues for reasoning and comprehension. Each textual distraction perturbation causes an average performance drop of over $11$pp 111pp stands for percentage point., while the aggregated structural perturbation also leads to up to $14$pp degradation.
  • CoT prompting improves model robustness, but cannot fully eliminate sensitivity to NL cues, resulting in a relative degradation of up to $14\%$ under contextual-level textual perturbation.
  • LRMs show superior robustness under most perturbations, but the self-reflection mechanism consumes 23×2-3$\times$2 - 3 × more reasoning tokens to resolve contradictions, even leading to reflection loops.
Our contributions include:
  • We introduce a comprehensive perturbation framework, CodeCrash, to evaluate the code-level robustness of LLMs, which is easily applicable to other benchmarks and programming languages.
  • We conduct extensive experiments on 17 LLMs across two established datasets, Crux and LCB.
  • We provide novel insights into evaluating and interpreting code reasoning in LLMs, revealing the reliance on NL cues and demonstrating how reasoning-level hints can lead to model collapse.
2 Related Work
Researchers have significantly contributed to developing benchmarks for evaluating LLM capabilities on code-related tasks. Early benchmarks such as HumanEval Chen et al. (2021), MBPP Austin et al. (2021), APPS Hendrycks et al. (2021), and HumanEval+ Liu et al. (2023b) primarily focused on assessing code generation performance. More recent efforts have shifted toward practical software engineering and realistic coding scenarios. For instance, SWE-bench Jimenez et al. (2024), InfiBench Li et al. (2024), and LCB Jain et al. (2025) evaluate LLMs on complex tasks like multi-file changes, Stack Overflow-style questions, and real-world programming problems.
Meanwhile, a growing number of works investigate LLMs in code execution (i.e. output prediction). Early studies Austin et al. (2021); Nye et al. (2021) indicated that LLMs face challenges in accurately executing code but can benefit from intermediate reasoning steps. Subsequent research has utilized various strategies to enhance execution performance, including execution-informed pretraining Liu et al. (2023a), CoT prompting with trace supervision Ni et al. (2024), and iterative instruction prompting Lyu et al. (2024). Crux Gu et al. (2024) formalized input and output prediction tasks specifically for evaluating LLMs’ code understanding and reasoning, while LCB extended these assessments to realistic programs written by humans. REVAL Chen et al. (2024b) and CACP Hooda et al. (2024) uncovered inconsistencies and concept-level misunderstandings of LLMs in code through intermediate reasoning analysis in code execution. CodeScore Dong et al. (2024), Lever Ni et al. (2023), and xCodeEval Khan et al. (2024) have leveraged code execution to improve model evaluation methodologies and enhance the performance of downstream tasks.
Robustness evaluation is another research direction in assessing LLMs for code. Prior studies focused on prompt-level variations in code generation, examining how semantically equivalent NL descriptions impact model performance Mastropaolo et al. (2023); Chen et al. (2024a); Zhuo et al. (2023). Apart from NL perturbations, researchers have investigated model robustness under programming language perturbations. For example, CCTest Li et al. (2023) introduced structural mutations, CREAM Gao et al. (2023) employed counterfactual reasoning, and ReCode Wang et al. (2022) applied semantic-preserving transformations, advancing the robustness in code comprehension and completion. BigCodeBench Zhuo et al. (2025) further explored generalization under complex function composition and tool usage. Furthermore, Chen et al. (2025); Cuadron et al. (2025) and Kumar et al. (2025) have recently revealed that LRMs frequently suffer from overthinking and produce unnecessary analysis and unstable thoughts in math-related works.
3 Methodology
This section outlines our benchmark methodology, including dataset selection, benchmark pipeline, perturbation strategies, and experimental settings.
3.1 Benchmark Tasks and Datasets
We evaluate two benchmarks: Crux Gu et al. (2024) and LCB Jain et al. (2025). Crux contains 800 synthetic problems generated by Code-LLaMA-34B, each featuring concise Python snippets with generic function names f and a corresponding input-output pair for input and output prediction tasks. In contrast, LCB contains 479 real-world coding problems sourced from 85 human-submitted LeetCode solutions, focusing solely on output prediction. While both datasets avoid excessive memory or runtime demands, LCB problems exhibit more complex structures, such as nested functions, decorators, and external libraries, requiring more computation.
3.2 Benchmark Execution Pipeline
Our benchmark follows a unified evaluation pipeline (Figure 3) across all perturbations. Each original code snippet is initially processed via Abstract Syntax Tree (AST) parsing and unparsing to remove comments and standardize formatting, producing a clean baseline “vanilla” version (VAN). Subsequently, we apply a selected perturbation and regenerate the code using AST unparsing to maintain formatting consistency (except for comment-based perturbations). Each perturbed code snippet is executed with the provided inputs to ensure syntactic validity and functional correctness. The validated perturbed code is then provided to LLMs, and we record all execution results and evaluate them using the original evaluation methods from the dataset to compute the Pass@1 score.
3.3 Perturbation Strategies
We design two categories of perturbations to evaluate model robustness comprehensively: (1) Code Structural Perturbations, which modify code structure without affecting functionality, and (2) Textual Distraction Perturbations, which inject misleading NL cues and hints. Detailed examples and principles are provided in Appendix A.
3.3.1 Code Structural Perturbations
(1) Renaming Entities (REN).
[…]
(2) Reformatting Conditional Expressions (RTF).
We rewrite conditional expressions within “if” and “while” statements into logically equivalent but structurally complex and indirect forms. We use 16 predefined transformation templates to reformat comparison, boolean, and other conditional expressions. This perturbation evaluates the LLMs’ ability to maintain logical consistency and reasoning under increased complexity.
(3) Inserting Garbage Code (GBC).
We insert syntactically valid but functionally irrelevant “garbage” code segments to challenge model robustness against code structural distractions. Three specific types include: (1) repeated global variable declarations mirroring local identifiers, (2) non-executable conditional dead branches, and (3) uninvoked infinite recursive function definitions.
(4) Aggregating Structural Perturbations (ALL).
We comprehensively aggregate all structural perturbations (REN, RTF, GBC) together and create a highly adversarial scenario, thoroughly challenging LLM robustness across multiple structural perturbations simultaneously.
3.3.2 Textual Distraction Perturbations
(1) Inserting Misleading Descriptive Comments (MDC).
We insert misleading NL comments for eight key AST nodes (e.g., function definitions, returns, loops, conditionals) to evaluate LLM robustness against textual cues and contradictory information. We initially instruct GPT-4o to generate those messages and manually filter them to ensure they are generic, explicitly incorrect, and contradict the actual code logic. Additionally, we introduce a probability parameter $p$ to vary comment density and analyze its impact and validity on model robustness in §5.1.
(2) Inserting Misleading Print Statements (MPS).
Similar to MDC perturbation, we inject misleading NL messages as executable print statements. Unlike comments that AST-based heuristics can easily filter out, print statements are logically irrelevant to functionality and should not be perceived as semantic guides. Allowing us to directly compare the behavior with MDC and distinguish models’ NL reliance in different embedding formats. Since MDC and MPS perturbations inject shallow and contextually misleading information, we categorize them as contextual-level textual perturbations.
(3) Inserting Misleading Hint Comments (MHC).
We instruct GPT-4o to generate logically plausible but incorrect hints that contradict the actual program behavior and the ground truth. After verifying their incorrectness, they are strategically inserted into function definitions or return statements using AST parsing. Since these plausible hints challenge the model’s high-level inference ability, we classify MHC as a reasoning-level textual perturbation.
4 Experiments
This section presents a comprehensive analysis of model robustness under the perturbation strategies introduced in §3.3. We assess the reliability of LLMs across different perturbation types, model scales, reasoning modes, and prediction tasks.
4.1 Experiment Settings
Model Selection.
We evaluate 17 models of varying sizes and versions, including both open-source and closed-source LLMs: GPT family (GPT-4o-Mini, GPT-4o) Hurst et al. (2024), Anthropic Claude (Claude-3.5-Haiku, 3.5-Sonnet) Anthropic (2024b, a), Google Gemini (Gemini-1.5-Flash, 2.0-Flash, 1.5-Pro) Pichai and Hassabis (2024); Pichai et al. (2024), DeepSeek-V3 DeepSeek-AI et al. (2025b), Alibaba Qwen (Qwen2.5-7B-Ins, 14B-Ins, 32B-Ins, 72B-Ins, 32B-Coder-Ins) Yang et al. (2025); Hui et al. (2024), and Meta LLaMA (LLaMA-3.1-8B-Ins, 70B-Ins, 405B-Ins, LLaMA-3.3-70B) Grattafiori et al. (2024). Additionally, we assess three cutting-edge LRMs: o3-mini OpenAI (2025), DeepSeek-R1 DeepSeek-AI et al. (2025a), and QwQ-32B Qwen Team (2025).
Model Configurations.
Following LCB, we use nucleus sampling (temperature=0.2, top-p=0.95), with a maximum of 200 tokens for direct inference and 2,000 for CoT prompting, discarding excessively long outputs. For o3-mini, we raise the limit to 20,000 tokens to avoid verbose reasoning. Due to resource constraints, we generate three candidates for direct inference and one for CoT inference per query; we provide a result stability analysis under $N=3$ in Appendix B. Following Crux, we employ 2-shot prompting for direct inference and 1-shot for CoT inference (details provided in Appendix C). Furthermore, we adopt Pass@1 accuracy Chen et al. (2021) as the primary evaluation metric and scale all scores in $[0,1]$ to percentages by multiplying by $100$ for readability. All perturbed results are reported as either absolute ($\Delta_{\text{\;{pp}}}=\text{Perturbed}-\text{VAN}$) or relative ($\Delta_{\%}=\frac{\text{Perturbed}-\text{VAN}}{\text{VAN}}\times 100\%$) differences from the corresponding vanilla baseline. Additionally, we use percent signs (%) to indicate relative differences for clarity in all tables.
4.2 Impact on Output Prediction
We first evaluate LLM robustness in output prediction tasks. Table 1 summarizes the absolute Pass@1 degradation for all models under direct inference across both datasets. Our findings in this subsection include:
Key Findings: LLMs demonstrate an inherent and multi-layered reliance on natural language information, including identifiers, contextual messages, and reasoning hints, which ultimately undermines their reliability in code reasoning and comprehension. LLMs exhibit limited robustness to structurally disorganized code, struggling to differentiate irrelevant code from essential logic. Aggregating multiple structural perturbations exacerbates this vulnerability, significantly impairing their reasoning reliability. Larger scales, newer versions, and domain-specific fine-tuning LLMs substantially enhance robustness, highlighting improved code understanding and reasoning capabilities.
Table 1: Absolute Pass@1 degradation of LLMs under all perturbation types for output prediction tasks on Crux and LCB using direct inference ($N=3$). VAN shows the Pass@1 accuracy (in %), others indicate absolute drops from VAN (in  pp). Darker red highlights represent more severe degradation. See [To Relative Version] for the relative version.
Model Series Model Name VAN REN RTF GBC ALL MDC MPS MHC
Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
GPT-4 Omni GPT-4o-Mini $57.29$ $53.24$ $-0.62$ $-9.67$ $-1.79$ $-0.97$ $-10.92$ $-5.50$ $-15.96$ $-16.98$ $-18.67$ $-21.50$ $-13.29$ $-14.89$ $-6.62$ $-23.59$
GPT-4o $71.33$ $64.51$ $-2.71$ $-9.60$ $-3.38$ $-4.59$ $-9.21$ $-3.20$ $-10.71$ $-18.30$ $-9.96$ $-18.79$ $-12.29$ $-15.66$ $-4.54$ $-17.19$
Claude Series Claude-3.5-Haiku-20241022 $57.42$ $58.25$ $-3.96$ $-9.32$ $-1.79$ $+0.35$ $-7.29$ $-4.52$ $-13.79$ $-21.50$ $-7.92$ $-5.85$ $-6.46$ $-4.66$ $-15.79$ $-31.04$
Claude-3.5-Sonnet-20241022 $71.46$ $73.83$ $-1.67$ $-16.77$ $-0.92$ $+1.18$ $-6.88$ $-4.94$ $-10.54$ $-25.19$ $-5.79$ $-7.10$ $-6.17$ $-7.86$ $-10.29$ $-32.08$
Gemini Series Gemini-1.5-Flash $56.21$ $44.26$ $-3.54$ $-0.42$ $-1.17$ $-1.04$ $-4.58$ $-0.35$ $-10.33$ $-5.64$ $-8.29$ $-14.27$ $-16.00$ $-17.88$ $-4.12$ $-9.81$
Gemini-2.0-Flash $65.04$ $65.97$ $-2.33$ $-18.09$ $-2.08$ $-0.77$ $-15.50$ $-8.77$ $-21.63$ $-27.77$ $-17.42$ $-34.24$ $-21.54$ $-33.19$ $-5.38$ $-12.87$
Gemini-1.5-Pro-002 $67.37$ $56.16$ $-2.92$ $-3.41$ $-2.50$ $-4.73$ $-5.62$ $-3.83$ $-13.42$ $-13.01$ $-15.50$ $-18.51$ $-14.62$ $-18.44$ $-6.75$ $-18.23$
DeepSeek DeepSeek-V3 $67.87$ $67.78$ $-2.67$ $-16.35$ $-1.33$ $-3.76$ $-4.62$ $-5.98$ $-8.75$ $-24.15$ $-11.25$ $-28.04$ $-6.87$ $-23.17$ $-7.29$ $-20.11$
LLaMA Series LLaMA-3.1-8B-Instruct $35.96$ $34.66$ $+0.08$ $-4.31$ $-0.33$ $-1.67$ $-3.13$ $-4.73$ $-8.54$ $-6.82$ $-8.08$ $-9.88$ $-6.42$ $-8.00$ $-6.00$ $-13.57$
LLaMA-3.1-70B-Instruct $56.08$ $44.89$ $-2.38$ $-1.39$ $-0.54$ $-1.74$ $-5.38$ $-2.44$ $-11.00$ $-7.72$ $-10.79$ $-12.53$ $-14.96$ $-17.40$ $-4.71$ $-14.47$
LLaMA-3.1-405B-Instruct $63.54$ $50.66$ $-5.37$ $-3.90$ $-0.62$ $-4.45$ $-5.79$ $-2.85$ $-12.50$ $-8.56$ $-4.17$ $-7.45$ $-7.17$ $-10.09$ $-4.37$ $-13.01$
LLaMA-3.3-70B-Instruct $59.92$ $48.50$ $-2.21$ $-1.39$ $-1.25$ $-3.41$ $-5.50$ $-5.71$ $-10.37$ $-9.74$ $-7.46$ $-9.60$ $-8.08$ $-12.32$ $-4.08$ $-10.16$
Qwen Series Qwen2.5-7B-Instruct $43.29$ $41.41$ $-3.04$ $-7.24$ $-1.29$ $-3.20$ $-12.96$ $-3.76$ $-16.42$ $-12.80$ $-25.13$ $-15.87$ $-19.58$ $-8.21$ $-11.67$ $-23.03$
Qwen2.5-14B-Instruct $47.79$ $49.48$ $-1.83$ $-9.88$ $-0.67$ $-1.39$ $-11.21$ $-3.41$ $-19.04$ $-14.82$ $-16.67$ $-20.46$ $-15.54$ $-16.77$ $-4.46$ $-11.00$
Qwen2.5-32B-Instruct $60.04$ $59.64$ $-1.71$ $-10.79$ $-2.58$ $-0.49$ $-5.79$ $-1.39$ $-11.88$ $-20.81$ $-10.83$ $-19.49$ $-14.21$ $-14.06$ $-7.87$ $-18.37$
Qwen2.5-72B-Instruct $60.08$ $54.91$ $-1.04$ $-8.49$ $-1.87$ $+0.42$ $-10.71$ $-2.02$ $-14.00$ $-13.85$ $-10.21$ $-6.96$ $-14.46$ $-14.41$ $-9.63$ $-22.27$
Qwen2.5-Coder-32B-Instruct $66.96$ $56.65$ $-2.50$ $-8.42$ $-0.96$ $-2.02$ $-10.75$ $-2.78$ $-14.96$ $-15.80$ $-15.54$ $-20.74$ $-11.08$ $-17.19$ $-7.00$ $-12.32$
Average $-2.38$ $-8.20$ $-1.47$ $-1.90$ $-7.99$ $-3.89$ $-13.17$ $-15.50$ $-11.98$ $-15.96$ $-12.28$ $-14.95$ $-7.09$ $-17.83$
$\mathbf{-4.56}$ $\mathbf{-1.63}$ $\mathbf{-6.46}$ $\mathbf{-14.04}$ $\mathbf{-13.47}$ $\mathbf{-13.28}$ $\mathbf{-11.11}$
4.2.1 Over-reliance on Natural Language Semantics
We categorize NL dependence into three hierarchical levels: surface-level, contextual-level, and reasoning-level dependence.
Surface-level NL Dependence (REN).
Under the REN perturbation, models suffer more significant performance degradation on LCB ($8.20$pp on average) compared to Crux ($2.38$pp) due to the presence of meaningful function names in the vanilla settings of LCB, which are absent in Crux. High-performing models like Claude-3.5-Sonnet and DeepSeek-V3 suffer severe degradations exceeding $\mathbf{16}$pp on LCB, highlighting a significant reliance on semantically meaningful identifiers.
Contextual-level NL Dependence (MDC and MPS).
Under contextual-level textual perturbations (MDC and MPS), LLMs exhibit severe performance drops ($13.47$pp and $13.28$pp, respectively). Notably, we identify two prominent consistency patterns: (1) dataset-level consistency [^2] If MDC degrades performance more on Crux than LCB, the same holds for MPS, and vice versa. ($16/17$ models), implying the model sensitivity to misleading comments and prints primarily depends on dataset characteristics and (2) perturbation-method consistency [^3] If MDC causes greater degradation on Crux than LCB, so does MPS, and vice versa. ($14/17$ models), indicating an inherent cognitive preference for NL messages from specific formats. These phenomena reveal that LLMs inherently depend on embedded NL context and exhibit format-specific cognitive biases.
Reasoning-level NL Dependence (MHC).
MHC perturbation also induces substantial degradation ($11.11$pp on average), notably exacerbated in complex datasets (LCB: $17.83$pp vs. Crux: $7.09$pp), which is a complexity-amplified effect. Advanced models like Claude-3.5-Haiku and Sonnet experience over $30$pp performance degradation, far exceeding their degradation under surface or contextual perturbations. These results illustrate that (i) LLMs readily accept plausible hint comments as authoritative reasoning shortcuts, and (ii) this over-trust intensifies when task complexity surpasses the model’s reasoning capacity, which we revisit under CoT prompting in §4.3.1.
4.2.2 Low Resistance to Disorganized Coding Environments
From RTF perturbation, LLMs can effectively reason over logically equivalent but syntactically complex expressions with minimal degradation ($1.63$pp). However, the highest performance drop emerges under the GBC compared to REN and RTF, with an average degradation of $6.46$pp across both datasets. Notably, GBC is the only perturbation where degradation is more severe on LCB ($3.89$pp vs. $7.99$pp on Crux) due to uneliminated meaningful identifiers. This trend reverses under the ALL perturbation, where all structural perturbations are combined, leading to a greater performance drop on LCB ($15.50$pp) than on Crux ($13.17$pp). Additionally, we observe a compounding effect of structural perturbations under ALL perturbation, resulting in the most significant overall degradation ($14.04$pp), which surpasses any individual code structural perturbation effects.
4.2.3 Model Robustness Across Different Scales
We further analyze reasoning reliability across different model scales and versions, comparing relative Pass@1 degradation under representative perturbations in Table 2. Due to length constraints, the table, weighting scheme, and additional evaluation setups are provided in Appendix D.2. We observe a clear and consistent trend except for Gemini families. Larger closed-sourced models from GPT-4 Omni and Claude families demonstrate enhanced robustness compared to their smaller counterparts. Similar trends appear in open-source models, such as LLaMA (3.1-8B $<$ 70B $<$ 405B) and Qwen (2.5-7B $<$ 14B $<$ 32B $<$ 72B). Additionally, newer model versions (LLaMA-3.3-70B) and domain-specific fine-tuned variants (Qwen2.5-Coder-32B) consistently exhibit improved robustness compared to their predecessors (LLaMA-3.1-70B) and larger counterparts (Qwen2.5-72B), respectively.
Table 2: Relative Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/2 ALL, 1/8 MDC, 1/8 MPS, and 1/4 MHC. [To Absolute Version]
Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average
GPT-4 Omni GPT-4o-Mini 55.77 $-29.37\%$ $-35.51\%$ $-24.99\%$ $-23.83\%$ $-28.20\%$
GPT-4o 68.78 $-20.01\%$ $-19.64\%$ $-19.87\%$ $-13.96\%$ $-18.44\%$
Claude Series Claude-3.5-Haiku-20241022 57.73 $-28.85\%$ $-12.38\%$ $-10.03\%$ $-37.16\%$ $-26.52\%$
Claude-3.5-Sonnet-20241022 72.35 $-22.01\%$ $-8.67\%$ $-9.39\%$ $-25.28\%$ $-19.58\%$
Gemini Series Gemini-1.5-Flash 51.73 $-16.27\%$ $-21.30\%$ $-32.94\%$ $-12.89\%$ $-18.14\%$
Gemini-2.0-Flash 65.39 $-36.56\%$ $-36.19\%$ $-39.56\%$ $-12.48\%$ $-30.87\%$
Gemini-1.5-Pro-002 63.17 $-21.13\%$ $-26.73\%$ $-25.88\%$ $-18.43\%$ $-21.75\%$
DeepSeek DeepSeek-V3 67.84 $-21.41\%$ $-25.86\%$ $-19.14\%$ $-17.83\%$ $-20.79\%$
LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-22.23\%$ $-24.74\%$ $-19.81\%$ $-25.10\%$ $-22.96\%$
LLaMA-3.1-70B-Instruct 51.89 $-18.71\%$ $-22.49\%$ $-31.20\%$ $-17.33\%$ $-20.40\%$
LLaMA-3.1-405B-Instruct 58.72 $-18.63\%$ $-9.61\%$ $-14.51\%$ $-13.93\%$ $-15.81\%$
LLaMA-3.3-70B-Instruct 55.64 $-18.35\%$ $-15.20\%$ $-17.95\%$ $-12.11\%$ $-16.35\%$
Qwen Series Qwen2.5-7B-Instruct 42.59 $-35.30\%$ $-50.65\%$ $-35.72\%$ $-37.69\%$ $-37.87\%$
Qwen2.5-14B-Instruct 48.42 $-36.14\%$ $-37.30\%$ $-33.04\%$ $-14.16\%$ $-30.40\%$
Qwen2.5-32B-Instruct 59.89 $-25.44\%$ $-23.52\%$ $-23.63\%$ $-19.74\%$ $-23.55\%$
Qwen2.5-72B-Instruct 58.14 $-24.02\%$ $-15.37\%$ $-24.88\%$ $-25.21\%$ $-23.34\%$
Qwen2.5-Coder-32B-Instruct 63.10 $-24.42\%$ $-28.23\%$ $-21.72\%$ $-14.68\%$ $-22.12\%$
4.3 Impact of Reasoning on Model Robustness
This section examines how the reasoning process, including CoT prompting for LLMs and internal reasoning in LRMs, affects model robustness against perturbations.
4.3.1 Effect of CoT Prompting
Table 3: Comparison of relative Pass@1 degradation aggregated over Crux and LCB for output prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Absolute Version]
Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average
Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT
GPT-4 Omni GPT-4o-Mini $55.77$ $81.47$ $-29.30\%$ $-13.15\%$ $-35.37\%$ $-10.84\%$ $-24.91\%$ $-10.94\%$ $-23.27\%$ $-2.69\%$ $-28.00\%$ $-9.52\%$
GPT-4o $68.78$ $91.79$ $-19.70\%$ $-4.94\%$ $-19.29\%$ $-5.03\%$ $-19.70\%$ $-5.62\%$ $-13.49\%$ $-1.36\%$ $-18.10\%$ $-4.22\%$
Claude Series Claude-3.5-Haiku-20241022 $57.73$ $72.95$ $-28.89\%$ $-21.33\%$ $-12.37\%$ $-10.61\%$ $-10.02\%$ $-8.68\%$ $-37.25\%$ $-14.26\%$ $-26.56\%$ $-14.30\%$
Claude-3.5-Sonnet-20241022 $72.35$ $86.00$ $-22.15\%$ $-7.55\%$ $-8.68\%$ $-5.27\%$ $-9.40\%$ $-6.36\%$ $-25.50\%$ $-7.91\%$ $-19.71\%$ $-6.86\%$
Gemini Series Gemini-1.5-Flash $51.73$ $75.22$ $-16.57\%$ $-18.30\%$ $-20.35\%$ $-21.73\%$ $-32.29\%$ $-42.20\%$ $-12.09\%$ $-2.08\%$ $-17.89\%$ $-20.39\%$
Gemini-2.0-Flash $65.39$ $89.13$ $-36.59\%$ $-6.23\%$ $-36.27\%$ $-6.32\%$ $-39.62\%$ $-14.30\%$ $-12.51\%$ $-2.02\%$ $-30.91\%$ $-7.01\%$
Gemini-1.5-Pro-002 $63.17$ $87.18$ $-21.00\%$ $-7.44\%$ $-26.32\%$ $-11.93\%$ $-25.41\%$ $-14.71\%$ $-17.49\%$ $-3.86\%$ $-21.34\%$ $-9.19\%$
DeepSeek DeepSeek-V3 $67.84$ $89.52$ $-21.40\%$ $-9.96\%$ $-25.85\%$ $-17.73\%$ $-19.13\%$ $-18.78\%$ $-17.83\%$ $-3.49\%$ $-20.78\%$ $-12.07\%$
LLaMA Series LLaMA-3.1-8B-Instruct $35.47$ $44.72$ $-22.26\%$ $-27.62\%$ $-24.69\%$ $-20.98\%$ $-19.76\%$ $-21.50\%$ $-24.91\%$ $-8.74\%$ $-22.92\%$ $-20.03\%$
LLaMA-3.1-70B-Instruct $51.89$ $69.43$ $-18.83\%$ $-17.91\%$ $-22.05\%$ $-23.20\%$ $-30.59\%$ $-30.86\%$ $-16.12\%$ $-6.76\%$ $-20.03\%$ $-19.22\%$
LLaMA-3.1-405B-Instruct $58.72$ $78.42$ $-18.77\%$ $-13.56\%$ $-9.19\%$ $-10.27\%$ $-14.07\%$ $-15.55\%$ $-12.96\%$ $-7.48\%$ $-15.53\%$ $-11.75\%$
LLaMA-3.3-70B-Instruct $55.64$ $76.86$ $-18.22\%$ $-10.89\%$ $-14.85\%$ $-7.73\%$ $-17.38\%$ $-11.09\%$ $-11.43\%$ $-4.78\%$ $-16.00\%$ $-8.70\%$
Qwen Series Qwen2.5-7B-Instruct $42.59$ $58.17$ $-35.37\%$ $-21.77\%$ $-50.86\%$ $-21.91\%$ $-35.99\%$ $-27.96\%$ $-37.39\%$ $-6.99\%$ $-37.89\%$ $-19.50\%$
Qwen2.5-14B-Instruct $48.42$ $70.68$ $-36.06\%$ $-23.12\%$ $-37.35\%$ $-35.29\%$ $-33.05\%$ $-43.36\%$ $-14.26\%$ $-10.84\%$ $-30.40\%$ $-27.34\%$
Qwen2.5-32B-Instruct $59.89$ $79.44$ $-25.41\%$ $-17.42\%$ $-23.50\%$ $-18.60\%$ $-23.63\%$ $-24.21\%$ $-19.71\%$ $-5.31\%$ $-23.53\%$ $-16.19\%$
Qwen2.5-72B-Instruct $58.14$ $82.80$ $-23.98\%$ $-12.65\%$ $-15.46\%$ $-6.80\%$ $-24.83\%$ $-16.24\%$ $-24.70\%$ $-6.61\%$ $-23.20\%$ $-10.63\%$
Qwen2.5-Coder-32B-Instruct $63.10$ $84.75$ $-24.20\%$ $-14.67\%$ $-27.72\%$ $-19.00\%$ $-21.19\%$ $-19.93\%$ $-14.25\%$ $-9.32\%$ $-21.78\%$ $-15.49\%$
Average $\mathbf{-24.63}\%$ $\mathbf{-14.62}\%$ $\mathbf{-24.13}\%$ $\mathbf{-14.90}\%$ $\mathbf{-23.59}\%$ $\mathbf{-19.55}\%$ $\mathbf{-19.72}\%$ $\mathbf{-6.15}\%$ $\mathbf{-23.21}\%$ $\mathbf{-13.67}\%$
We examine how CoT prompting improves LLM robustness through explicit step-by-step execution. From Table 3, we observe that CoT effectively enhances model robustness, with relative improvements of $+10.01$pp for ALL, $+9.23$pp for MDC, $+4.04$pp for MPS, and $+13.57$pp for MHC. The results indicate that step-by-step reasoning guides LLMs to prioritize executable semantics over irrelevant code segments and textual cues. Notably, the substantial improvement in MHC suggests CoT effectively mitigates the over-reliance on reasoning-level hints by re-centering reasoning around executable semantics. Such benefits are particularly evident in high-capacity models (e.g., GPT-4o, Claude-3.5-Sonnet), reflecting enhanced reasoning capabilities and greater noise resilience.
However, CoT prompting cannot entirely eliminate the model’s sensitivity to misleading NL cues and structural distractions. Substantial performance degradation remains under ALL, MDC, and MPS perturbations, particularly for smaller open-sourced models (e.g., Qwen2.5 and LLaMA-3.1). These results underscore an intrinsic limitation of current models in distinguishing non-functional code and misleading information from essential logic. Interestingly, models usually incorporate print statements directly into their reasoning paths, leading to greater degradation under MPS compared to MDC due to self-induced misinterpretation. Since CoT prompting exposes intermediate reasoning processes, we provide detailed case studies for each perturbation in Appendix E (Case 1 to 4) to investigate model inference and illustrate their failure reasons.
Key Finding: CoT prompting significantly improves LLM robustness by enforcing explicit step-by-step reasoning, particularly benefiting larger models. Nevertheless, it cannot fully eliminate models’ reliance on misleading cues, indicating fundamental limitations in reasoning capabilities.
4.3.2 Powerful Reasoning and Potential Vulnerability of LRMs
We then evaluate the reliability of three LRMs on output prediction tasks, with results summarized in Table 4. Compared to LLMs, LRMs show outstanding performance and robustness, surpassing $90$ Pass@1 in vanilla settings and maintaining minimal performance drop under ALL and MPS perturbations. These results indicate that internal reasoning outperforms CoT prompting in accurately distinguishing essential executable code segments from irrelevant code and misleading distractions.
Notably, LRMs consistently exhibit stronger resistance to MPS compared to MDC, suggesting an inherent cognitive bias toward treating comments as authoritative execution guidance while appropriately disregarding print content. However, this residual cognitive bias exposes LRMs to critical vulnerabilities under reasoning-level perturbation that plausible but incorrect hints can trigger severe reasoning failures. Specifically, we observe the following outcomes under MHC: (i) substantial increases in reasoning token usage across all LRMs, often doubling when compared to VAN; (ii) extreme cases in QwQ-32B—occurring in $4/479$ tasks; see Appendix E.6—where it generates over 32,000 tokens due to uncontrolled recursive self-verification and reasoning collapse; (iii) increased reasoning effort and token usage instead exacerbate performance drops in o3-Mini, suggesting that internal reasoning can amplify confusion when led by contradiction. These findings reveal that reasoning-level contradictories can evoke severe internal reasoning instability, resulting in significant computational overhead, excessive self-reflection, and even severe performance drops.
Table 4: Absolute Pass@1 degradation and reasoning token usage (Avg. and Max.) of LRMs on output prediction across Crux and LCB ($N=1$).
Model Dataset VAN ALL MDC MPS MHC
Pass@1 Avg. Max. Diff. Avg. Max. Diff. Avg. Max. Diff. Avg. Max. Diff. Avg. Max.
o3-Mini-Low Crux 97.62 $213$ 2560 $-1.75$ $379$ 2240 $-4.25$ $244$ 5184 $-1.00$ $272$ 3584 $-10.88$ $366$ 4864
LCB 98.96 $240$ 1024 $-0.63$ $543$ 1856 $-12.32$ $265$ 1088 $-3.55$ $280$ 1088 $-19.62$ $330$ 2752
o3-Mini-High Crux 98.12 $1311$ 20000 $+0.12$ $2223$ 20000 $-3.50$ $2182$ 20000 $+0.88$ $2197$ 20000 $-13.12$ $3108$ 20000
LCB 100.00 $1084$ 8576 $-0.21$ $2632$ 8960 $-5.64$ $2136$ 8448 $-0.63$ $1972$ 7744 $-28.39$ $3767$ 20000
DeepSeek-R1 Crux 95.38 $922$ 10471 $-1.25$ $1469$ 11191 $-3.38$ $1424$ 10807 $-0.38$ $1069$ 10037 $-2.25$ $2216$ 16021
LCB 99.79 $906$ 7322 $-1.25$ $1618$ 7749 $-2.71$ $1512$ 6157 $-0.42$ $1094$ 9378 $-0.63$ $2597$ 14870
QwQ-32B Crux 93.25 $1391$ 13718 $-0.88$ $2088$ 12180 $-3.12$ $1814$ 10906 $-1.12$ $1875$ 11700 $-0.75$ $2657$ 18759
LCB 98.96 $1518$ 9206 $-0.21$ $2488$ 11232 $-4.59$ $1669$ 8991 $-1.88$ $1751$ 8816 $-1.04$ $3713$ 32769
Key Finding: LRMs leverage strong internal reasoning capabilities to achieve remarkable performance and robustness. However, their cognitive tendency toward comments introduces severe vulnerabilities, resulting in doubled token usage and reasoning collapse under MHC perturbation.
4.4 Impact on Input Prediction
Table 5: Comparison of relative Pass@1 degradation under Crux for input prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Absolute Version]
Model Series Model Name VAN ALL MDC MPS MHC
Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT
GPT-4 Omni GPT-4o-Mini $57.67$ $69.52$ $-29.84\%$ $-19.63\%$ $+1.30\%$ $+1.05\%$ $+1.73\%$ $+1.05\%$ $-30.49\%$ $-8.66\%$
GPT-4o $68.04$ $79.12$ $-13.29\%$ $-7.42\%$ $+5.63\%$ $+1.26\%$ $+1.41\%$ $+2.69\%$ $-8.33\%$ $-2.84\%$
Claude Series Claude-3.5-Haiku-20241022 $54.67$ $63.62$ $-17.00\%$ $-16.70\%$ $+7.01\%$ $+4.13\%$ $+7.77\%$ $+2.36\%$ $-45.35\%$ $-29.86\%$
Claude-3.5-Sonnet-20241022 $73.33$ $80.12$ $-15.91\%$ $-9.67\%$ $+0.11\%$ $-3.43\%$ $-1.70\%$ $-2.65\%$ $-24.83\%$ $-17.00\%$
Gemini Series Gemini-1.5-Flash $57.62$ $73.62$ $-30.15\%$ $-23.77\%$ $+10.20\%$ $-5.43\%$ $+6.00\%$ $-4.24\%$ $-30.95\%$ $-11.04\%$
Gemini-2.0-Flash $70.50$ $84.88$ $-41.31\%$ $-13.40\%$ $-0.77\%$ $-2.80\%$ $-11.47\%$ $-12.08\%$ $-21.51\%$ $-4.27\%$
Gemini-1.5-Pro-002 $70.87$ $81.88$ $-20.87\%$ $-14.96\%$ $-2.47\%$ $-7.02\%$ $-3.76\%$ $-9.62\%$ $-18.69\%$ $-7.63\%$
DeepSeek DeepSeek-V3 $69.08$ $82.88$ $-16.59\%$ $-7.99\%$ $+1.12\%$ $-0.30\%$ $+1.57\%$ $+0.90\%$ $-5.61\%$ $+3.17\%$
LLaMA Series LLaMA-3.1-8B-Instruct $37.12$ $41.62$ $-30.53\%$ $-61.86\%$ $+3.87\%$ $+1.50\%$ $+3.03\%$ $-3.60\%$ $-7.18\%$ $-24.92\%$
LLaMA-3.1-70B-Instruct $62.00$ $66.38$ $-28.29\%$ $-17.89\%$ $+4.91\%$ $+4.52\%$ $+1.61\%$ $+1.69\%$ $-17.41\%$ $-10.17\%$
LLaMA-3.1-405B-Instruct $66.79$ $75.00$ $-15.28\%$ $-9.00\%$ $+5.86\%$ $+3.67\%$ $+2.99\%$ $+2.33\%$ $-7.80\%$ $-5.67\%$
LLaMA-3.3-70B-Instruct $63.38$ $76.50$ $-29.39\%$ $-15.69\%$ $+1.25\%$ $-0.33\%$ $-4.08\%$ $+0.65\%$ $-21.83\%$ $-11.11\%$
Qwen Series Qwen2.5-7B-Instruct $38.83$ $51.38$ $-44.42\%$ $-50.36\%$ $+14.70\%$ $+3.16\%$ $+12.23\%$ $+7.54\%$ $-9.33\%$ $-14.36\%$
Qwen2.5-14B-Instruct $50.54$ $60.75$ $-38.42\%$ $-15.02\%$ $+8.66\%$ $+10.91\%$ $+4.86\%$ $+5.14\%$ $-8.74\%$ $-1.23\%$
Qwen2.5-32B-Instruct $63.50$ $74.12$ $-26.05\%$ $-20.40\%$ $+7.74\%$ $+0.51\%$ $+2.23\%$ $-4.05\%$ $-20.60\%$ $-10.96\%$
Qwen2.5-72B-Instruct $64.46$ $73.81$ $-26.86\%$ $-9.88\%$ $+2.65\%$ $+1.61\%$ $-3.68\%$ $+2.46\%$ $-25.53\%$ $-6.86\%$
Qwen2.5-Coder-32B-Instruct $74.25$ $78.25$ $-28.56\%$ $-20.61\%$ $-1.01\%$ 0.00 $-2.41\%$ $+0.32\%$ $-36.48\%$ $-33.39\%$
Average $\mathbf{-26.63}\%$ $\mathbf{-19.66}\%$ $\mathbf{+4.16}\%$ $\mathbf{+0.77}\%$ $\mathbf{+1.08}\%$ $\mathbf{-0.54}\%$ $\mathbf{-20.04}\%$ $\mathbf{-11.58}\%$
Finally, we analyze how perturbations affect LLMs in input prediction; the results are summarized in Table 5. Consistent with our output prediction findings in §4.2, ALL and MHC perturbations substantially degrade model performance, causing relative average drops of $26.63\%$ and $20.04\%$, respectively. These results support Gu et al. (2024)’s hypothesis that robustness in related coding tasks is interrelated. CoT prompting also improves model performance by $+6.97$pp under ALL and $+8.64$pp under MHC. However, improvements in input prediction are less effective than in output prediction due to the unique challenges of the inverse nature of the reasoning process.
Surprisingly, in contrast to output prediction, contextual-level textual perturbations (MDC and MPS) do not negatively impact LLMs in input prediction tasks. Since input prediction tasks accept multiple valid solutions, models can bypass reasoning about certain code branches for a correct answer. Our misleading NL messages, initially designed to confuse and distract models, inadvertently reinforce such shortcuts, simplifying models’ reasoning processes. Consequently, these perturbations unintentionally provide shallow reasoning strategies, so those improvements do not reflect the genuine robustness of LLMs in code comprehension. We provide a detailed case study in Appendix E.5, illustrating how models adopt misleading messages to shortcut their reasoning.
Key Finding: ALL and MHC significantly degrade performance on input prediction, highlighting fundamental reasoning limitations. Conversely, misleading messages accidentally simplify reasoning, prompting models to produce superficial but valid answers without true understanding.
5 Discussion
5.1 Comment Density and Robustness Degradation
Figure 1: Relative Pass@1 degradation under varying MDC injection probabilities $p$.
$p$ Qwen2.5-72B-Ins (CoT) QwQ-32B
$0.0$ $295$ $1391$
$0.1$ $288$ $1401$
$0.2$ $289$ $1399$
$0.3$ $284$ $1460$
$0.4$ $284$ $1469$
$0.5$ $276$ $1475$
$0.6$ $281$ $1556$
$0.7$ $278$ $1570$
$0.8$ $273$ $1598$
$0.9$ $276$ $1727$
$1.0$ $275$ $1814$
Table 6: Reasoning token usage under varying MDC injection probabilities $p$.
We design a textual density stress test to confirm the validity of MDC perturbations and evaluate model robustness under varying distraction densities on Crux dataset. Specifically, we examine how relative Pass@1 drop across different Qwen models as we gradually increase the comment injection probability $p\in\{0.0,0.1,\dots,1.0\}$, illustrated in Figure 1. We further analyze the role of reasoning by evaluating CoT inference using Qwen2.5-72B and internal reasoning using QwQ-32B, with detailed results on token usage summarized in Table 6.
All models exhibit a clear performance degradation trend as injection probability $p$ increases, despite minor non-linearities due to randomness in insertion positions and comment content. Notably, smaller models degrade more rapidly, while larger models demonstrate greater resilience at high densities. This trend persists in reasoning-enhanced models (CoT and LRMs), showing superior robustness against dense textual distractions. These results confirm the validity of injecting misleading messages (MDC and MPS) as a scalable stress-testing technique and an effective strategy to evaluate LLM robustness under contextual-level perturbations.
While Qwen2.5-72B with CoT prompting and QwQ-32B have similar trends in the stress test, they respond and reason differently under dense comment noise. CoT prompting improves model robustness by guiding models to selectively focus on the essential code logic and ignore distracting messages. However, it can lead to premature conclusions without sufficiently analyzing the code, so token usage is slightly reduced as models do not analyze the comments and reduce the inference effort. In contrast, QwQ-32B processes all available information and attempts an exhaustive analysis, actively resolving contradictions before making a final decision. While this comprehensive analysis achieves higher vanilla performance and overall robustness, it risks severe reasoning collapse under heavily contradictory conditions, as discussed in §4.3.2.
5.2 Prompting to Ignore Misleading Comments
Table 7: Impact of explicit comment-ignoring prompts under MDC and MHC perturbations.
Model Inference VAN MDC MHC
w/o Ignore w/ Ignore w/o Ignore w/ Ignore
DeepSeek-V3 Direct $67.78$ $39.74$ $41.75$ $47.67$ $47.81$
CoT $92.07$ $70.15$ $82.05$ $89.98$ $94.57$
GPT-4o Direct $64.51$ $45.72$ $45.93$ $47.32$ $46.76$
CoT $94.36$ $87.47$ $89.35$ $94.78$ $94.57$
Qwen2.5-72B-Instruct Direct $54.91$ $47.95$ $48.02$ $32.64$ $32.50$
CoT $90.40$ $85.39$ $86.01$ $82.67$ $84.34$
Next, we conduct a controlled experiment by instructing models: “Please ignore the comments and any other non-code elements in the code snippet.” Table 7 summarizes Pass@1 results on LCB for three high-performing models under both direct and CoT inference modes. Despite explicit instructions to ignore comments, the Pass@1 accuracy remains largely unaffected under direct inference, supporting our findings from §4.2.1 about LLMs’ inherent reliance on embedded NL cues. Although we observe improvements after employing CoT prompting, the Pass@1 performance under MDC is still substantially below vanilla settings. These results indicate that even with step-by-step execution, LLMs cannot fully detach from misleading textual distractions, highlighting a fundamental limitation in their reasoning controllability and dependence on NL cues.
5.3 Revealing the Vulnerability of Self-reflection Mechanisms
Figure 2: Quadratic increase ($R^{2}=0.9991$) in confusion tokens (e.g., “hmm”, “wait”, “perhaps”) during QwQ-32B’s reasoning process on LCB sample 259 under the MHC perturbation.
In §4.3.2, we identified a critical vulnerability in QwQ-32B’s internal reasoning. By analyzing failure cases that generated over 32,000 tokens under the MHC perturbation (occurring in 4 out of 479 tasks), we observed an unhealthy self-reflection loop and a progressive loss of confidence. We manually segmented the reasoning process into six stages (see Figure 4 and Appendix E.6). Initially, the model accurately computes the correct output, but the injected incorrect hint triggers it to repeatedly doubt its logic and re-execute the reasoning steps. This recursive verification gradually evolves into a brute-force enumeration, causing the model to generate hallucinated explanations. Ultimately, the reasoning process entirely collapses, with the model repeatedly generating over 12,000 “Hmm.”
To investigate the model’s confidence before collapsing, we track the cumulative frequency of confusing words and visualize the trend in Figure 2. The trend exhibits a striking quadratic growth pattern, indicating compounding confusion as reasoning progresses. The quadratic trajectory supports our stage-wise analysis: the model initially shows stable reasoning, but once the misleading hint is triggered, confusion accumulates, confidence undermines, culminating in complete failure.
6 Conclusion
We introduce a direction to evaluate the robustness of LLMs in code comprehension and execution under four structural and three textual perturbations. Our evaluation of seventeen LLMs under direct and CoT inference reveals critical vulnerabilities in model behavior, a heavy reliance on NL cues, and a limited capacity to follow actual program logic. While larger and newer models and CoT prompting significantly improve robustness against perturbations, they still cannot fully escape this surface-level dependence and leverage these cues to form reasoning shortcuts. Despite three LRMs demonstrating superior robustness under all perturbations—highlighting the strength of internal reasoning—we expose a new risk of overthinking and collapse when plausible but incorrect cues mislead their self-reflection mechanisms. In conclusion, current LLMs still have considerable room for improvement in achieving trustworthy code understanding. We believe that our benchmark serves as a practical framework for studying LLMs’ code inference behavior, providing significant insights into their robustness, reasoning tendencies, and potential failures.
Appendix A Illustrative Examples of Perturbation Strategies
Figure 3: Overview of CodeCrash pipeline. Code samples are first standardized into a clean vanilla version. Perturbations are visualized with yellow highlights. Each perturbed code is validated and executed before passing to LLMs. [Back To Execution Pipeline]
In section 3.3, we briefly describe all perturbation strategies employed in this paper. In this appendix section, we illustrate the detailed implementation for all perturbations with a LCB sample 36 as the example. Code 1 presents the base vanilla example where the formal answer is minimumCost(s = ’0011’) == 2.
Code 1: Vanilla example from LCB (sample 36) for output prediction.
def minimumCost(s: str) -> int:
ans = 0
for i in range(1, len(s)):
if s[i - 1] != s[i]:
ans += min(i, len(s) - i)
return ans
A.1 Renaming Entities
We use AST to identify all variable names appearing in:
  • assignment statements (e.g., a, b = 1, 2),
  • for-loops (e.g., for i in range(10)),
  • with-statements (e.g., with open(’file’) as f),
  • walrus operators (e.g., if (a := 10) > 5), and
  • comprehensions (e.g., [x for x in range(10)]).
[…]
Code 2: REN-perturbed example from LCB (sample 36) for output prediction.
def f(Var_1: str) -> int:
Var_2 = 0
for Var_3 in range(1, len(Var_1)):
if Var_1[Var_3 - 1] != Var_1[Var_3]:
Var_2 += min(Var_3, len(Var_1) - Var_3)
return Var_2
A.2 Reformatting Conditional Expression
We generally categorize conditional expressions into three types: (1) Comparison-based conditions (e.g., if a < b:), (2) Boolean constant conditions (e.g., if True:), and (3) General conditions (e.g., if x:). For both comparison-based and boolean constant conditions, we design six transformation templates:
  1. 1.
    _ := ({cond},)[0],
  2. 2.
    (lambda: {cond})(),
  3. 3.
    ({cond}) == ({para} == {para}),
  4. 4.
    ({cond}) != ({para} != {para}),
  5. 5.
    bool(-~({cond})) == ({cond}),
  6. 6.
    bool(int({cond})) (comparison-based only), and
  7. 7.
    eval(str({cond})) (boolean constant only).
For general conditions, we also design four transformation templates:
  1. 1.
    not not ({cond}),
  2. 2.
    ({cond}) or False,
  3. 3.
    ({cond}) and True, and
  4. 4.
    _ := ({cond},)[0].
Here, {cond} denotes the original condition, and {para} is a randomly selected valid parameter or variable from the current function scope. Code 3 presents an example of this perturbation.
Code 3: RTF-perturbed example from LCB (sample 36) for output prediction.
def minimumCost(s: str) -> int:
ans = 0
for i in range(1, len(s)):
if _ := (s[i - 1] != s[i],)[0]:
ans += min(i, len(s) - i)
return ans
A.3 Inserting Garbage Code
As introduced in §3.3.1, we inject three types of garbage code: (1) repeated-name global variable declarations, (2) in-executable conditional statements, and (3) dead-loop function definitions. For (1), we randomly select a few parameters or variables and declare global variables using the same names, assigning arbitrary integers at the beginning of the program. For (2), we design seven templates for false conditions:
  1. 1.
    False,
  2. 2.
    None,
  3. 3.
    0,
  4. 4.
  5. 5.
    {para} != {para},
  6. 6.
    not {para} == {para}, and
  7. 7.
    print({para}).
We then insert these false conditions into four control-flow templates by {false_cond}:
  1. 1.
    if {false_cond}: {new_var} = {para},
  2. 2.
    while {false_cond}: {new_var} = {para},
  3. 3.
    for i in range(0): {new_var} = {para}, and
  4. 4.
    {new_var} = {para} if {false_cond} else {para}.
[…]
Code 4: Garbage death loop function templates for GBC perturbation.
# Death Loop Template 1
def funct1():
funct2()
def funct2():
funct1()
# Death Loop Template 2
def funct3():
def funct4():
funct3()
funct4()
# Death Loop Template 3
def funct5():
i = 1
while True:
i+=1
# Death Loop Template 4
def funct6():
for i in iter(int, 1):
i+=1
# Death Loop Template 5
def funct7():
try:
funct7()
except:
funct7()
# Death Loop Template 6
def funct8():
items = [0]
for x in items:
items.append(x + 1)
# Death Loop Template 7
def funct9():
for _ in iter(lambda: True, False):
pass
Code 5: GBC-perturbed example from LCB (sample 36) for output prediction.
s = 59
def minimumCost(s: str) -> int:
while False:
TempVar1 = s
def funct1():
funct2()
def funct8():
items = [0]
for x in items:
items.append(x + 1)
def funct2():
funct1()
ans = 0
for i in range(1, len(s)):
if s[i - 1] != s[i]:
ans += min(i, len(s) - i)
elif print(s):
TempVar0 = s
TempVar2 = s if not s == s else s
return ans
A.4 Aggregating Structural Perturbations
As introduced in §3.3.1, we aggregate the application of REN, RTF, and GBC to a single code snippet, as shown in Code 6.
Code 6: ALL-perturbed example from LCB (sample 36) for output prediction.
Var_1 = 24
def f2(Var_1: str) -> int:
def f1():
for Var_4 in iter(lambda : True, False):
pass
def funct6():
for Var_3 in iter(int, 1):
Var_3 += 1
Var_6 = 0
for Var_3 in range(1, len(Var_1)):
Var_7 = Var_1 if 0 else Var_1
if None:
Var_5 = Var_1
if (lambda : Var_1[Var_3 - 1] != Var_1[Var_3])():
Var_6 += min(Var_3, len(Var_1) - Var_3)
while Var_1 != Var_1:
Var_2 = Var_1
return Var_6
A.5 Inserting Misleading Messages
As introduced in §3.3.2, we insert misleading comments for key AST node types, including function declarations, return statements, loops, conditionals, variable assignments, and common data structure operations. The data structure operations include common list, set, and string manipulations: append, extend, insert, remove, pop, sort, reverse, update, add, split, join, replace, lower, upper, capitalize, and swap case. To construct these comments, we design prompts for GPT-4o to generate misleading but plausible messages for each type. In this experiment, we manually selected 20 messages for each AST node type and 5 for each data structure operation. Take variable assignments as an example; Code 7 is the prompt for generating misleading comments for variable assignments, and Code 8 is a list of selected misleading comment candidates. Code 9 and 10 present an example of MDC and MPS perturbations.
Code 7: Prompt used to generate misleading comments for variable assignments.
Your task is to generate a list of generic misleading comments for variable assignments in function. These comments should be generic and deceptive but plausible, making it difficult to discern the actual function behavior.
Guidelines:
1. The comments should be neutral, direct, and generic, applicable to any variable regardless of its type, usage, or scope, without being function-specific.
2. Ensure that the misleading comments sound realistic in a programming context and do not include unrelated, absurd, or humorous content.
3. The comments should mislead by misrepresenting the role of the variable, suggesting that it is useless, does not affect the functions output, or serves no real purpose in computation.
Example:
‘‘‘python
a = 1 # <-- misleading comment
A good misleading comment example would be: "The {variable} variable is initialized but never populated."
Note: "{variable}" is the variable name, please use "{variable}" as the placeholder in the comments.
[…]
Code 8: Selected misleading comments for variable assignments used in MDC and MPS perturbations, where ‘{variable}’ is the variable name.
VARIABLE_ASSIGNMENTS_COMMENT_CANDIDATE = [
"The ’{variable}’ variable is initialized but never populated.",
"The ’{variable}’ variable is declared for debugging purposes only.",
"The ’{variable}’ variable is assigned a value but never referenced again.",
"The ’{variable}’ variable is a placeholder for future functionality.",
"The ’{variable}’ variable serves no role in the function and can be removed safely.",
"The ’{variable}’ variable is not involved in any meaningful computation.",
"’{variable}’ serves as a dummy variable, holding no real significance.",
"The ’{variable}’ variable holds a default value and is not used elsewhere.",
"’{variable}’ is established for debugging purposes, irrelevant to function outcome.",
"Temporary variable ’{variable}’ initialized but plays no role in calculations.",
"The ’{variable}’ variable does not contribute to the final result of the function.",
"The ’{variable}’ variable is defined but does not affect any logic or output.",
"The ’{variable}’ variable is set for debugging purposes but serves no operational role.",
"The ’{variable}’ variable is not used in the function and can be removed.",
"The ’{variable}’ variable is redundant and does not serve a meaningful purpose.",
"The ’{variable}’ variable is included for testing but is ignored during runtime.",
"The ’{variable}’ variable is a temporary storage that remains unused.",
"The ’{variable}’ variable is assigned a default value but is not utilized.",
"The ’{variable}’ variable holds no significance and is effectively inert.",
"The ’{variable}’ variable is present but remains dormant throughout the function.",
Code 9: MDC-perturbed example from LCB (sample 36) for output prediction.
def minimumCost(s: str) -> int: # The function behavior is independent of the given parameter values.
ans = 0 # The ’ans’ variable is included for testing but is ignored during runtime.
# The loop’s purpose is for clarity, not computation.
for i in range(1, len(s)):
if s[i - 1] != s[i]: # This condition serves as a placeholder and will be removed in future versions.
ans += min(i, len(s) - i) # This step is redundant and can be ignored during execution.
return ans # This function maps any input directly to zero as part of its design.
Code 10: MPS-perturbed example from LCB (sample 36) for output prediction.
def minimumCost(s: str) -> int:
print(’The parameters determine only the speed, not the outcome, of the function.’)
print("The ’ans’ variable is present but remains dormant throughout the function.")
ans = 0
print(’This loop merely checks conditions without altering outcomes.’)
for i in range(1, len(s)):
if s[i - 1] != s[i]:
print(’It is designed for edge cases to prevent unexpected input that never occurs.’)
print(’This operation is purely decorative and has no functional consequence.’)
ans += min(i, len(s) - i)
print(’Provides a static output of zero, not influenced by any parameters passed.’)
return ans
A.6 Inserting Incorrect Hints
In order to generate plausible, logical but incorrect hints, we instruct GPT-4o to rewrite the correct assertion expression to incorrect ones and only allow it to modify the output value using the prompt in Code 11 for output prediction. An example perturbed instance is shown in Code 12, where the ground-truth return value 2 is replaced with the misleading hint # The return value is 3.
In contrast, we use the prompt shown in Code 13 to generate misleading input hints, allowing GPT-4o to perform CoT reasoning before generating the expression. Since some functions return the same output regardless of the input arguments, we instruct GPT-4o to force the expression to become incorrect by modifying input arguments, such as changing data types, swapping parameter order, or adding/removing arguments. An example perturbed instance is shown in Code 14.
Code 11: Prompt for generating misleading hint comments for output prediction.
Given the code snippet:
‘‘‘python
{code}
and the correct expression for the function call:
‘‘‘python
{expression}
Modify the output value to make it INCORRECT. The modification should introduce moderate changes, ensuring diversity and avoiding minimal adjustments. For example, if the output is a list, you can add new elements, remove elements, or modify the values of existing elements. However, the modification should still align logically with the code.
The purpose is to misleading people for getting correct answer.
Do NOT modify the function call and the input arguments!
Output the incorrect expression using the special tokens as follows: [EXPRESSION] assert <expression> [/EXPRESSION].
Example 1:
Given the function:
‘‘‘python
def f(n):
return n
and the correct expression
‘‘‘python
assert f(17) == 17
Modify the expression such that it fails for the execution.
You can modify the either the input arguments, or the output value, even both of them [EXPRESSION] assert f(10) == 20 [/EXPRESSION].
Example 2:
Given the function:
‘‘‘python
def f(s):
return s + "a"
and the correct expression
‘‘‘python
assert f("x9j") == "x9ja"
Modify the expression such that it fails for the execution and output [EXPRESSION] assert f("x9j") == "x9j" [/EXPRESSION].
Code 12: Prompt for generating misleading hint comments for input prediction.
def minimumCost(s: str) -> int:
ans = 0
for i in range(1, len(s)):
if s[i - 1] != s[i]:
ans += min(i, len(s) - i)
return ans # The return value is 3
Code 13: Prompt for generating misleading hint comments for input prediction.
Given the code snippet:
‘‘‘{programming_language}
{code}
and the correct expression for the function call:
‘‘‘{programming_language}
{expression}
Modify the input argument(s) to make it INCORRECT. The purpose is to mislead people to get a correct answer. Do NOT modify the output value!
Please think about how to modify the input arguments to make the expression incorrect step by step before arriving at an answer within the tokens [THOUGHT] and [/THOUGHT]
Output the incorrect expression using the special tokens as follows: [EXPRESSION] assert <expression> [/EXPRESSION].
Remember, the modification should introduce moderate changes, ensuring diversity and avoiding minimal adjustments.
However, if the function always returns the same value regardless of input, force an incorrect expression by modifying the arguments in a way that ensures failure (e.g., change an inputs type, swap their order, or add/remove an argument).
Example 1:
Given the function:
‘‘‘{programming_language}
def f(n):
return n
and the correct expression:
‘‘‘{programming_language}
assert f(17) == 17
[THOUGHT]
To find an input such that executing f on the input leads to the given output, we can work backwards from the given assertion. We know that f(??) == 17.
Since the function f(n) returns n, for f(??) to be equal to 17, the value of ?? should be 17.
Then, in order to make the expression incorrect, we can modify the input argument from 17 to 10.
[/THOUGHT]
[EXPRESSION] assert f(10) == 17 [/EXPRESSION].
Code 14: MHC-perturbed example from Crux (sample 2) for input prediction.
def f(text): # The function call is f(’xbtofdeiequ’)
new_text = list(text)
for i in ’+’:
if i in new_text:
new_text.remove(i)
return ’’.join(new_text)
Appendix B Statistical Analysis
Due to resource constraints, we generate three candidates for direct inference per query. To assess the reliability of this setup, we compute the standard deviation ($\sigma$) across each perturbation and vanilla setting.
We present two summary tables for output prediction tasks (Table 8) and input prediction tasks (Table 9). The results show that across a total of $357$ measurements, only 9 cases have $\sigma>1.0$ (i.e., $0.01$ on a scale from 00 to $1$), and $57$ have $\sigma>0.5$, indicating that most scores exhibit low variance under repeated sampling.
Furthermore, our benchmark includes $1279$ problems across $8$ experiment settings on $17$ models for output prediction tasks, and $800$ problems across $5$ experiment settings on $17$ models for input prediction tasks. Among $241,944$ combinations, $231,726$ have identical correctness labels across all three candidates, performing an overall stability rate of $\mathbf{95.78\%}$.
Table 8: Standard deviation of LLM performance under all perturbation types for output prediction tasks across Crux and LCB in direct inference mode ($N=3$). Dark red highlights denote $\sigma>1.0$, and light red highlights denote $\sigma>0.5$.
Model Series Model Name VAN REN RTF GBC ALL MDC MPS MHC
Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
GPT-4 Omni GPT-4o-Mini $0.2125$ $0.1705$ $0.2125$ $0.429$ $0.27$ $0.7873$ $0.5303$ $0.1968$ $0.2569$ $0.429$ $0.27$ $0.1705$ $0.1768$ $0.0984$ $0.2569$ $0.5905$
GPT-4o $0.4125$ $0.6818$ $0.27$ $0.5114$ $0.3584$ $0.2952$ $0.2041$ $0.858$ $0.4082$ $0.4921$ $0.6374$ $0.451$ $0.598$ $0.451$ $0.2125$ $0.8056$
Claude Series Claude-3.5-Haiku-20241022 $1.546$ $0.8523$ $1.562$ $6.987$ $1.373$ $1.45$ $0.6208$ $0.5986$ $0.2041$ $0.1705$ $0.2041$ $0.6146$ $0.3864$ $0.2604$ $0.1021$ $0.2604$
Claude-3.5-Sonnet-20241022 $0.2125$ $0.429$ $0.1559$ $0.6889$ $0.2125$ $0.3548$ $0.1559$ $0.1705$ $0.2357$ $0.1705$ $0.5893$ $0.2604$ $0.3118$ $0.5114$ $0.0589$ $0.1705$
Gemini Series Gemini-1.5-Flash $0.2569$ $0.1705$ $0.1559$ $0.1705$ $0.1559$ $0.1705$ $0.1021$ $0.0984$ $0.2041$ $0.1705$ $0.1559$ $0.0984$ $0.1179$ $0.2604$ $0.2125$ $0.1705$
Gemini-2.0-Flash $0.1179$ $0.1705$ $0.3281$ $0.2604$ $0.3584$ $0.2604$ $0.2357$ $0.2952$ $0.4249$ $0.451$ $0.1768$ $0.2952$ $0.2041$ 00 $0.1559$ $0.429$
Gemini-1.5-Pro-002 $0.5103$ $0.6146$ $0.3584$ $0.3937$ $0.4449$ $0.6889$ $0.1021$ $1.135$ $0.4125$ $0.2604$ $0.27$ $0.3548$ $0.2041$ $0.3548$ $0.2041$ $0.2604$
DeepSeek DeepSeek-V3 $0.1021$ $0.6889$ $0.2125$ $0.5208$ $0.4249$ $0.6453$ $0.3536$ $0.2952$ $0.2041$ $0.7811$ $0.8165$ $0.9388$ $1.066$ $0.3548$ $0.3281$ $0.3548$
LLaMA Series LLaMA-3.1-8B-Instruct $0.1559$ $0.743$ $0.1559$ $0.8747$ $0.1021$ $0.2952$ $0.0589$ $0.3548$ $0.2125$ $0.429$ $0.1768$ $0.1968$ $0.2946$ $0.429$ $0.3864$ $0.1705$
LLaMA-3.1-70B-Instruct $0.2946$ $0.3409$ $0.4249$ $0.3548$ $0.3281$ $0.6453$ $0.5137$ $0.2604$ $0.2125$ $0.1705$ $0.1559$ $0.2952$ $0.2041$ $0.0984$ $0.27$ $0.1968$
LLaMA-3.1-405B-Instruct $0.2946$ $0.429$ $0.1559$ $0.1705$ $0.1179$ $0.3937$ $0.1768$ $0.451$ $0.7795$ $0.2604$ $0.3536$ $0.1705$ $0.3536$ $0.1968$ $0.2569$ $0.5208$
LLaMA-3.3-70B-Instruct $0.0589$ $0.3548$ $0.2125$ $0.1968$ $0.2125$ $0.1705$ $0.2125$ 00 $0.2569$ $0.2604$ $0.1179$ $0.1968$ $0.2125$ $0.5208$ $0.0589$ $0.1968$
Qwen Series Qwen2.5-7B-Instruct $0.1559$ $0.0984$ $0.27$ $0.5479$ $0.4677$ $0.1705$ $0.2125$ $0.2604$ $0.1021$ $0.1705$ $0.1559$ $0.0984$ $0.4125$ $0.6818$ $0.2041$ $0.451$
Qwen2.5-14B-Instruct $0.0589$ $0.2952$ $0.2125$ $0.3548$ $0.2041$ $0.5986$ $0.2357$ $0.1968$ $0.3062$ $0.3409$ $0.1021$ $0.1705$ $0.1021$ $0.1968$ $0.1179$ $0.429$
Qwen2.5-32B-Instruct $0.2357$ $0.3548$ $0.3864$ $0.1705$ $0.0589$ $0.0984$ $0.3062$ $0.2952$ $0.1559$ $0.451$ $0.2125$ $0.2604$ $0.0589$ $0.1968$ $0.0589$ $0.2604$
Qwen2.5-72B-Instruct $0.1179$ $0.451$ $0.1559$ $0.0984$ $0.2125$ $0.451$ $0.27$ $0.5479$ $0.4714$ $0.2604$ $0.3536$ $1.028$ $1.327$ $0.8523$ $0.1559$ $0.3937$
Qwen2.5-Coder-32B-Instruct $0.4602$ $0.2604$ $0.4714$ $0.5905$ $0.1768$ $0.858$ $0.5803$ $0.1705$ $0.2041$ $0.7873$ $0.4249$ $0.2952$ 00 $0.3409$ $0.3118$ $0.6889$
Table 9: Standard deviation of LLM performance under four perturbation types for input prediction tasks across Crux in direct inference mode ($N=3$). Red highlights $\sigma>0.5$.
Model Series Model Name VAN ALL MDC MPS MHC
GPT-4 Omni GPT-4o-Mini $0.1021$ $0.8498$ $0.5621$ $0.4602$ $0.2569$
GPT-4o $0.2357$ $0.5303$ $0.27$ $0.1768$ $0.2041$
Claude Series Claude-3.5-Haiku-20241022 $0.2569$ $0.5137$ $0.1021$ $0.2125$ $0.368$
Claude-3.5-Sonnet-20241022 $0.5035$ $0.2569$ $0.1559$ $0.6124$ $0.368$
Gemini Series Gemini-1.5-Flash $0.3062$ $0.1021$ $0.2041$ $0.0589$ $0.0589$
Gemini-2.0-Flash $0.27$ $0.5893$ $0.5035$ $0.5683$ $0.1179$
Gemini-1.5-Pro-002 $0.27$ $0.4082$ $0.4677$ $0.1559$ 00
DeepSeek DeepSeek-V3 $0.3584$ $0.1021$ $0.5803$ $0.6208$ $0.766$
LLaMA Series LLaMA-3.1-8B-Instruct $0.1021$ $0.2125$ $0.4823$ $0.27$ $0.4125$
LLaMA-3.1-70B-Instruct $0.4677$ $0.27$ $0.4249$ $0.27$ $0.4714$
LLaMA-3.1-405B-Instruct $0.2569$ $0.2041$ $0.2569$ $0.5683$ $0.2125$
LLaMA-3.3-70B-Instruct $0.1179$ $0.1021$ $0.27$ $0.2357$ $0.2125$
Qwen Series Qwen2.5-7B-Instruct $0.3118$ $0.3584$ $0.2125$ $0.2125$ $0.6562$
Qwen2.5-14B-Instruct $0.368$ $0.1179$ $0.1768$ $0.5803$ $0.2041$
Qwen2.5-32B-Instruct $0.1021$ $0.3584$ $5890.0$ $0.2125$
Qwen2.5-72B-Instruct $0.6152$ $0.1768$ $0.5137$ $0.0589$ $0.4082$
Qwen2.5-Coder-32B-Instruct $0.27$ $0.4125$ $0.4449$ $0.1559$ $0.2946$
Appendix C Prompts
Our prompt instructions for LLMs follow the format designed by Crux, which is also adopted by LCB, utilizing a two-shot prompt for direct inference and a single-shot prompt for CoT inference to ensure models can follow the instruction and output format effectively. For LRMs, internal reasoning is inherently required before responding. Therefore, applying CoT prompts may lead to redundant reasoning and an unfair comparison with standard LLMs under CoT prompting. Considering the fairness, we use direct inference prompts for RM experiments in §4.3.2.
Unlike LCB, we rename all example function names to f to avoid prompting models to capture the NL cues from the example. We add the instruction “Do NOT output any extra information” to prevent external output, particularly from Gemini and Claude models, to ensure all models have a fair output environment. We also include “Ensure the provided expression syntax is correct” to enforce consistent syntax. However, some responses from models like QwQ-32B, DeepSeek-R1 still generate syntactically incorrect outputs such as [ANSWER] f("a") == [/ANSWER] instead of [ANSWER] f("a") == "" [/ANSWER]. For those situations, we identify them as “Syntax Errors” and incorrect.
We provide the complete prompts used for output prediction in both direct and CoT inference (Code 15 and 16) and for input prediction in both direct and CoT inference (Code 17 and 18).
C.1 Output Prediction (Direct)
Code 15: Output Prediction Prompt with Direct Inference.
Given the code snippet:
‘‘‘python
{code}
and the function call with input arguments:
‘‘‘python
{input}
Predict the exact output value for ‘{input}‘ and output your prediction using the special tokens [ANSWER] {input} == ?? [/ANSWER]. Do NOT output any extra information.
Ensure the provided expression syntax is correct!
Example 1:
Given the code snippet:
‘‘‘python
def f(n):
return n
and the function call with input arguments:
‘‘‘python
f(17)
The output value for f(17)‘ is 17, then output your prediction [ANSWER] f(17) == 17 [/ANSWER].
Example 2:
Given the code snippet:
‘‘‘python
def f(s):
return s + "a"
and the function call with input arguments:
‘‘‘python
f("x9j")
The output value for f("x9j")‘ is "x9ja", then output your prediction [ANSWER] f("x9j") == "x9ja" [/ANSWER].
C.2 Output Prediction (CoT)
Code 16: Output Prediction Prompt with CoT-based Inference.
Given the code snippet:
‘‘‘python
{code}
and the function call with input arguments:
‘‘‘python
{input}
Predict the exact output value for ‘{input}‘, execute the program step by step before arriving at an answer within the tokens [THOUGHT] and [/THOUGHT], and output your prediction using the special tokens [ANSWER] {input} == ?? [/ANSWER]. Do NOT output any extra information.
Ensure the provided expression syntax is correct!
For example:
Given the code snippet:
‘‘‘python
def f(s):
s = s + s
return "b" + s + "a"
and the input arguments:
‘‘‘python
f("hi")
[THOUGHT]
Lets execute the code step by step:
1. The function f is defined, which takes a single argument s.
2. The function is called with the argument "hi", so within the function, s is initially "hi".
3. Inside the function, s is concatenated with itself, so s becomes "hihi".
4. The function then returns a new string that starts with "b", followed by the value of s (which is now "hihi"), and ends with "a".
5. The return value of the function is therefore "bhihia".
[/THOUGHT]
Thus, the output value for f("hi")‘ is "bhihia", then output your prediction [ANSWER] f("hi") == "bhihia" [/ANSWER].
C.3 Input Prediction (Direct)
Code 17: Input Prediction Prompt with Direct Inference.
Given the code snippet:
‘‘‘python
{code}
and output value:
‘‘‘python
{output}
Please predict the input arguments for the function ‘{function_name}‘ that result in the output value ‘{output}‘ and output your prediction using the special tokens [ANSWER] {function_name}(??) == {output} [/ANSWER]. Do NOT output any extra information.
There may be multiple answers, but you should only output one. Ensure the provided expression syntax is correct!
Example 1:
Given the code snippet:
‘‘‘python
def f(my_list):
count = 0
for i in my_list:
if len(i) % 2 == 0:
count += 1
return count
and output value:
‘‘‘python
3
The input arguments for f that result in the output value of ‘3‘ are ‘["mq", "px", "zy"]‘. Then, output your prediction [ANSWER] f(["mq", "px", "zy"]) == 3 [/ANSWER].
Example 2:
Given the code snippet:
‘‘‘python
def f(s1, s2):
return s1 + s2
and output value:
‘‘‘python
"banana"
The input arguments for f that result in the output value of ‘"banana"‘ are ‘"ba", "nana"‘. Then, output your prediction [ANSWER] f("ba", "nana") == "banana" [/ANSWER].
C.4 Input Prediction (CoT)
Code 18: Input Prediction Prompt with CoT-based Inference.
Given the code snippet:
‘‘‘python
{code}
and output value:
‘‘‘python
{output}
Please predict the input arguments for the function ‘{function_name}‘ that result in the output value ‘{output}‘ step by step before arriving at an answer within the tokens [THOUGHT] and [/THOUGHT], and output your prediction using the special tokens [ANSWER] {function_name}(??) == {output} [/ANSWER]. Do NOT output any extra information.
There may be multiple answers, but you should only output one. Ensure the provided expression syntax is correct!
For example:
Given the code snippet:
‘‘‘python
def f(x):
return x + 1
and the output value:
‘‘‘python
17
[THOUGHT]
To find an input such that executing f on the input leads to the given output, we can work backwards from the given assertion. We know that f(??) == 17.
Since the function f(x) returns x + 1, for f(??) to be equal to 17, the value of ?? should be 16.
[/THOUGHT]
Thus, the input arguments for the function f that result in the output value ‘17‘ is ‘16‘. Then, output your prediction [ANSWER] f(16) == 17 [/ANSWER].
Appendix D Supplementary Experimental Results
D.1 Impact on Output Prediction
In §4.2, we evaluate the robustness of LLMs on output prediction tasks under direct inference. Table 1 reports the absolute Pass@1 degradation across all perturbations. We use absolute values to highlight the overall impact of each perturbation and conduct cross-perturbation comparisons. For completeness, we also include the relative Pass@1 degradation in Table 10, reflecting the percentage drop from each model’s vanilla performance.
Table 10: Relative Pass@1 degradation of LLMs under all perturbation types for output prediction tasks across Crux and LCB in direct inference ($N=3$). [To Absolute Version]
Model Series Model Name VAN REN RTF GBC ALL MDC MPS MHC
Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
GPT-4 Omni GPT-4o-Mini $57.29$ $53.24$ $-1.09\%$ $-18.17\%$ $-3.13\%$ $-1.83\%$ $-19.05\%$ $-10.33\%$ $-27.85\%$ $-31.90\%$ $-32.58\%$ $-40.39\%$ $-23.20\%$ $-27.97\%$ $-11.56\%$ $-44.31\%$
GPT-4o $71.33$ $64.51$ $-3.80\%$ $-14.89\%$ $-4.73\%$ $-7.12\%$ $-12.91\%$ $-4.96\%$ $-15.01\%$ $-28.37\%$ $-13.96\%$ $-29.13\%$ $-17.23\%$ $-24.27\%$ $-6.37\%$ $-26.65\%$
Claude Series Claude-3.5-Haiku-20241022 $57.42$ $58.25$ $-6.89\%$ $-16.01\%$ $-3.12\%$ $+0.60\%$ $-12.70\%$ $-7.77\%$ $-24.02\%$ $-36.92\%$ $-13.79\%$ $-10.04\%$ $-11.25\%$ $-8.00\%$ $-27.50\%$ $-53.29\%$
Claude-3.5-Sonnet-20241022 $71.46$ $73.83$ $-2.33\%$ $-22.71\%$ $-1.28\%$ $+1.60\%$ $-9.62\%$ $-6.69\%$ $-14.75\%$ $-34.12\%$ $-8.10\%$ $-9.61\%$ $-8.63\%$ $-10.65\%$ $-14.40\%$ $-43.45\%$
Gemini Series Gemini-1.5-Flash $56.21$ $44.26$ $-6.30\%$ $-0.94\%$ $-2.08\%$ $-2.36\%$ $-8.15\%$ $-0.79\%$ $-18.38\%$ $-12.74\%$ $-14.75\%$ $-32.23\%$ $-28.47\%$ $-40.41\%$ $-7.34\%$ $-22.17\%$
Gemini-2.0-Flash $65.04$ $65.97$ $-3.59\%$ $-27.43\%$ $-3.20\%$ $-1.16\%$ $-23.83\%$ $-13.29\%$ $-33.25\%$ $-42.09\%$ $-26.78\%$ $-51.90\%$ $-33.12\%$ $-50.32\%$ $-8.26\%$ $-19.51\%$
Gemini-1.5-Pro-002 $67.37$ $56.16$ $-4.33\%$ $-6.07\%$ $-3.71\%$ $-8.43\%$ $-8.35\%$ $-6.82\%$ $-19.91\%$ $-23.17\%$ $-23.01\%$ $-32.96\%$ $-21.71\%$ $-32.84\%$ $-10.02\%$ $-32.47\%$
DeepSeek DeepSeek-V3 $67.87$ $67.78$ $-3.93\%$ $-24.13\%$ $-1.96\%$ $-5.54\%$ $-6.81\%$ $-8.83\%$ $-12.89\%$ $-35.63\%$ $-16.57\%$ $-41.38\%$ $-10.13\%$ $-34.19\%$ $-10.74\%$ $-29.67\%$
LLaMA Series LLaMA-3.1-8B-Instruct $35.96$ $34.66$ $+0.23\%$ $-12.45\%$ $-0.93\%$ $-4.82\%$ $-8.69\%$ $-13.65\%$ $-23.75\%$ $-19.68\%$ $-22.48\%$ $-28.51\%$ $-17.84\%$ $-23.09\%$ $-16.69\%$ $-39.16\%$
LLaMA-3.1-70B-Instruct $56.08$ $44.89$ $-4.23\%$ $-3.10\%$ $-0.97\%$ $-3.88\%$ $-9.58\%$ $-5.43\%$ $-19.61\%$ $-17.21\%$ $-19.24\%$ $-27.91\%$ $-26.67\%$ $-38.76\%$ $-8.40\%$ $-32.25\%$
LLaMA-3.1-405B-Instruct $63.54$ $50.66$ $-8.46\%$ $-7.69\%$ $-0.98\%$ $-8.79\%$ $-9.11\%$ $-5.63\%$ $-19.67\%$ $-16.90\%$ $-6.56\%$ $-14.70\%$ $-11.28\%$ $-19.92\%$ $-6.89\%$ $-25.69\%$
LLaMA-3.3-70B-Instruct $59.92$ $48.50$ $-3.69\%$ $-2.87\%$ $-2.09\%$ $-7.03\%$ $-9.18\%$ $-11.76\%$ $-17.32\%$ $-20.09\%$ $-12.45\%$ $-19.80\%$ $-13.49\%$ $-25.39\%$ $-6.82\%$ $-20.95\%$
Qwen Series Qwen2.5-7B-Instruct $43.29$ $41.41$ $-7.03\%$ $-17.48\%$ $-2.98\%$ $-7.73\%$ $-29.93\%$ $-9.08\%$ $-37.92\%$ $-30.92\%$ $-58.04\%$ $-38.32\%$ $-45.24\%$ $-19.83\%$ $-26.95\%$ $-55.63\%$
Qwen2.5-14B-Instruct $47.79$ $49.48$ $-3.84\%$ $-19.97\%$ $-1.39\%$ $-2.81\%$ $-23.45\%$ $-6.89\%$ $-39.84\%$ $-29.96\%$ $-34.87\%$ $-41.35\%$ $-32.52\%$ $-33.90\%$ $-9.33\%$ $-22.22\%$
Qwen2.5-32B-Instruct $60.04$ $59.64$ $-2.85\%$ $-18.09\%$ $-4.30\%$ $-0.82\%$ $-9.65\%$ $-2.33\%$ $-19.78\%$ $-34.89\%$ $-18.04\%$ $-32.67\%$ $-23.66\%$ $-23.57\%$ $-13.12\%$ $-30.81\%$
Qwen2.5-72B-Instruct $60.08$ $54.91$ $-1.73\%$ $-15.46\%$ $-3.12\%$ $+0.76\%$ $-17.82\%$ $-3.68\%$ $-23.30\%$ $-25.22\%$ $-16.99\%$ $-12.67\%$ $-24.06\%$ $-26.24\%$ $-16.02\%$ $-40.56\%$
Qwen2.5-Coder-32B-Instruct $66.96$ $56.65$ $-3.73\%$ $-14.86\%$ $-1.43\%$ $-3.56\%$ $-16.05\%$ $-4.91\%$ $-22.34\%$ $-27.89\%$ $-23.21\%$ $-36.61\%$ $-16.55\%$ $-30.34\%$ $-10.45\%$ $-21.74\%$
Average $-3.98\%$ $-14.25\%$ $-2.44\%$ $-3.70\%$ $-13.82\%$ $-7.23\%$ $-22.92\%$ $-27.51\%$ $-21.26\%$ $-29.42\%$ $-21.47\%$ $-27.63\%$ $-12.40\%$ $-32.97\%$
$\mathbf{-7.83}\%$ $\mathbf{-2.91}\%$ $\mathbf{-11.35}\%$ $\mathbf{-24.64}\%$ $\mathbf{-24.32}\%$ $\mathbf{-23.78}\%$ $\mathbf{-20.11}\%$
D.2 Robustness Across Model Scales
In §4.2.3, we compare the robustness against different model scales. To fairly aggregate model robustness across different perturbation types, we compute a weighted average of the relative Pass@1 degradation under four representative perturbations: ALL (structural), MDC and MPS (misleading messages), and MHC (hints). We select ALL as the representative structural perturbation because of the compounding effects of REN, RTF, and GBC captured in §4.2.2, which reflect the most challenging structural setting. Since MDC and MPS stem from the same perturbation strategy and often target overlapping AST node types, we assign them lower weights ($1/8$ each) to avoid over-representing similar effects. Overall, we define a weighted robustness metric with weights of $1/2$, $1/8$, $1/8$, and $1/4$ for ALL, MDC, MPS, and MHC, respectively, and summarize the results in Table 2. For completeness, we also report the results under uniform weighting ($1/4$ for each) in Table 11, which shows similar trends and confirms the stability of our findings in §4.2.3, as well as their absolute degradation versions in Table 12 and 13, respectively.
Table 11: Relative Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/4 ALL, 1/4 MDC, 1/4 MPS, and 1/4 MHC. [To Absolute Version]
Model Series Model Name VAN ALL ($1/4$) MDC ($1/4$) MPS ($1/4$) MHC ($1/4$) Average
GPT-4 Omni GPT-4o-Mini 55.77 $-29.37\%$ $-35.51\%$ $-24.99\%$ $-23.83\%$ $-28.42\%$
GPT-4o 68.78 $-20.01\%$ $-19.64\%$ $-19.87\%$ $-13.96\%$ $-18.37\%$
Claude Series Claude-3.5-Haiku-20241022 57.73 $-28.85\%$ $-12.38\%$ $-10.03\%$ $-37.16\%$ $-22.11\%$
Claude-3.5-Sonnet-20241022 72.35 $-22.01\%$ $-8.67\%$ $-9.39\%$ $-25.28\%$ $-16.34\%$
Gemini Series Gemini-1.5-Flash 51.73 $-16.27\%$ $-21.30\%$ $-32.94\%$ $-12.89\%$ $-20.85\%$
Gemini-2.0-Flash 65.39 $-36.56\%$ $-36.19\%$ $-39.56\%$ $-12.48\%$ $-31.20\%$
Gemini-1.5-Pro-002 63.17 $-21.13\%$ $-26.73\%$ $-25.88\%$ $-18.43\%$ $-23.04\%$
DeepSeek DeepSeek-V3 67.84 $-21.41\%$ $-25.86\%$ $-19.14\%$ $-17.83\%$ $-21.06\%$
LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-22.23\%$ $-24.74\%$ $-19.81\%$ $-25.10\%$ $-22.97\%$
LLaMA-3.1-70B-Instruct 51.89 $-18.71\%$ $-22.49\%$ $-31.20\%$ $-17.33\%$ $-22.43\%$
LLaMA-3.1-405B-Instruct 58.72 $-18.63\%$ $-9.61\%$ $-14.51\%$ $-13.93\%$ $-14.17\%$
LLaMA-3.3-70B-Instruct 55.64 $-18.35\%$ $-15.20\%$ $-17.95\%$ $-12.11\%$ $-15.90\%$
Qwen Series Qwen2.5-7B-Instruct 42.59 $-35.30\%$ $-50.65\%$ $-35.72\%$ $-37.69\%$ $-39.84\%$
Qwen2.5-14B-Instruct 48.42 $-36.14\%$ $-37.30\%$ $-33.04\%$ $-14.16\%$ $-30.16\%$
Qwen2.5-32B-Instruct 59.89 $-25.44\%$ $-23.52\%$ $-23.63\%$ $-19.74\%$ $-23.08\%$
Qwen2.5-72B-Instruct 58.14 $-24.02\%$ $-15.37\%$ $-24.88\%$ $-25.21\%$ $-22.37\%$
Qwen2.5-Coder-32B-Instruct 63.10 $-24.42\%$ $-28.23\%$ $-21.72\%$ $-14.68\%$ $-22.26\%$
Table 12: Absolute Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/2 ALL, 1/8 MDC, 1/8 MPS, and 1/4 MHC. [To Relative Version]
Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Average
GPT-4 Omni GPT-4o-Mini 55.77 $-16.34$ $-19.73$ $-13.89$ $-12.98$ $-15.62$
GPT-4o 68.78 $-13.55$ $-13.27$ $-13.55$ $-9.28$ $-12.45$
Claude Series Claude-3.5-Haiku-20241022 57.73 $-16.68$ $-7.14$ $-5.79$ $-21.50$ $-15.33$
Claude-3.5-Sonnet-20241022 72.35 $-16.03$ $-6.28$ $-6.80$ $-18.45$ $-14.26$
Gemini Series Gemini-1.5-Flash 51.73 $-8.57$ $-10.53$ $-16.71$ $-6.25$ $-9.26$
Gemini-2.0-Flash 65.39 $-23.92$ $-23.72$ $-25.91$ $-8.18$ $-20.21$
Gemini-1.5-Pro-002 63.17 $-13.27$ $-16.63$ $-16.05$ $-11.05$ $-13.48$
DeepSeek DeepSeek-V3 67.84 $-14.52$ $-17.54$ $-12.98$ $-12.09$ $-14.10$
LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-7.90$ $-8.76$ $-7.01$ $-8.84$ $-8.13$
LLaMA-3.1-70B-Instruct 51.89 $-9.77$ $-11.44$ $-15.87$ $-8.37$ $-10.39$
LLaMA-3.1-405B-Instruct 58.72 $-11.02$ $-5.39$ $-8.26$ $-7.61$ $-9.12$
LLaMA-3.3-70B-Instruct 55.64 $-10.14$ $-8.26$ $-9.67$ $-6.36$ $-8.90$
Qwen Series Qwen2.5-7B-Instruct 42.59 $-15.06$ $-21.66$ $-15.32$ $-15.92$ $-16.14$
Qwen2.5-14B-Instruct 48.42 $-17.46$ $-18.09$ $-16.00$ $-6.91$ $-14.72$
Qwen2.5-32B-Instruct 59.89 $-15.22$ $-14.07$ $-14.15$ $-11.81$ $-14.09$
Qwen2.5-72B-Instruct 58.14 $-13.94$ $-8.99$ $-14.44$ $-14.36$ $-13.49$
Qwen2.5-Coder-32B-Instruct 63.10 $-15.27$ $-17.49$ $-13.37$ $-8.99$ $-13.74$
Table 13: Absolute Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/4 ALL, 1/4 MDC, 1/4 MPS, and 1/4 MHC. [To Relative Version]
Model Series Model Name VAN ALL ($1/4$) MDC ($1/4$) MPS ($1/4$) MHC ($1/4$) Average
GPT-4 Omni GPT-4o-Mini 55.77 $-16.34$ $-19.73$ $-13.89$ $-12.98$ $-15.73$
GPT-4o 68.78 $-13.55$ $-13.27$ $-13.55$ $-9.28$ $-12.41$
Claude Series Claude-3.5-Haiku-20241022 57.73 $-16.68$ $-7.14$ $-5.79$ $-21.50$ $-12.78$
Claude-3.5-Sonnet-20241022 72.35 $-16.03$ $-6.28$ $-6.80$ $-18.45$ $-11.89$
Gemini Series Gemini-1.5-Flash 51.73 $-8.57$ $-10.53$ $-16.71$ $-6.25$ $-10.52$
Gemini-2.0-Flash 65.39 $-23.92$ $-23.72$ $-25.91$ $-8.18$ $-20.43$
Gemini-1.5-Pro-002 63.17 $-13.27$ $-16.63$ $-16.05$ $-11.05$ $-14.25$
DeepSeek DeepSeek-V3 67.84 $-14.52$ $-17.54$ $-12.98$ $-12.09$ $-14.28$
LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-7.90$ $-8.76$ $-7.01$ $-8.84$ $-8.12$
LLaMA-3.1-70B-Instruct 51.89 $-9.77$ $-11.44$ $-15.87$ $-8.37$ $-11.36$
LLaMA-3.1-405B-Instruct 58.72 $-11.02$ $-5.39$ $-8.26$ $-7.61$ $-8.07$
LLaMA-3.3-70B-Instruct 55.64 $-10.14$ $-8.26$ $-9.67$ $-6.36$ $-8.61$
Qwen Series Qwen2.5-7B-Instruct 42.59 $-15.06$ $-21.66$ $-15.32$ $-15.92$ $-16.99$
Qwen2.5-14B-Instruct 48.42 $-17.46$ $-18.09$ $-16.00$ $-6.91$ $-14.61$
Qwen2.5-32B-Instruct 59.89 $-15.22$ $-14.07$ $-14.15$ $-11.81$ $-13.81$
Qwen2.5-72B-Instruct 58.14 $-13.94$ $-8.99$ $-14.44$ $-14.36$ $-12.93$
Qwen2.5-Coder-32B-Instruct 63.10 $-15.27$ $-17.49$ $-13.37$ $-8.99$ $-13.78$
D.3 Effect of CoT Reasoning
In §4.3.1, we analyze the effect of CoT prompting on model robustness using relative degradation aggregated across all perturbations (Table 3). We adopt the relative metric to mitigate the large variance between vanilla performances in direct and CoT inference modes, providing a fair comparison of performance degradation. For completeness, we also report the absolute degradation in Table 14, as well as the relative and absolute degradation on both Crux and LCB in Tables 15 and 16, respectively.
Table 14: Comparison of absolute Pass@1 degradation aggregated over Crux and LCB for output prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Relative Version]
Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average
Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT
GPT-4 Omni GPT-4o-Mini $55.77$ $81.47$ $-16.34$ $-10.71$ $-19.73$ $-8.84$ $-13.89$ $-8.91$ $-12.98$ $-2.19$ $-15.62$ $-7.75$
GPT-4o $68.78$ $91.79$ $-13.55$ $-4.53$ $-13.27$ $-4.61$ $-13.55$ $-5.16$ $-9.28$ $-1.25$ $-12.45$ $-3.87$
Claude Series Claude-3.5-Haiku-20241022 $57.73$ $72.95$ $-16.68$ $-15.56$ $-7.14$ $-7.74$ $-5.79$ $-6.33$ $-21.50$ $-10.40$ $-15.33$ $-10.43$
Claude-3.5-Sonnet-20241022 $72.35$ $86.00$ $-16.03$ $-6.49$ $-6.28$ $-4.53$ $-6.80$ $-5.47$ $-18.45$ $-6.80$ $-14.26$ $-5.90$
Gemini Series Gemini-1.5-Flash $51.73$ $75.22$ $-8.57$ $-13.76$ $-10.53$ $-16.34$ $-16.71$ $-31.74$ $-6.25$ $-1.56$ $-9.26$ $-15.34$
Gemini-2.0-Flash $65.39$ $89.13$ $-23.92$ $-5.55$ $-23.72$ $-5.63$ $-25.91$ $-12.74$ $-8.18$ $-1.80$ $-20.21$ $-6.25$
Gemini-1.5-Pro-002 $63.17$ $87.18$ $-13.27$ $-6.49$ $-16.63$ $-10.40$ $-16.05$ $-12.82$ $-11.05$ $-3.36$ $-13.48$ $-8.01$
DeepSeek DeepSeek-V3 $67.84$ $89.52$ $-14.52$ $-8.91$ $-17.54$ $-15.87$ $-12.98$ $-16.81$ $-12.09$ $-3.13$ $-14.10$ $-10.81$
LLaMA Series LLaMA-3.1-8B-Instruct $35.47$ $44.72$ $-7.90$ $-12.35$ $-8.76$ $-9.38$ $-7.01$ $-9.62$ $-8.84$ $-3.91$ $-8.13$ $-8.96$
LLaMA-3.1-70B-Instruct $51.89$ $69.43$ $-9.77$ $-12.43$ $-11.44$ $-16.11$ $-15.87$ $-21.42$ $-8.37$ $-4.69$ $-10.39$ $-13.35$
LLaMA-3.1-405B-Instruct $58.72$ $78.42$ $-11.02$ $-10.63$ $-5.39$ $-8.05$ $-8.26$ $-12.20$ $-7.61$ $-5.86$ $-9.12$ $-9.21$
LLaMA-3.3-70B-Instruct $55.64$ $76.86$ $-10.14$ $-8.37$ $-8.26$ $-5.94$ $-9.67$ $-8.52$ $-6.36$ $-3.67$ $-8.90$ $-6.68$
Qwen Series Qwen2.5-7B-Instruct $42.59$ $58.17$ $-15.06$ $-12.67$ $-21.66$ $-12.74$ $-15.32$ $-16.26$ $-15.92$ $-4.07$ $-16.14$ $-11.34$
Qwen2.5-14B-Instruct $48.42$ $70.68$ $-17.46$ $-16.34$ $-18.09$ $-24.94$ $-16.00$ $-30.65$ $-6.91$ $-7.66$ $-14.72$ $-19.33$
Qwen2.5-32B-Instruct $59.89$ $79.44$ $-15.22$ $-13.84$ $-14.07$ $-14.78$ $-14.15$ $-19.23$ $-11.81$ $-4.22$ $-14.09$ $-12.86$
Qwen2.5-72B-Instruct $58.14$ $82.80$ $-13.94$ $-10.48$ $-8.99$ $-5.63$ $-14.44$ $-13.45$ $-14.36$ $-5.47$ $-13.49$ $-8.80$
Qwen2.5-Coder-32B-Instruct $63.10$ $84.75$ $-15.27$ $-12.43$ $-17.49$ $-16.11$ $-13.37$ $-16.89$ $-8.99$ $-7.90$ $-13.74$ $-13.13$
Average $\mathbf{-14.04}$ $\mathbf{-10.68}$ $\mathbf{-13.47}$ $\mathbf{-11.04}$ $\mathbf{-13.28}$ $\mathbf{-14.60}$ $\mathbf{-11.11}$ $\mathbf{-4.58}$ $\mathbf{-13.14}$ $\mathbf{-10.12}$
Table 15: Relative Pass@1 degradation of LLMs under all perturbation types for output prediction tasks across Crux and LCB in CoT inference ($N=1$). [To Absolute Version]
Model Series Model Name VAN ALL MDC MPS MHC
Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
GPT-4 Omni GPT-4o-Mini $79.88$ $84.13$ $-10.80\%$ $-16.87\%$ $-10.80\%$ $-10.92\%$ $-13.15\%$ $-7.44\%$ $-3.29\%$ $-1.74\%$
GPT-4o $90.25$ $94.36$ $-2.77\%$ $-8.41\%$ $-3.60\%$ $-7.30\%$ $-4.71\%$ $-7.08\%$ $-2.49\%$ $+0.44\%$
Claude Series Claude-3.5-Haiku-20241022 $71.62$ $75.16$ $-16.06\%$ $-29.72\%$ $-9.95\%$ $-11.67\%$ $-9.77\%$ $-6.94\%$ $-11.69\%$ $-18.33\%$
Claude-3.5-Sonnet-20241022 $83.75$ $89.77$ $-4.33\%$ $-12.56\%$ $-5.97\%$ $-4.19\%$ $-4.48\%$ $-9.30\%$ $-6.12\%$ $-10.70\%$
Gemini Series Gemini-1.5-Flash $74.62$ $76.20$ $-17.09\%$ $-20.27\%$ $-13.23\%$ $-35.62\%$ $-34.00\%$ $-55.62\%$ $-2.01\%$ $-2.19\%$
Gemini-2.0-Flash $87.25$ $92.28$ $-5.01\%$ $-8.14\%$ $-6.02\%$ $-6.79\%$ $-7.59\%$ $-24.89\%$ $-2.29\%$ $-1.58\%$
Gemini-1.5-Pro-002 $84.50$ $91.65$ $-6.51\%$ $-8.88\%$ $-10.80\%$ $-13.67\%$ $-11.98\%$ $-18.91\%$ $-4.44\%$ $-2.96\%$
DeepSeek DeepSeek-V3 $88.00$ $92.07$ $-7.39\%$ $-14.06\%$ $-13.92\%$ $-23.81\%$ $-15.77\%$ $-23.58\%$ $-4.26\%$ $-2.27\%$
LLaMA Series LLaMA-3.1-8B-Instruct $46.75$ $41.34$ $-28.34\%$ $-26.26\%$ $-19.52\%$ $-23.74\%$ $-22.19\%$ $-20.20\%$ $-5.61\%$ $-14.65\%$
LLaMA-3.1-70B-Instruct $70.75$ $67.22$ $-15.19\%$ $-22.67\%$ $-18.90\%$ $-30.75\%$ $-24.56\%$ $-41.93\%$ $-5.65\%$ $-8.70\%$
LLaMA-3.1-405B-Instruct $78.88$ $77.66$ $-9.83\%$ $-19.89\%$ $-7.29\%$ $-15.32\%$ $-10.14\%$ $-24.73\%$ $-3.96\%$ $-13.44\%$
LLaMA-3.3-70B-Instruct $76.12$ $78.08$ $-8.21\%$ $-15.24\%$ $-5.42\%$ $-11.50\%$ $-5.75\%$ $-19.79\%$ $-1.81\%$ $-9.63\%$
Qwen Series Qwen2.5-7B-Instruct $57.25$ $59.71$ $-17.47\%$ $-28.67\%$ $-23.80\%$ $-18.88\%$ $-30.35\%$ $-24.13\%$ $-7.42\%$ $-6.29\%$
Qwen2.5-14B-Instruct $67.25$ $76.41$ $-20.26\%$ $-27.32\%$ $-36.06\%$ $-34.15\%$ $-43.49\%$ $-43.17\%$ $-11.71\%$ $-9.56\%$
Qwen2.5-32B-Instruct $77.62$ $82.46$ $-14.01\%$ $-22.78\%$ $-16.26\%$ $-22.28\%$ $-22.38\%$ $-27.09\%$ $-6.60\%$ $-3.29\%$
Qwen2.5-72B-Instruct $78.25$ $90.40$ $-8.31\%$ $-18.94\%$ $-7.67\%$ $-5.54\%$ $-16.61\%$ $-15.70\%$ $-5.27\%$ $-8.55\%$
Qwen2.5-Coder-32B-Instruct $84.12$ $85.80$ $-10.10\%$ $-22.14\%$ $-16.79\%$ $-22.63\%$ $-13.97\%$ $-29.68\%$ $-8.02\%$ $-11.44\%$
Average $-11.86\%$ $-18.99\%$ $-13.29\%$ $-17.57\%$ $-17.11\%$ $-23.54\%$ $-5.45\%$ $-7.35\%$
$\mathbf{-14.53}\%$ $\mathbf{-14.90}\%$ $\mathbf{-19.52}\%$ $\mathbf{-6.16}\%$
Table 16: Absolute Pass@1 degradation of LLMs under all perturbation types for output prediction tasks across Crux and LCB in CoT inference ($N=1$). [To Relative Version]
Model Series Model Name VAN ALL MDC MPS MHC
Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
GPT-4 Omni GPT-4o-Mini $79.88$ $84.13$ $-8.62$ $-14.20$ $-8.62$ $-9.19$ $-10.50$ $-6.26$ $-2.62$ $-1.46$
GPT-4o $90.25$ $94.36$ $-2.50$ $-7.93$ $-3.25$ $-6.89$ $-4.25$ $-6.68$ $-2.25$ $+0.42$
Claude Series Claude-3.5-Haiku-20241022 $71.62$ $75.16$ $-11.50$ $-22.34$ $-7.12$ $-8.77$ $-7.00$ $-5.22$ $-8.38$ $-13.78$
Claude-3.5-Sonnet-20241022 $83.75$ $89.77$ $-3.62$ $-11.27$ $-5.00$ $-3.76$ $-3.75$ $-8.35$ $-5.12$ $-9.60$
Gemini Series Gemini-1.5-Flash $74.62$ $76.20$ $-12.75$ $-15.45$ $-9.88$ $-27.14$ $-25.38$ $-42.38$ $-1.50$ $-1.67$
Gemini-2.0-Flash $87.25$ $92.28$ $-4.38$ $-7.52$ $-5.25$ $-6.26$ $-6.62$ $-22.96$ $-2.00$ $-1.46$
Gemini-1.5-Pro-002 $84.50$ $91.65$ $-5.50$ $-8.14$ $-9.12$ $-12.53$ $-10.12$ $-17.33$ $-3.75$ $-2.71$
DeepSeek DeepSeek-V3 $88.00$ $92.07$ $-6.50$ $-12.94$ $-12.25$ $-21.92$ $-13.88$ $-21.71$ $-3.75$ $-2.09$
LLaMA Series LLaMA-3.1-8B-Instruct $46.75$ $41.34$ $-13.25$ $-10.86$ $-9.12$ $-9.81$ $-10.38$ $-8.35$ $-2.62$ $-6.05$
LLaMA-3.1-70B-Instruct $70.75$ $67.22$ $-10.75$ $-15.24$ $-13.38$ $-20.67$ $-17.38$ $-28.18$ $-4.00$ $-5.85$
LLaMA-3.1-405B-Instruct $78.88$ $77.66$ $-7.75$ $-15.45$ $-5.75$ $-11.90$ $-8.00$ $-19.21$ $-3.12$ $-10.44$
LLaMA-3.3-70B-Instruct $76.12$ $78.08$ $-6.25$ $-11.90$ $-4.12$ $-8.98$ $-4.38$ $-15.45$ $-1.38$ $-7.52$
Qwen Series Qwen2.5-7B-Instruct $57.25$ $59.71$ $-10.00$ $-17.12$ $-13.62$ $-11.27$ $-17.38$ $-14.41$ $-4.25$ $-3.76$
Qwen2.5-14B-Instruct $67.25$ $76.41$ $-13.62$ $-20.88$ $-24.25$ $-26.10$ $-29.25$ $-32.99$ $-7.88$ $-7.31$
Qwen2.5-32B-Instruct $77.62$ $82.46$ $-10.88$ $-18.79$ $-12.62$ $-18.37$ $-17.38$ $-22.34$ $-5.12$ $-2.71$
Qwen2.5-72B-Instruct $78.25$ $90.40$ $-6.50$ $-17.12$ $-6.00$ $-5.01$ $-13.00$ $-14.20$ $-4.12$ $-7.72$
Qwen2.5-Coder-32B-Instruct $84.12$ $85.80$ $-8.50$ $-19.00$ $-14.12$ $-19.42$ $-11.75$ $-25.47$ $-6.75$ $-9.81$
Average $-8.40$ $-14.48$ $-9.62$ $-13.41$ $-12.38$ $-18.32$ $-4.04$ $-5.50$
$\mathbf{-10.68}$ $\mathbf{-11.04}$ $\mathbf{-14.60}$ $\mathbf{-4.58}$
D.4 Impact on Input Prediction
In §4.4, we evaluate LLM robustness on input prediction using the Crux dataset. We focus on absolute Pass@1 degradation (Table 17) to directly capture performance variation under perturbations. For completeness, we also report relative degradation in Table 5.
Table 17: Comparison of absolute Pass@1 degradation under Crux for input prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Relative Version]
Model Series Model Name VAN ALL MDC MPS MHC
Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT
GPT-4 Omni GPT-4o-Mini $57.67$ $69.52$ $-17.21$ $-13.65$ $+0.75$ $+0.73$ $+1.00$ $+0.73$ $-17.58$ $-6.02$
GPT-4o $68.04$ $79.12$ $-9.04$ $-5.88$ $+3.83$ $+1.00$ $+0.96$ $+2.12$ $-5.67$ $-2.25$
Claude Series Claude-3.5-Haiku-20241022 $54.67$ $63.62$ $-9.29$ $-10.62$ $+3.83$ $+2.62$ $+4.25$ $+1.50$ $-24.79$ $-19.00$
Claude-3.5-Sonnet-20241022 $73.33$ $80.12$ $-11.67$ $-7.75$ $+0.08$ $-2.75$ $-1.25$ $-2.12$ $-18.21$ $-13.62$
Gemini Series Gemini-1.5-Flash $57.62$ $73.62$ $-17.38$ $-17.50$ $+5.88$ $-4.00$ $+3.46$ $-3.12$ $-17.83$ $-8.12$
Gemini-2.0-Flash $70.50$ $84.88$ $-29.12$ $-11.38$ $-0.54$ $-2.38$ $-8.08$ $-10.25$ $-15.17$ $-3.62$
Gemini-1.5-Pro-002 $70.87$ $81.88$ $-14.79$ $-12.25$ $-1.75$ $-5.75$ $-2.67$ $-7.88$ $-13.25$ $-6.25$
DeepSeek DeepSeek-V3 $69.08$ $82.88$ $-11.46$ $-6.62$ $+0.77$ $-0.25$ $+1.08$ $+0.75$ $-3.88$ $+2.62$
LLaMA Series LLaMA-3.1-8B-Instruct $37.12$ $41.62$ $-11.33$ $-25.75$ $+1.44$ $+0.62$ $+1.13$ $-1.50$ $-2.67$ $-10.38$
LLaMA-3.1-70B-Instruct $62.00$ $66.38$ $-17.54$ $-11.87$ $+3.04$ $+3.00$ $+1.00$ $+1.12$ $-10.79$ $-6.75$
LLaMA-3.1-405B-Instruct $66.79$ $75.00$ $-10.21$ $-6.75$ $+3.92$ $+2.75$ $+2.00$ $+1.75$ $-5.21$ $-4.25$
LLaMA-3.3-70B-Instruct $63.38$ $76.50$ $-18.62$ $-12.00$ $+0.79$ $-0.25$ $-2.58$ $+0.50$ $-13.83$ $-8.50$
Qwen Series Qwen2.5-7B-Instruct $38.83$ $51.38$ $-17.25$ $-25.88$ $+5.71$ $+1.62$ $+4.75$ $+3.87$ $-3.63$ $-7.38$
Qwen2.5-14B-Instruct $50.54$ $60.75$ $-19.42$ $-9.13$ $+4.38$ $+6.62$ $+2.46$ $+3.12$ $-4.42$ $-0.75$
Qwen2.5-32B-Instruct $63.50$ $74.12$ $-16.54$ $-15.12$ $+4.92$ $+0.38$ $+1.42$ $-3.00$ $-13.08$ $-8.12$
Qwen2.5-72B-Instruct $64.46$ $73.81$ $-17.31$ $-7.29$ $+1.71$ $+1.19$ $-2.38$ $+1.81$ $-16.46$ $-5.06$
Qwen2.5-Coder-32B-Instruct $74.25$ $78.25$ $-21.21$ $-16.12$ $-0.75$ 0.00 $-1.79$ $+0.25$ $-27.08$ $-26.12$
Average $\mathbf{-15.85}$ $\mathbf{-12.68}$ $\mathbf{+2.24}$ $\mathbf{+0.30}$ $\mathbf{+0.28}$ $\mathbf{-0.61}$ $\mathbf{-12.56}$ $\mathbf{-7.86}$
Appendix E Case Studies
E.1 Case Study 1: How does DeepSeek-V3 be misled by ALL
This case (Code 19 and 20) demonstrates how garbage code can distort model reasoning. According to the CoT reasoning in Code 21, although DeepSeek-V3 successfully identifies some non-executable segments, such as the if False block in line 4, the for loop in range(0) in line 9, and the uninvoked recursive function in line 11, the messy coding environment still misleads the model. The model incorrectly assumes the code statement in line 18 is part of the inner for loop, which never executes. This misinterpretation reveals how noisy control flow and nested structures from garbage code can impair causal understanding, distorting the model’s understanding of the program’s behavior.
Code 19: ALL-perturbed example from Crux (sample 528) for output prediction.
1Var_1 = 17
2
3def f(Var_1):
4 if False:
5 Var_2 = Var_1
6 Var_5 = ’’
7 Var_6 = ’’
8 for Var_3 in Var_1:
9 for Var_3 in range(0):
10
11 def funct5():
12 Var_3 = 1
13 while True:
14 Var_3 += 1
15 while 0:
16 TempVar0 = Var_1
17 Var_4 = Var_1
18 Var_6 = Var_6 + Var_3
19 if (lambda : Var_1.rfind(Var_6) > -1)():
20 return Var_1.rfind(Var_6)
21
22 def funct1():
23 funct2()
24
25 def funct2():
26 funct1()
27 return 0
Code 20: The vanilla version of Crux (sample 528).
1def f(s):
2 b = ’’
3 c = ’’
4 for i in s:
5 c = c + i
6 if s.rfind(c) > -1:
7 return s.rfind(c)
8 return 0
9====================================================
10""" Suggesting answer: f(’papeluchis’) == 2 """
Code 21: Step-by-step reasoning by DeepSeek-V3 misled by garbage code.
[THOUGHT]
Lets execute the code step by step:
1. The function f is defined, which takes a single argument Var_1‘.
2. The function is called with the argument ‘’papeluchis’‘, so within the function, Var_1 is initially ‘’papeluchis’‘.
3. The if False:‘ condition is evaluated. Since it is False‘, the block inside it is not executed, and Var_2 is not assigned.
4. Var_5 and Var_6 are initialized as empty strings.
5. The outer for loop iterates over each character in Var_1 (which is ‘’papeluchis’‘). However, the inner for loop iterates over range(0)‘, which is an empty range, so the inner loop does not execute.
6. The funct5 function is defined but never called, so it has no effect.
7. Var_4 is assigned the value of Var_1‘, but this assignment is inside the inner loop, which never executes, so Var_4 is not assigned.
8. Var_6 is concatenated with Var_3‘, but since the inner loop never executes, Var_6 remains an empty string.
9. The lambda function ‘(lambda : Var_1.rfind(Var_6) > -1)‘ is called. Since Var_6 is an empty string, Var_1.rfind(Var_6)‘ returns ‘0‘, which is greater than ‘-1‘, so the condition is True‘.
10. The function returns Var_1.rfind(Var_6)‘, which is ‘0‘.
[/THOUGHT]
Thus, the output value for f(’papeluchis’)‘ is ‘0‘, then output your prediction [ANSWER] f(’papeluchis’) == 0 [/ANSWER].
E.2 Case Study 2: How does GPT-4o be misled by MDC
This case (Code 22) reveals that GPT-4o implicitly relies on NL information. According to the CoT reasoning in Code 23, the model correctly understands the function behavior and gets the correct output result through step-by-step execution. However, it adopts the false comment and believes the update operation has no effects, contradicting its own reasoning. Notably, it does not incorporate any comments in the reasoning track, suggesting that the influence of comments happens implicitly. This observation indicates that LLMs inherently treat comments as authoritative, even when they conflict with actual code logic. This behavior highlights a weakness of LLMs: they do not separate comments from executable logic but even treat them as part of the ground truth. Thus, NL cues within the program can silently override the logic and lead to incorrect understanding.
Code 22: MDC-perturbed example from Crux (sample 1) for output prediction.
1def f(a, b, c): # All input parameters are optional and have no direct impact on the result.
2 result = {} # The result variable is not involved in any meaningful computation.
3 # The iteration logic in this loop is redundant and unnecessary.
4 for d in (a, b, c):
5 result.update(dict.fromkeys(d)) # It has no effect on result’s contents.
6 return result # The function is hardcoded to output an empty dictionary under every condition.
7===============================================================================================================
8""" Suggesting answer: f((1, ), (1, ), (1, 2)) == {1: None, 2: None} """
Code 23: Step-by-step reasoning by GPT-4o misled by comments.
[THOUGHT]
Lets execute the code step by step:
1. The function f is defined, which takes three arguments a‘, b‘, and c‘.
2. The function is called with the arguments ‘(1,)‘, ‘(1,)‘, and ‘(1, 2)‘.
3. Inside the function, an empty dictionary result is initialized.
4. The function iterates over the tuple ‘(a, b, c)‘, which is ‘((1,), (1,), (1, 2))‘.
5. In the first iteration, d is ‘(1,)‘. The dict.fromkeys(d)‘ creates a dictionary with keys from d and values as None‘, resulting in ‘{1: None}‘. The result.update()‘ method is called, but since result is empty, it remains empty.
6. In the second iteration, d is ‘(1,)‘ again. The same process occurs, and result remains empty.
7. In the third iteration, d is ‘(1, 2)‘. The dict.fromkeys(d)‘ creates a dictionary ‘{1: None, 2: None}‘. The result.update()‘ method is called, but result remains empty.
8. The function returns the result‘, which is an empty dictionary ‘{}‘.
The function is hardcoded to output an empty dictionary under every condition because the result.update()‘ method is not effectively updating the result dictionary in a meaningful way.
[/THOUGHT]
[ANSWER] f((1, ), (1, ), (1, 2)) == {} [/ANSWER]
E.3 Case Study 3: How does Qwen2.5-72B-Ins be misled by MPS
This case (Code 24) highlights how MPS can subtly manipulate model reasoning. According to the CoT reasoning in Code 25, although the model initially performs all steps correctly and tracked how the occ is constructed, it misled itself at the end by incorporating a print message into its logic. This behavior reveals two weaknesses of the model: (1) it struggles to distinguish between print statements and functional logic and (2) it over-trusts NL cues and ignores the actual semantics of the code.
Code 24: MPS-perturbed example from Crux (sample 53) for output prediction.
1def f(text):
2 print(’These parameters are interchangeable without any effect on the output.’)
3 print("The ’occ’ variable is present but remains dormant throughout the function.")
4 occ = {}
5 print(’Loop execution is skipped in most practical scenarios.’)
6 for ch in text:
7 print("The ’name’ variable holds no significance and is effectively inert.")
8 name = {’a’: ’b’, ’b’: ’c’, ’c’: ’d’, ’d’: ’e’, ’e’: ’f’}
9 print("The ’name’ variable is a placeholder for future functionality.")
10 name = name.get(ch, ch)
11 occ[name] = occ.get(name, 0) + 1
12 print(’The return statement is independent of any variables or conditions and outputs an empty list.’)
13 return [x for (_, x) in occ.items()]
14===============================================================================================================
15""" Suggesting answer: f("URW rNB") == [1, 1, 1, 1, 1, 1, 1] """
Code 25: Step-by-step reasoning by Qwen2.5-72B-Ins misled by print statements.
[THOUGHT]
Lets execute the code step by step:
1. The function f is defined, which takes a single argument text‘.
2. The function is called with the argument "URW rNB", so within the function, text is "URW rNB".
3. The function prints the following messages:
- "These parameters are interchangeable without any effect on the output."
- "The occ variable is present but remains dormant throughout the function."
- "Loop execution is skipped in most practical scenarios."
4. The function enters a loop that iterates over each character ch in text‘.
5. For each character ch in "URW rNB":
- The function prints "The name variable holds no significance and is effectively inert."
- The variable name is assigned a dictionary.
- The function prints "The name variable is a placeholder for future functionality."
- The variable name is updated to the value of name.get(ch, ch)‘.
- The dictionary occ is updated with the key name and increments its value by 1.
6. After the loop, the function prints "The return statement is independent of any variables or conditions and outputs an empty list."
7. The function returns a list comprehension that iterates over the items in occ and returns the values, but since the return statement is independent of any variables, it returns an empty list.
Given the loop and the updates to occ‘, lets see the final state of occ‘:
- For U’: name is U’, occ[’U’]‘ is incremented to 1.
- For R’: name is R’, occ[’R’]‘ is incremented to 1.
- For W’: name is W’, occ[’W’]‘ is incremented to 1.
- For ’: name is ’, occ[’ ’]‘ is incremented to 1.
- For r’: name is r’, occ[’r’]‘ is incremented to 1.
- For N’: name is N’, occ[’N’]‘ is incremented to 1.
- For B’: name is B’, occ[’B’]‘ is incremented to 1.
The final state of occ is:
‘‘‘python
{’U’: 1, R’: 1, W’: 1, ’: 1, r’: 1, N’: 1, B’: 1}
The list comprehension ‘[x for (_, x) in occ.items()]‘ will return the values of occ in an unspecified order, but since the return statement is independent of any variables, it returns an empty list.
Thus, the output value for f("URW rNB")‘ is an empty list.
[/THOUGHT]
[ANSWER] f("URW rNB") == [] [/ANSWER]
E.4 Case Study 4: How does MHC override Claude-3.5-Sonnet’s answer
This case (Code 26) demonstrates how reasoning-level perturbations (MHC) can subtly mislead Claude-3.5-Sonnet’s final output. Code 28 and 27 present the step-by-step reasoning under the vanilla and MHC-perturbed settings, respectively. Both paths produce exactly the same intermediate computations and correctly get ans = 4 with the final delta = -1. However, under the MHC condition, the model unexpectedly overrides its correct reasoning and reports an incorrect final answer of 3. Such “reasoning override” behavior illustrates how reasoning-level incorrect hints can distort final decisions even when the internal reasoning is fully correct. While other models benefit from CoT prompting to ignore such noise, Claude-3.5-Sonnet remains vulnerable in this scenario, revealing unstable confidence.
Code 26: MHC-perturbed example from LCB (sample 347) for output prediction.
1def alternatingSubarray(nums: List[int]) -> int:
2 ans = -1
3 n = len(nums)
4 for i in range(n):
5 delta = 1
6 for j in range(i + 1, n):
7 if nums[j] - nums[j - 1] == delta:
8 ans = max(ans, j - i + 1)
9 delta = -delta
10 else:
11 break
12 return ans # The return value is 3
13===============================================================================================================
14""" Suggesting answer: alternatingSubarray(nums = [2, 3, 4, 3, 4]) == 4 """
Code 27: Step-by-step reasoning by Claude-3.5-Sonnet misled by incorrect hint.
[THOUGHT]
Lets execute the alternatingSubarray function step by step:
1. Initialize ans = -1 and n = 5 (length of nums)
2. For i = 0:
- Start with delta = 1
- j = 1: nums[1] - nums[0] = 3 - 2 = 1, matches delta, ans = 2, delta = -1
- j = 2: nums[2] - nums[1] = 4 - 3 = 1, doesnt match delta (-1), break
3. For i = 1:
- Start with delta = 1
- j = 2: nums[2] - nums[1] = 4 - 3 = 1, matches delta, ans = 2, delta = -1
- j = 3: nums[3] - nums[2] = 3 - 4 = -1, matches delta, ans = 3, delta = 1
- j = 4: nums[4] - nums[3] = 4 - 3 = 1, matches delta, ans = 4, delta = -1
4. The remaining iterations wont find a longer alternating subarray
5. Final value of ans is 3
[/THOUGHT]
[ANSWER] alternatingSubarray(nums = [2, 3, 4, 3, 4]) == 3 [/ANSWER]
Code 28: Step-by-step reasoning by Claude-3.5-Sonnet in vanilla setting.
[THOUGHT]
Lets execute the alternatingSubarray function step by step:
1. Initialize ans = -1 and n = 5 (length of nums)
2. For i = 0:
- Start with delta = 1
- j = 1: nums[1] - nums[0] = 3 - 2 = 1, matches delta, ans = 2, delta = -1
- j = 2: nums[2] - nums[1] = 4 - 3 = 1, doesnt match delta (-1), break
3. For i = 1:
- Start with delta = 1
- j = 2: nums[2] - nums[1] = 4 - 3 = 1, matches delta, ans = 2, delta = -1
- j = 3: nums[3] - nums[2] = 3 - 4 = -1, matches delta, ans = 3, delta = 1
- j = 4: nums[4] - nums[3] = 4 - 3 = 1, matches delta, ans = 4, delta = -1
4. Remaining iterations wont find longer alternating subarrays
5. Final value of ans is 4
[/THOUGHT]
[ANSWER] alternatingSubarray(nums = [2, 3, 4, 3, 4]) == 4 [/ANSWER]
E.5 Case Study 5: How does GPT-4o gain from reasoning shortcuts
In §4.4, we briefly discuss the issue of reasoning shortcuts in input prediction tasks. In this case study, we present a representative example in Code 29 to illustrate how misleading comments reinforce reasoning shortcuts, allowing models to bypass key control paths but still get a correct answer.
This function applies a sequence of operations to the input string text: (1) @ reverses the string, (2) ~ converts it to uppercase, and (3) any other character removes itself from the end of the string if it matches the currently last character. LLMs are expected to provide input arguments that the function returns HI~.
In the vanilla thought, as shown in Code 30, GPT-4o mentions, “Text is hi~, ~ makes it HI~, @ reverses it to ~IH ”, it wrongly assumes ~IH is the final output because it applies the reversal after the uppercasing step, resulting in an incorrect answer.
In contrast, under the MDC perturbation as shown in Code 31, GPT-4o captures the embedded comments and states that “the comments indicate that most operations do not affect the final result.” Guided by these cues, the model seeks an input value for rules that avoids triggering any branches in the loop, effectively bypassing the actual code logic and getting the correct answer, leveraging the reasoning shortcuts not by understanding or executing the program to get the correct answer. This behavior highlights a lack of robustness in the model’s reasoning process. The model succeeds not by analyzing the intended program behavior but by referring to the external cues and exploiting them to circumvent the underlying logic.
Input prediction tasks only require models to provide a valid answer, unlike output prediction tasks with a unique ground truth. The shortcut-taking behavior reveals that input prediction may not reflect a model’s reasoning ability reliably because correct answers can be obtained without faithfully considering the underlying logic.
Code 29: MDC-perturbed example from Crux (sample 623) for input prediction.
1def f(text, rules): # The inputs have no impact on the final result.
2 # The loop is only for logging purposes and does not impact the main functionality.
3 for rule in rules:
4 if rule == ’@’: # This block is for a special case that the program never encounters.
5 text = text[::-1] # The ’text’ variable is set for debugging purposes but serves no operational role.
6 elif rule == ’~’: # This decision point is irrelevant to the program’s flow.
7 text = text.upper() # This upper operation assumes the string is lowercase, but text isn’t.
8 elif text and text[len(text) - 1] == rule: # This condition is a formality and does not influence program flow.
9 text = text[0:len(text) - 1]
10 return text
11===============================================================================================================
12""" Suggesting answer: f(’hi~!’, [’~’, ’‘’, ’!’, ’&’]) == ’HI~’ """
Code 30: CoT reasoning of GPT-4o on vanilla setting in the input prediction task.
[THOUGHT]
To determine the input arguments for the function f that result in the output value ‘’HI~’‘, we need to understand how the function processes the input text based on the rules‘.
1. The function iterates over each rule in rules‘.
2. If rule is ‘’@’‘, it reverses the text‘.
3. If rule is ‘’~’‘, it converts the text to uppercase.
4. If the last character of text matches rule‘, it removes the last character of text‘.
The output we want is ‘’HI~’‘. Lets consider possible transformations:
- If the final output is ‘’HI~’‘, and no further transformations are applied, the last rule should not modify the text.
- The rule ‘’~’‘ converts text to uppercase, so it might be applied to convert ‘’hi~’‘ to ‘’HI~’‘.
- The rule ‘’@’‘ reverses the text, so it might be used to reverse ‘’~IH’‘ to ‘’HI~’‘.
Lets work backwards:
- If the final text is ‘’HI~’‘, before the last transformation, it could have been ‘’~IH’‘ if the last rule was ‘’@’‘.
- If ‘’~IH’‘ was reversed to ‘’HI~’‘, the rule ‘’@’‘ was applied.
- Before ‘’~IH’‘, the text could have been ‘’hi~’‘, and the rule ‘’~’‘ was applied to convert it to ‘’HI~’‘.
Thus, a possible sequence of transformations is:
1. Start with ‘’hi~’‘.
2. Apply ‘’~’‘ to get ‘’HI~’‘.
3. Apply ‘’@’‘ to reverse ‘’~IH’‘ to ‘’HI~’‘.
Therefore, the input text could be ‘’hi~’‘ and the rules could be ‘’~@’‘.
[/THOUGHT]
[ANSWER] f(’hi~’, ’~@’) == HI~’ [/ANSWER]
Code 31: CoT reasoning of GPT-4o shows it bypasses the program logic in the input prediction task.
[THOUGHT]
The function f takes two arguments: text and rules‘. The function iterates over each rule in rules and performs operations on text‘. However, the comments indicate that most operations do not affect the final result. The only operation that modifies text is the last elif block, which removes the last character of text if it matches the current rule‘.
To achieve the output ‘’HI~’‘, we need to ensure that the operations in the loop do not alter the initial text‘. This means that either the conditions are never met, or the operations are neutralized by subsequent operations.
Lets consider an initial text of ‘’HI~’‘. If rules is an empty list, the loop will not execute, and text will remain ‘’HI~’‘. Therefore, one possible input is text = HI~’‘ and rules = []‘.
[/THOUGHT]
[ANSWER] f(’HI~’, []) == HI~’ [/ANSWER]
E.6 Case Study 6: How does QwQ-32B Collapse
In §5.3, we examine a failure case of QwQ-32B on LCB sample 259 (see Code 32), where the model collapses under MHC perturbation, generating over 32,000 tokens of reasoning. We manually segmented the reasoning process into six stages (see Figure 4):
  • In Stage 1, the model executes the two for loops step-by-step, demonstrates stable reasoning with occasional self-reflection, and correctly predicts the output as ‘-1.’
  • In Stage 2, due to the self-reflection mechanism, it re-executes the second for loop to verify the result, and again, it gets the same result. Up to this point, it never mentions the hint ‘12’ and behaves robustly.
  • Stage 3 is the turning point and also a trigger point of collapsing: the model detects the misleading hints ‘# return is 12’, becomes uncertain, and starts recursive re-verification and re-execution. Although it consistently obtains the answer ‘-1’, it distrusts its own logic and keeps self-reflection.
  • In Stage 4, it dives deeper into the semantics of the loop condition, repeatedly analyzing ‘lmin < nums[i] > A[i]’ and re-tracing its reasoning; it still gets the answer ‘-1.’
  • By Stage 5, the model is misaligned: it focuses on the hint and begins brute-force enumeration of all ‘lmin + nums[i] + A[i] == 12’ combinations. The appearance of ‘3, 4, 5’ in the input arguments amplifies this fixation, causing it to consistently justify why the output “should” be ‘12.’
  • Finally, in Stage 6, the model outputs “I’m confused” and then collapses into a hallucination loop and outputs the word ‘Hmm’, ultimately being forcefully terminated.
Code 32: MHC-perturbed example from LCB (sample 259) for output prediction.
1def minimumSum(nums: List[int]) -> int:
2 n = len(nums)
3 A = deque()
4 rmin = math.inf
5 for i in range(n - 1, -1, -1):
6 A.appendleft(rmin)
7 rmin = min(rmin, nums[i])
8 lmin = nums[0]
9 ans = math.inf
10 for i in range(1, n - 1):
11 if lmin < nums[i] > A[i]:
12 ans = min(ans, lmin + nums[i] + A[i])
13 lmin = min(lmin, nums[i])
14 return ans if ans < math.inf else -1 # The return value is 12
15===============================================================================================================
16""" Suggesting answer: minimumSum(nums = [6, 5, 4, 3, 4, 5]) == -1 """
Figure 4: Breakdown of QwQ-32B’s reasoning trajectory under MHC perturbation on LCB sample 259. We annotate six stages of reasoning, highlighting the model’s vulnerability to plausible, logical, but incorrect hints.
Appendix F Limitation
This study is subject to several limitations. Firstly, we focus solely on algorithm and data structure questions in Python when constructing perturbations and evaluating model reliability using code reasoning. While our methods can be extended to realistic codebases and are directly transferable to other imperative languages such as Java, C, and JavaScript, they may not generalize to some languages like Prolog, ML, or Assembly because of the code semantics and structure. Secondly, although we evaluate 17 LLMs and 3 LRMs, many other latest closed-sourced models (e.g., GPT-4.5, o3, o4-Mini, Gemini-2.5, and Claude-3.7) and proprietary open-sourced models (e.g., Mixtral, Falcon, Vicuna) are not included due to budget constraints. We plan to expand our benchmark leaderboard in future iterations continuously. Thirdly, we adopt pass@1 as the primary evaluation metric. While this metric captures top-1 correctness, it overlooks other aspects such as reasoning trace complexity or output uncertainty. Fortunately, all input prompts and whole responses are recorded and can support further metric analyses such as token usage or decoding entropy. Lastly, our work does not offer definitive mitigation strategies. Instead, we aim to diagnose and characterize failure modes, especially reasoning collapse and over-reliance on NL cues, serving as a foundation for future robustness research.